### v1

# Complements of discriminants of real parabolic function singularities. II

Abstract.
We provide the complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$ and $J_{10}$. Thus, we prove (and improve in one particular case) all corresponding conjectures from the previous work [12] with the same title.
Key words and phrases:
discriminant, singularity, surgery, versal deformation, isotopy classification
1991 Mathematics Subject Classification:
Primary: 14Q30. Secondary: 14B07, 14P25
1. Introduction
Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular. These varieties appear in many problems of PDE theory and integral geometry (often under the name of wavefronts), see e.g. [1], [2], [3], [10], [11].
Many important functions in these theories are defined by integral transforms and are regular outside the discriminant varieties of deformations of corresponding generating functions. Their analytic behavior near the singular points of the discriminants depends on the singularity type of these functions and on the connected component of the complement of the discriminant. Thus, enumerating these components is one of the first steps in the study of their analytical properties.
The list of function singularities begins with simple singularities, which were distinguished by V. Arnold, see [2]. E. Looijenga [8] implicitly listed the local connected components of complements of their discriminants. O. Viro [18] listed all possible topological types of non-discriminant functions representing these components. (By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood. The topological type is an invariant of the components of the discriminant complements.) In [13], it was proven that each topological type of this list represents exactly one connected component of the discriminant complement.
The next important class of function singularities are the parabolic ones, which were also distinguished by V. Arnold. The classification of complex parabolic singularities of functions consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$; the classification of their real forms has more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
In [12], a system of invariants of the connected components of the complements of the discriminant varieties of real function singularities was described. The complete lists of the possible values of these invariants for parabolic singularities were found, and all of these values were realized by concrete polynomials. We also conjectured that the resulting lists of connected components thus found in [12] are complete. These lists consist of 7, 14, 52, 11, and 33 components for the singularities of classes $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two mutually symmetric sets of perturbations of the $J_{10}^{1}$ singularities indicated in [12] consist of pairs of different connected components. Thus, the number 11 above is replaced by 13. These results are summarized in Theorems 5, 13, 18, 28, and 31 below. The classes $P_{8}^{*}$ are not considered yet.
After the publication of [12] and [13], I discovered that O. Viro [18] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities (as well as of simple singularities).
Table 1. Numbers of isotopy and topological classes of non-discriminant perturbations of $X_{9}$ and $J_{10}$ singularities
Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$
Topology 6 12 50 10 31
Isotopy 7 14 52 13 33
It turns out that the “isotopy” classification of parabolic singularity perturbations considered in the present work is only slightly stricter than the topological classification, see Table 1.
E. Looijenga [8] proved that all local components of discriminant complements of simple singularities are contractible. Conversely, we show that at least some three such components of parabolic singularities of classes $X_{9}^{+}$ and $X_{9}^{-}$ have infinite one-dimensional homology groups, see Theorem 12 on page 12.
An application to hyperbolic PDE’s. In the theory of hyperbolic PDE’s, the local Petrovskii lacunas are the local connected components of the complements of wavefronts, in which the wave functions behave regularly, see e.g. [1], [10].
In [11], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete. The enumeration of all local components of wavefront complements given below proves this conjecture for all parabolic singularities of classes $X_{9}$ and $J_{10}$, see § 4.
2. Results for $X_{9}^{*}$ classes
Table 2. Normal forms of $X_{9}$ and $J_{10}$ singularities in two variables
Notation Normal form Restriction
$X_{9}^{+}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ $|A|<1$
$X_{9}^{1}$ $xy(x^{2}+2Axy+y^{2})$ $|A|<1$
$X_{9}^{2}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ A<1A<-1
$J_{10}^{1}$ $(x^{2}+y^{2})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $(x^{2}-y^{2})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
Definition 1.
A function singularity is a germ of a $C^{\infty}$-smooth function ${\mathbb{R}}^{2}\to{\mathbb{R}}$ at its critical point.
Real parabolic function singularities of type $X_{9}$ form four families $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$. Any function singularity of classes $X_{9}^{+}$, $X_{9}^{1}$ or $X_{9}^{2}$ has the normal form indicated in Table 2 in appropriate local coordinates centered at the critical point. The class $X_{9}^{-}$ (and all considerations concerning it) can be reduced to the case $X_{9}^{+}$ by multiplying by 1-1.
The standard miniversal deformations of functions of all three classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$ have the same form: they are families of polynomials
(1) $f_{A}(x)+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}x^{2}+\lambda_{5}xy+\lambda_{6}y^{2}+\lambda_{7}x^{2}y+\lambda_{8}xy^{2},$
depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 2. By the versality theory (see, e.g., [2]), the study of all other deformations of these singularities can be reduced to these.
Definition 2.
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (1) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ consists of four different lines in ${\mathbb{C}}^{2}$.
The spaces $\Phi_{+}$, $\Phi_{1}$, and $\Phi_{2}$ consist of all polynomials of the form (1) with principal parts of classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$, respectively.
The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$. The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value. These varieties divide the spaces $\Phi_{*}$ into connected components. The enumeration of these components is our goal in this section.
In §2 of [12] we introduced the set-valued invariant of these components. Let us briefly review this construction.
The virtual function associated with a generic real polynomial is a collection of certain topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set. Namely, these data of a function $f(x,y)$ consist of
  • the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented basic vanishing cycles in the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
  • the string of intersection indices of these cycles with the set of real points,
  • the Morse indices of real critical points of $f$, and
  • the numbers of positive and negative critical values of $f$.
The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions. The formal graph of a singularity class $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips. The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries crossing the discriminant are removed. The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component $S(f)$ of the formal graph containing the virtual functions associated with this polynomial.
These objects have the following obvious property.
Proposition 1.
If two polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same. $\Box$
Enumerating the virtual components of a finite formal graph is a combinatorial problem that can be solved by a computer program. The work [12] is essentially a list of the results of these computations for all parabolic singularities and of their immediate consequences in the terms of real functions. The remaining problem is to find, for each virtual component, all real connected components with this value of the set-valued invariant.
It follows from the results of [5] that, for parabolic singularities, all standard flips of virtual functions can be realized by real surgeries of generic real functions. In particular, the following statement holds.
Proposition 2 (see [15], Proposition 3).
For any class of polynomials $\Phi_{+}$, $\Phi_{1}$, or $\Phi_{2}$,
a)) every virtual function of this class is associated with a generic real polynomial $\tilde{f}$ of the same class;
b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the complement of the discriminant. $\Box$
Remark 3.
In the statement of Proposition 3 of [15], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four polynomials with non-degenerate principal homogeneous parts. However, the entire construction proving this proposition can be performed within the spaces of the versal deformations (1). Thus, it also proves Proposition 2.
These considerations enable us to estimate the number of isotopy classes of generic Morse perturbations associated with any given virtual function. These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately. Consequently, we obtain estimates of the number of connected components of the complement of a real discriminant which are associated with each virtual component. Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight. In the case $X_{9}^{1}$ the analogous group is of order four. Then, for each virtual component in the list from [12], we compare all real components represented by the polynomials associated with some its virtual function and forming an orbit of this group, and determine which of these components are identical. This comparison provides the results of the present work concerning the $X_{9}$ singularities. The strategy for the $J_{10}$ singularities is exactly the same.
Remark 4.
In [15], [16], and [17], we use the same strategy with a different definition of virtual components. In these works, we classify the Morse perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions. Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
Definition 3.
For any generic polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$, $\mbox{Ind}(f)$ is the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
It is easy to see that this number is an invariant of the connected components of the discriminant complements.
2.1. Computations for the $X_{9}^{+}$ singularities
In this subsection, we prove the following statement, conjectured in [12].
++$(a)$
(b)(b)++
(c)(c)++
++$(d)$
(e)(e)++
++$(f)$
++$(g)$
Figure 1. Perturbations of $X_{9}$ singularity
Theorem 5.
There are exactly seven connected components in the space of non-discriminant functions of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials representing these components are shown in Fig. 1.
Remark 6.
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine isotopy classes of bounded degree four algebraic plane curves. In our case, however, the isotopies are allowed only in a much smaller space $\Phi_{+}$. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
The starting point of the proof of Theorem 5 is the following proposition.
Proposition 7 (see [12]).
There are exactly seven virtual components of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials associated with virtual functions representing these virtual components are shown in Fig. 1. $\Box$
Definition 4.
Denote by $G_{0}$ the group of linear transformations of ${\mathbb{R}}^{2}$ generated by reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(y,x)$
Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight. This group acts on the space $\Phi_{+}$ of polynomials (1) with principal parts of class $X_{9}^{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$). This action preserves the set of generic polynomials.
Proposition 8.
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Phi_{+}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of class $\Phi_{+}$ containing $f$ and $\tilde{f}$ are mapped to each other by an element of the group $G_{0}$.
Definition 5.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part $x^{4}+2Ax^{2}y^{2}+y^{4}$, denote by ${\bf T}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of ${\mathbb{R}}^{2}$. (This is the translation by the vector $\frac{1}{4}(a,b),$ where $a$ and $b$ are the coefficients of $f$ at the monomials $x^{3}$ and $y^{3}$).
Lemma 9 (see [15], Lemma 4).
A function of the form ((1)) with a positive principal homogeneous part has a critical point of class $E_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{27}{256}$ if and only if it is one of the four polynomials
(2) $\hskip 82.51282pt{\bf T}(x^{4}\pm x^{3}+y^{4})\qquad\mbox{ or }\qquad{\bf T}(x^{4}+y^{4}\pm y^{3})\ .\hskip 82.51282pt\Box$
Proof of Proposition 8. This proof almost coincides with the proof of Proposition 3 in §3.5 of [15]. In the next paragraph, we present the outline of this proof; we refer to [15] for all common details.
Using the Lyashko–Looijenga covering (in the form described in [5]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that commutes with the Lyashko–Looijenga map: the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this reflection. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this reflection, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads to the previous situation. $\Box$
By Proposition 8, any virtual component mentioned in Proposition 7 is associated with at most eight real connected components. Let us prove that in fact each virtual component represented by a polynomial shown in Fig. 1 is associated with a single real connected component.
Pictures $(a)$, $(b)$, $(c)$, $(f)$ of Fig. 1 are realized, respectively, by four polynomials
$x^{4}+y^{4}+t,$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\hskip 23.49976pt\alpha\in(0,1/4),$
$x^{4}+y^{4}-t,$
$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(1/4,1/2),$
with arbitrary $t>0$. All of these polynomials are invariant under the entire group $G_{0}$.
Picture (d) can be represented by the polynomial
(3) $x^{4}+2Ax^{2}y^{2}+y^{4}+2x^{2}-2y^{2}+1/2$
with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
(4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2$
The polynomials (3) and (4) are connected by the family of polynomials
(5) $x^{4}+2Ax^{2}y^{2}+y^{4}+2(x\cos\varphi+y\sin\varphi)^{2}-2(x\sin\varphi-y\cos\varphi)^{2}+1/2,$
depending on the parameter $\varphi\in[0,\pi/2]$.
Lemma 10.
If $A<1$ is sufficiently close to $1$, then all polynomials of this family are non-discriminant, in particular their zero-level sets are all smooth and isotopic to each other.
Proof. If we replace $A$ by 1 in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$. Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle. If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $x^{2}+y^{2}\leq 2$ and maintains its topological structure in this domain for any $\varphi$. It also remains empty outside this domain. $\Box$
Thus, the connected component of the space of non-discriminant polynomials containing (3) is also invariant under the transformation $(x,y)\mapsto(y,x)$.
The topological type of the curve shown in picture $(e)$ can be realized by any of the four polynomials
(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa)$
and
(7) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(y^{3}-3x^{2}y)+\varkappa)$
with arbitrary $A\in(-1,1)$ and sufficiently small $\varkappa>0$ (where the restriction on $\varkappa$ depends on $A$). These four polynomials form an orbit of the group $G_{0}$. All of them belong to the one-parameter family of polynomials
(8) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}+\mbox{Re}(e^{i\tau}(x+iy)^{3}))+\varkappa),$
$\tau\in[0,2\pi]$. For $A=1$, all these polynomials are obtained from each other by the rotations of ${\mathbb{R}}^{2}$. They are bounded from below, have a critical point of class $D_{4}^{-}$ with zero critical value at the origin and three real minimum points with critical value $\frac{27}{256}+\varkappa$, and no other critical points. Therefore, for $\varkappa<\frac{27}{256}$ all of them realize the topological picture $(e)$. Analogously to Lemma 10, taking $A<1$ very close to 1, we will not change these topological pictures for any $\tau$. So, the connected component of the discriminant complement containing these polynomials is invariant under the action of the group $G_{0}$.
The picture $(g)$ can be realized by either of the four polynomials
(9) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10x(x^{2}-3y^{2})+15(x^{2}+y^{2})-5)$
and
(10) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10y(y^{2}-3x^{2})+15(x^{2}+y^{2})-5).$
with sufficiently small $\varepsilon>0$. These four polynomials also form an orbit of the group $G_{0}$.
The family of polynomials
(11) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}+10\mbox{Re}\left(e^{i\tau}(x+iy))^{3}\right)+15(x^{2}+y^{2})-5),$
$\tau\in[0,2\pi],$ contains all four polynomials (9), (10) for $\tau=0,$ $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. All elements of this family lie in the same connected component of the complement of the discriminant. This completes the proof of Theorem 5. $\Box$
Remark 11.
To realize the pictures ($d$), ($e$), and ($g$) by arbitrarily small perturbations of the $X_{9}^{+}$ singularities, we can multiply by $t^{4-k}$ all monomials of degree $k$ in the formulas (3), (6), and (9).
The 1-cohomology of discriminant complements. Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. Let $C$ be a connected component of the discriminant complement in the space $\Phi_{+}$ consisting of polynomials with $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(C)$ is well-defined as follows. For any such polynomial $f$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $C$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $C\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
Define the class $W\in H^{1}(C)$ as the class obtained by this map from the generator of the group $H^{1}(B({\mathbb{R}}^{2},k))$.
Theorem 12 (cf. [14], § 1.8.3).
The class $W$ takes all integer values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
Proof. These cycles are the families (5), (8), and (11). $\Box$
2.2. The components for the $X_{9}^{1}$ singularities
Define the group $G_{1}$ as the group of linear transformations of ${\mathbb{R}}^{2}$ generated by the symmetries $(x,y)\mapsto(y,x)$ and $(x,y)\mapsto(-x,-y)$. Clearly, this group is isomorphic to ${\mathbb{Z}}_{2}^{2}$. It acts on the space of polynomials $\Phi_{1}$ and preserves the set of generic polynomials. Polynomials obtained from each other by this action have equal associated virtual functions and so are not separated by the set-valued invariant.
Additionally, the involution
(12) $f(x,y)\mapsto-f(x,-y)$
acts on the space of polynomials (1) of class $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement. This permutation has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components related by this involution is always 1-1. This involution naturally extends to the set of associated virtual components: each virtual component associated with the polynomials from the connected component containing the polynomial $f(x,y)$ goes to a virtual component associated with the polynomials isotopic to f(x,y)-f(x,-y). The resulting involution on the set of virtual components of class $\Phi_{1}$ also has no fixed elements.
For any polynomial $f$ with a principal homogeneous part of the form
(13) $xy(x^{2}+2Axy+y^{2}),\qquad|A|<1,$
its zero-level set goes to the infinity along four branches: two vertical and two horizontal. If the polynomial is non-discriminant, then each vertical branch is matched with the horizontal branch belonging to the same component of the zero-level set. The involution (12) acts on the zero-level sets as the reflection in the line $\{y=0\}$. In particular, it always alters the matching of the vertical and horizontal branches.
++--++$(a)$
++--++$(b)$
++--++$(c)$
++--++$(d)$
++--++$(e)$
++--++$(f)$
++--++$(g)$
Figure 2. Perturbations of $X_{9}^{1}$ singularity
Theorem 13.
There are exactly 14 connected components of the complement of the discriminant in the space of polynomials ((1)) of class $\Phi_{1}$. The shapes of the zero-level sets of polynomials representing some seven of them are shown in Fig. 2, and the remaining seven can be obtained from these by the involution ((12)).
The proof of this theorem is based on the following two propositions.
Proposition 14 (see [12]).
There are exactly ten virtual components of type $X_{9}^{1}$. They split into five pairs of virtual components related to each other by the involution induced by ((12)). The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d), and ((f)) of Fig. 2. $\Box$
Proposition 15.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{1}$ are associated with the same virtual function, then their connected components in the space of generic polynomials are moved to each other by an element of the group $G_{1}$.
Definition 6.
For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part (13), denote by ${\bf\tilde{T}}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of the plane. (This is the translation by the vector $(\alpha,\beta),$ where $\alpha$ and $\beta$ are the coefficients of $f$ at the monomials $y^{3}$ and $x^{3}$).
Lemma 16.
A polynomial of the form ((1)) with the principal part ((13)) has a critical point of class $E_{7}$ with zero critical value and a critical point of class $A_{2}$ with critical value $3^{\frac{15}{2}}2^{-8}$ if and only if it is one of the following four polynomials:
(14) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm x^{3})$
or
(15) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm y^{3})\ .$
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [15], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (13). In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $ax^{3}$ or $ay^{3}$, $a\neq 0$. The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 15. The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with one simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [15], § 4.5). Therefore, the statement follows immediately from the fact that all points (14) and (15) belong to the same orbit of the group $G_{1}$. $\Box$
Proof of Theorem 13. Proposition 15 implies that any virtual component of the formal graph of the $X_{9}^{1}$ singularity is associated with at most four real connected components mapped to each other by the group $G_{1}$. Below we prove that six of the virtual components described in Proposition 14 are associated with single real connected component each, while the four remaining virtual components are each associated with a pair of real connected components that are mapped into each other by the central symmetry in ${\mathbb{R}}^{2}$.
Pictures $(a)$, $(b)$, and $(c)$ of Fig. 2 are realized by the polynomials
$xy(x^{2}+y^{2})-t,$
$xy(x^{2}+y^{2})-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/2),$
and
$xy(x^{2}+y^{2}-t)-\alpha t^{2},\qquad\alpha\in(0,1/8)$
for any $t>0$. All of these polynomials are invariant under the action of the group $G_{1}$. The polynomials obtained from them by the involution (12) and representing other three connected components of the discriminant complement are also invariant under this action.
Picture $(d)$ (respectively, $(e)$) is realized by the polynomial
(16) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),$
$\alpha\in(0,0.003)$, with positive (respectively, negative) $t$. These two pictures are topologically equivalent. However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement, see [15].
Each polynomial (16) is invariant under the transformation $(x,y)\mapsto(y,x)$. The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the polynomial with the opposite value of parameter $t$. In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
Pictures $(f)$ and $(g)$ of Fig. 2 are realized by polynomials
(17) $xy\left((x-2t)^{2}+(y-2t)^{2}-t^{2}\right)-t^{4}=0$
with positive and negative values of $t$, respectively. Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$; the transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (17) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
The same considerations hold for components obtained from these by the involution (12). This concludes the proof of Theorem 13. $\Box$
2.3. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.1) acts on the space $\Phi_{2}$ of all polynomials (1) with A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way. All components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
In addition, the operation
(18) $f(x,y)\mapsto-f\left(\frac{x+y}{\sqrt{2}},\frac{x-y}{\sqrt{2}}\right)$
acts on the space $\Phi_{2}$ (and on the entire space of polynomials of degree four). This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $I$ of the invariant Ind to a polynomial with value 3I-3-I. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation. The cardinalities of all these orbits are divisors of 8.
----++$(a)$
----++$(b)$
----++$(c)$
----++++$(d)$
----++++$(e)$
----++++$(f)$
----++++$(g)$
----++++$(h)$
++++++--$(i)$
Figure 3. Non-discriminant perturbations of $X_{9}^{2}$
Proposition 17 (see [12]).
There are exactly 18 virtual components of class $\Phi_{2}$. The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig. 3 (see also Fig. 33 in [18]), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((18)).
Theorem 18.
There are exactly 52 connected components of the discriminant complement in the space $\Phi_{2}$.
Specifically, the virtual components represented by pictures $(a)$ and $(b)$ of Fig. 3 are associated with single real connected components of the discriminant complement; the virtual components represented by pictures $(c)$ and $(i)$ are associated with pairs of such components ((whose pictures are obtained from each other by rotating the plane by the angle $\frac{\pi}{2})$; each of the virtual components represented by pictures $(d)$, $(e)$, $(f)$, $(g)$, and $(h)$ is associated with four real connected components of the discriminant complement, whose pictures are obtained from each other by rotating the plane by the angles $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$. The same is true for the components obtained from these nine by the operation ((18)).
Remark 19.
Picture ($c$) and its rotation by the angle $\frac{\pi}{2}$ are topologically equivalent, but the connected components of the discriminant complement realized by polynomials with these shapes are different.
Theorem 18 was conjectured in [12], and the difference of all the connected components mentioned there was proven. It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated by this theorem.
Proposition 20.
If two generic polynomials $f$ and $\tilde{f}$ of class $\Phi_{2}$ are associated with the same virtual function, then their connected components of the space of generic polynomials of the form ((1)) are mapped to each other by an element of the group $G_{0}$.
Lemma 21.
There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$. These eight polynomials form an orbit of the natural action of the group $G_{0}$.
Lemma 22.
Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by a linear transformation to the form
(19) $\pm\left[t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right]\ ,\hskip 23.49976ptt>0\ .$
In particular, the j-invariant of its principal homogeneous part is equal to 5.
Proof of Lemma 22. This lemma follows from Proposition 22 of [15] by the substitution $\tilde{x}=|a|^{\frac{-3}{4}}X$, $\tilde{y}=sign(a)|a|^{\frac{1}{4}}Y$, $t=|a|^{\frac{-5}{4}}$. $\Box$
Corollary 23.
If a polynomial of degree four in ${\mathbb{R}}^{2}$ has a critical point of class $D_{6}$ with zero critical value and a point of class $A_{3}$ with critical value $\frac{-1}{324}$, then it has the form
(20) $\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}$
in some affine coordinates.
Proof. Of all the polynomials (19), only this one has critical value $\frac{-1}{324}$. $\Box$
Corollary 24.
If a polynomial $f(x,y)$ of degree four has critical points of classes $D_{6}$ and $A_{3}$ and principal homogeneous part
(21) $x^{4}+2Ax^{2}y^{2}+y^{4}\ ,$
then $A$ is a root of the equation
(22) $\hskip 136.5733pt\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.\hskip 136.5733pt\Box$
[…]
The rest of the proof of Proposition 20 repeats the proof of Proposition 3 of [15] in § 5.4 of [15]. $\Box$
Proof of Theorem 18. Pictures ($a$) and ($b$) of Fig. 3 are realized, respectively, by the polynomials (23) and (24) with arbitrary $t>0$.
(23) $x^{4}-6x^{2}y^{2}+y^{4}+t$
(24) $x^{4}-6x^{2}y^{2}+y^{4}-2t(x^{2}+y^{2})+\frac{1}{2}t^{2}$
These polynomials are invariant under the entire group $G_{0}$. Therefore, their virtual components are associated with single real connected components of the discriminant complement.
Pictures ($c$) and ($i$) are realized respectively by polynomials (25) and (26) with arbitrary $t>0$.
(25) $(x^{2}-3t)^{2}-6(x^{2}-3t)(y^{2}-t/2)+(y^{2}-t/2)^{2}+t^{4}$
(26) $x^{4}-6x^{2}(y^{2}+t)+(y^{2}+t)^{2}-t^{2}/2.$
These polynomials are invariant under the reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y),$ which generate a subgroup of index two in $G_{0}$. Thus, in each of these two cases there are no more than two connected components of the discriminant complement associated with the corresponding virtual component.
Pictures ($d$), ($e$), ($f$), ($g$), and ($h$) are realized by the equations (27), (28), (29), (30), and (31) for sufficiently small $\varepsilon>0$ and arbitrary $t>0$.
(27) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}+\varepsilon t^{3}x\right)$
(28) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-\varepsilon t^{3}x\right)$
(29) ${\bf T}\left(x^{4}-6x^{2}y^{2}+t(x^{3}-\varepsilon^{4}t^{2}x)+(y^{2}-\varepsilon^{3}t^{2})^{2}\right)$
(30) ${\bf T}\left(x^{4}-6x^{2}y^{2}+y^{4}+tx^{3}-6\varepsilon txy^{2}-2\varepsilon^{4}t^{3}x+8\varepsilon^{3}t^{2}y^{2}\right)$
(31) ${\bf T}\left((x+0.7t)(x+1.3t)(x+6t)(x+7t)-6x^{2}y^{2}+y^{4}\right)$
These five polynomials are all constructed as follows. First we take the polynomial
$x^{4}-6x^{2}y^{2}+y^{4}+x^{3},$
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at the point closest to the center of the picture. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [18]. All the resulting polynomials are invariant under the reflection in the axis $y=0$. Thus, in each case, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons, see [12].
For the remaining nine virtual components, the realizations of the corresponding real connected components that prove Theorem 18 in these cases can be obtained from the polynomials (23)–(31) via the operation (18). $\Box$
3. Components for $J_{10}$
3.1.
The canonical versal deformations of the singularities of the classes $J_{10}^{1}$ and $J_{10}^{3}$ consist, respectively, of polynomials
(32) $(x-Ay^{2})(x^{2}+y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad A\in(-\infty,+\infty),$
and
(33) $(x-Ay^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad A\in(-1,1),$
with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, see [5].
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (32) and (33), respectively.
Proposition 25 (see [17]).
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials are either the same or mapped to each other by the involution
(34) $f(x,y)\leftrightarrow f(x,-y)\ .$
In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components. $\Box$
In the following pictures, the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
3.2. Components for $J_{10}^{1}$ singularities
++-$(a)$
++-$(b)$
++-$(c)$
++-$(d)$
++-$(e)$
++-$(f)$
++-($\tilde{b}$)
++-($\tilde{c}$)
++-($\tilde{d}$)
++-($\tilde{e}$)
Figure 4. Non-discriminant perturbations of $J_{10}^{1}$ singularities
Proposition 26 (see §4.2 in [18]).
There are exactly ten topological types of zero-level sets of non-discriminant perturbations of $J_{10}^{1}$ singularities, see Fig. 4. $\Box$
Proposition 27 (see [12]).
There are exactly ten virtual components of the formal graph of $J_{10}^{1}$ singularities. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
Theorem 28.
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types shown in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((34)).
Proof of Theorem 28. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
$x(x^{2}+y^{4}+t),$
$(x-2t)((x+t)^{2}+y^{4}-t^{2}),$
$(x-2t)((x+t)^{2}+y^{4}-3ty^{2}+2t^{2}),$
$(x-y^{2}+2t)(x^{2}-2tx+y^{4}-5t^{2})-t^{3}/10$
with arbitrary $t>0$. These polynomials are all invariant under the involution (34). Therefore, the corresponding virtual components are associated with single connected components of the discriminant complement.
The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from the realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
(35) $f(x,y)\mapsto-f(-x,y).$
All of them are also invariant under the involution (34).
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (34). Indeed, the polynomials (32) are of degree three in the variable $x$. Therefore, any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times.
An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows. Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take three arbitrary points inside the rightmost three ovals. There is a single parabola of the form $x=ay^{2}+by+c$ that passes through these three points. The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$. This restriction has six roots, so it is of degree exactly six. This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate. The position of the parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement. However, these relative positions are different for any polynomial with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y). The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
3.3. Components for $J_{10}^{3}$ singularities
---++++++$(a)$
---++++++$(b)$
---++++++$(c)$
---++++++$(d)$
---++++++$(e)$
---++++++$(f)$
---++++++$(g)$
---++++++$(h)$
---++++++$(i)$
---++++++$(j)$
---++++++$(k)$
---++++++$(l)$
---++++++$(m)$
Figure 5. Perturbations of $J_{10}^{3}$
Proposition 29 (see [12]).
There are exactly 23 virtual components of class $\Omega_{3}$. The shapes of the zero-level sets of polynomials $f(x,y)$ representing some thirteen of them are shown in Fig. 5 (see also Fig. 29 in [18]). In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((35)), represent different virtual components.
We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),…by the involution induced by the transformation (35), that is, by the reflection in a horizontal line and simultaneous change of all signs.
Remark 30.
Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however it follows from the Bezout’s theorem that they represent different connected components.
Theorem 31.
There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((33)). Each virtual component represented by a picture of Fig. 5 in the list
(36) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
[…]
Proof. The difference of all these 33 connected components easily follows from the topological considerations and Bezout’s theorem, see [12]. By Proposition 25, it remains to be shown that, in all cases not covered by the list (36), the involution (34) does not change the real connected component.
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$. All of these polynomials are even with respect to the variable $y$, so their virtual components are represented by single connected components of the complements of the discriminants.
Picture $(m)$ is represented by the polynomials
$x(x-y^{2}-t)(x+y^{2}+t),\qquad t>0,$
which are also self-symmetric.
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
realize picture $(c)$ for any $C\in(0,1/2))$. They are also invariant under the involution (34).
The polynomials
$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$
realize picture $(f)$.
Figure 6.
To realize the picture $(e)$, we start with the polynomial
(37) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)$
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal quasihomogeneous part will be of class $J_{10}^{3}$. The polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The topological picture of the zero-level set of the resulting polynomial will be the desired one. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see Table 2. Finally, we apply the operator $\bf T$} to reduce the polynomial to the form (1). The resulting polynomial realizes the picture $(e)$, is of class $\Omega_{3}$, and is invariant under the involution (34). $\Box$
4. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at the $X_{9}$ and $J_{10}$ singularities proves all of the conjectures in [11] concerning the number of local Petrovskii lacunas at the corresponding singular points of wavefronts.
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity when the dimension $N$ and the index $i_{+}$ are both even is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The four lacunas at the $X_{9}^{2}$ singularity, with even $N$ and odd $i_{+}$, are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (18) to these two. The unique lacuna at the $J_{10}^{3}$ singularity, with even $N$ and odd $i_{+}$, is shown in picture $(m)$ of Fig. 5.

### v2

@@
 Abstract.
-We provide the complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$ and $J_{10}$.
-Thus, we prove (and improve in one particular case) all corresponding conjectures from the previous work [12] with the same title.
+We provide a complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$, $J_{10}$ and $P_{8}^{1}$ (as well as a partial list for the remaining class, $P_{8}^{2}$).
+Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [18] with the same title.
+As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs.
+In particular, we discover a new local lacuna at the $P_{8}^{2}$ singularities.
+We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, unlike all simple singularities.
 Key words and phrases:
@@
 1. Introduction
+$1.$
 Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular.
-These varieties appear in many problems of PDE theory and integral geometry (often under the name of wavefronts), see e.g.
-[1], [2], [3], [10], [11].
-Many important functions in these theories are defined by integral transforms and are regular outside the discriminant varieties of deformations of corresponding generating functions.
-Their analytic behavior near the singular points of the discriminants depends on the singularity type of these functions and on the connected component of the complement of the discriminant.
-Thus, enumerating these components is one of the first steps in the study of their analytical properties.
-The list of function singularities begins with simple singularities, which were distinguished by V.
+These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g.
+[1], [2], [3], [8], [15][16], [17].
+Many important functions in these theories arise as integral transforms.
+These functions are defined on the parameter spaces of certain function deformations (called generating families) and are equal to the integrals along the cycles in the zero-level sets of the corresponding functions.
+These integrals are regular outside the discriminant varieties of the deformations.
+Their analytic behavior near the discriminants depends on the singularity type of the deformed discriminant function and on the connected component of the complement of the discriminant.
+Thus, enumerating these components is one of the first steps in studying the analytical properties of these integral transforms.
+Definition 1 (see [2]).
+A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},0)\to({\mathbb{R}},0)$ at its critical point.
+Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$.
+Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
+$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
+An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin in ${\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic.
+The discriminant of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value (or, equivalently, the zero-level set $f_{\lambda}^{-1}(0)$ is singular near real space).
+For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [2]) deformation.
+The discriminant sets of all versal deformations of stably equivalent singularities are topologically equivalent, and the components of their complements are in a one-to-one correspondence.
+Therefore, for any stable equivalence class of singularities it is sufficient to enumerate the components for only one versal deformation of one function of this class.
+The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V.
 Arnold, see [2].
-E.
-Looijenga [8] implicitly listed the local connected components of complements of their discriminants.
-O.
-Viro [18] listed all possible topological types of non-discriminant functions representing these components.
-(By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood.
-The topological type is an invariant of the components of the discriminant complements.) In [13], it was proven that each topological type of this list represents exactly one connected component of the discriminant complement.
-The next important class of function singularities are the parabolic ones, which were also distinguished by V.
-Arnold.
-The classification of complex parabolic singularities of functions consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$; the classification of their real forms has more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
-In [12], a system of invariants of the connected components of the complements of the discriminant varieties of real function singularities was described.
-The complete lists of the possible values of these invariants for parabolic singularities were found, and all of these values were realized by concrete polynomials.
-We also conjectured that the resulting lists of connected components thus found in [12] are complete.
-These lists consist of 7, 14, 52, 11, and 33 components for the singularities of classes $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively.
-Below, we prove that almost all of these lists are indeed complete.
-However, some two mutually symmetric sets of perturbations of the $J_{10}^{1}$ singularities indicated in [12] consist of pairs of different connected components.
-Thus, the number 11 above is replaced by 13.
-These results are summarized in Theorems 5, 13, 18, 28, and 31 below.
-The classes $P_{8}^{*}$ are not considered yet.
-After the publication of [12] and [13], I discovered that O.
-Viro [18] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities (as well as of simple singularities).
-Table 1. Numbers of isotopy and topological classes of non-discriminant perturbations of $X_{9}$ and $J_{10}$ singularities
-Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$
-Topology 6 12 50 10 31
-Isotopy 7 14 52 13 33
-It turns out that the “isotopy” classification of parabolic singularity perturbations considered in the present work is only slightly stricter than the topological classification, see Table 1.
-E.
-Looijenga [8] proved that all local components of discriminant complements of simple singularities are contractible.
-Conversely, we show that at least some three such components of parabolic singularities of classes $X_{9}^{+}$ and $X_{9}^{-}$ have infinite one-dimensional homology groups, see Theorem 12 on page 12.
-An application to hyperbolic PDE’s.
-In the theory of hyperbolic PDE’s, the local Petrovskii lacunas are the local connected components of the complements of wavefronts, in which the wave functions behave regularly, see e.g.
-[1], [10].
-In [11], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete.
-The enumeration of all local components of wavefront complements given below proves this conjecture for all parabolic singularities of classes $X_{9}$ and $J_{10}$, see § 4.
-2. Results for $X_{9}^{*}$ classes
-Table 2. Normal forms of $X_{9}$ and $J_{10}$ singularities in two variables
+All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of functions in two variables.
+The components of discriminant complements for these singularities were listed earlier, see § 1.3.
+Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and $P_{8}$ singularities in three variables
 Notation Normal form Restriction
-$X_{9}^{+}$ $x^{4}+2Ax^{2}y^{2}+y^{4}$ $|A|<1$
+$A_{2k-1}$ $\pm x^{2k}\pm y^{2}$ $k\geq 1$
+$A_{2k}$ $x^{2k+1}\pm y^{2}$ $k\geq 1$
+$D_{2k}^{\pm}$ $x^{2}y\/\pm y^{2k-1}$ $k\geq 2$
+$D_{2k-1}$ $\pm(x^{2}y\/+y^{2k-2})$ $k\geq 3$
+$E_{6}^{\pm}$ $x^{3}\pm y^{4}$
+$E_{7}$ $x^{3}+xy^{3}$
+$E_{8}$ $x^{3}+y^{5}$
+$X_{9}^{\pm+}$ $\pm(x^{4}+2Ax^{2}y^{2}+y^{4})$ $|A|<1$
 $X_{9}^{1}$ $xy(x^{2}+2Axy+y^{2})$ $|A|<1$
@@
 $J_{10}^{3}$ $(x^{2}-y^{2})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
-Definition 1.
-A function singularity is a germ of a $C^{\infty}$-smooth function ${\mathbb{R}}^{2}\to{\mathbb{R}}$ at its critical point.
-Real parabolic function singularities of type $X_{9}$ form four families $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
-Any function singularity of classes $X_{9}^{+}$, $X_{9}^{1}$ or $X_{9}^{2}$ has the normal form indicated in Table 2 in appropriate local coordinates centered at the critical point.
-The class $X_{9}^{-}$ (and all considerations concerning it) can be reduced to the case $X_{9}^{+}$ by multiplying by 1-1.
-The standard miniversal deformations of functions of all three classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$ have the same form: they are families of polynomials
+$P_{8}^{1}$ $x^{3}+y^{3}+z^{3}-3Axyz$ $A<1$
+$P_{8}^{2}$ $x^{3}+y^{3}+z^{3}-3Axyz$ $A>1$
+In this work we list analogous components for the next natural class in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V.
+Arnold.
+The classification of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$.
+The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
+Functions of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables.
+Any function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any function in three variables of one of classes $P_{8}^{*}$ can be reduced to a normal form indicated in Table 1 by an appropriate choice of local coordinates centered at the critical point.
+Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
+We give complete lists of components for all parabolic singularity classes except $P_{8}^{2}$, for which we have only partial results.
+Table 2. Numbers of rigid isotopy and topological classes of non-discriminant perturbations of parabolic singularities
+Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$ $P_{8}^{1}$ $P_{8}^{2}$
+Isotopy 7 14 52 13 33 7 $\geq 15$
+Topology 6 12 50 10 31
+O.
+Viro [24] found all topological types of non-discriminant perturbations of the $X_{9}$ and $J_{10}$ singularities.
+By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood.
+The topological type is an invariant of the components of the discriminant complements.
+The “rigid isotopy” classification of parabolic singularity perturbations given in the present work is only slightly stricter than the topological classification, see Table 2.
+Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [7].
+This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
+1.2. On the method
+In [18], a system of invariants of connected components of the complements of the discriminant varieties of function singularities is described.
+Complete lists of possible values of these invariants for parabolic singularities are given in [18], and all of these values are realized by concrete polynomials.
+In the same paper, we also conjectured that the resulting lists of connected components are complete.
+These lists consist of 7, 15, 7, 14, 52, 11, and 33 components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively.
+Below, we prove that almost all of these lists are indeed complete.
+However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [18] consist of pairs of different connected components.
+Also, we present below partial results for the case $P_{8}^{2}$, which prove that at least 10 out of 15 classes listed in [18] are indeed represented by only one connected component each.
+Thus, the number 11 above is replaced in Table 2 by 13, and the number 15 is still under suspicion.
+These results are summarized in Theorems 5, 13, 18, 28, 31, 37, and 41 below.
+1.3. On the discriminant complements for simple singularities
+E.
+Looijenga [13] implicitly listed the local connected components of the complements of the discriminants of simple singularities.
+O.
+Viro [24] listed all possible topological types of non-discriminant functions representing these components.
+In [19], it was shown that each topological type on this list corresponds to exactly one connected component of the discriminant complement.
+For simple singularities with Milnor numbers at most six this fact also follows from the results of [14].
+Unfortunately, I only read the work [24] containing the enumeration of topological types of simple and parabolic function perturbations in two variables after the publication of [18] and [19].
+1.4. 1-cohomology of discriminant complements
+E.
+Looijenga [13] proved that all local components of discriminant complements of real simple singularities are topologically trivial.
+Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 12 and 39.
+1.5. An application to hyperbolic PDEs.
+In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [1], [8], [16]).
+Near the non-singular points of wavefronts (corresponding to the Morse functions) the lacunas were enumerated in [6] and [5].
+Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [8], and near all simple singularities in [15].
+In [17], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete.
+We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
+For the singularities of classes $P_{8}^{1}$, $X_{9}^{*}$ and $J_{10}^{*}$, these results follow directly from our complete lists of components of discriminant complements.
+Our partial results on enumerating components for the $P_{8}^{2}$ singularities are sufficient to prove that our new list of local lacunas is also complete in that case.
+2. Results for the $X_{9}^{*}$ classes
+2.1.
+Real parabolic function singularities of type $X_{9}$ form four stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
+The standard versal deformations of functions in all these classes have the same form: they are families of polynomials
 (1) $f_{A}(x)+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}x^{2}+\lambda_{5}xy+\lambda_{6}y^{2}+\lambda_{7}x^{2}y+\lambda_{8}xy^{2},$
 depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 1.
-By the versality theory (see, e.g., [2]), the study of all other deformations of these singularities can be reduced to these.
+According to versality theory (see, for example, [2]), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families.
 Definition 2.
 A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (1) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ consists of four different lines in ${\mathbb{C}}^{2}$.
-The spaces $\Phi_{+}$, $\Phi_{1}$, and $\Phi_{2}$ consist of all polynomials of the form (1) with principal parts of classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$, respectively.
+Denote by $\Phi_{+}$, $\Phi_{1}$, and $\Phi_{2}$ the spaces of all polynomials of the form (1) with principal parts of classes $X_{9}^{+}$, $X_{9}^{1}$, and $X_{9}^{2}$, respectively.
 The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$.
@@
 These varieties divide the spaces $\Phi_{*}$ into connected components.
-The enumeration of these components is our goal in this section.
-In §2 of [12] we introduced the set-valued invariant of these components. Let us briefly review this construction.
-The virtual function associated with a generic real polynomial is a collection of certain topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set.
-Namely, these data of a function $f(x,y)$ consist of
-  • the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented basic vanishing cycles in the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
-  • the string of intersection indices of these cycles with the set of real points,
+Enumerating of these components is our goal in this section.
+2.2. Virtual functions and virtual components
+In §2 of [18] we introduced the set-valued invariant of connected components of the discriminant complements. Let us briefly review this construction in the case of $X_{9}$ singularities.
+The virtual function associated with a generic real polynomial is a collection of topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set.
+Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
+  • the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
+  • the string of intersection indices of these nine cycles with the set of real points,
   • the Morse indices of real critical points of $f$, and
@@
 The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions.
-The formal graph of a singularity class $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips.
-The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries crossing the discriminant are removed.
-The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component $S(f)$ of the formal graph containing the virtual functions associated with this polynomial.
+The formal graph of any type $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips.
+The virtual components of the formal graph are the connected subgraphs, into which the formal graph splits when all edges modeling the surgeries that cross the discriminant are removed.
+The set-valued invariant of a generic polynomial $f$ of class $\Phi_{*}$ is the virtual component of the formal graph containing the virtual functions associated with this polynomial.
+This virtual component is denoted by $S(f)$.
 These objects have the following obvious property.
 Proposition 1.
-If two polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same. $\Box$
-Enumerating the virtual components of a finite formal graph is a combinatorial problem that can be solved by a computer program.
-The work [12] is essentially a list of the results of these computations for all parabolic singularities and of their immediate consequences in the terms of real functions.
-The remaining problem is to find, for each virtual component, all real connected components with this value of the set-valued invariant.
-It follows from the results of [5] that, for parabolic singularities, all standard flips of virtual functions can be realized by real surgeries of generic real functions.
+If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same.
+$\Box$
+The formal graphs of all simple and parabolic singularities are finite.
+Enumerating the virtual components of such a graph is a combinatorial problem that can be solved using a computer program.
+The work [18] provides the results of these computations for all parabolic singularities, as well as the immediate consequences of these computations in the terms of real functions.
+The remaining problem is to find all real connected components with this value of the set-valued invariant for each virtual component.
+The results of [10] imply that in the case of parabolic singularities all standard flips of virtual functions can be realized by real surgeries of generic real functions.
 In particular, the following statement holds.
@@
 For any class of polynomials $\Phi_{+}$, $\Phi_{1}$, or $\Phi_{2}$,
-a)) every virtual function of this class is associated with a generic real polynomial $\tilde{f}$ of the same class;
-b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the complement of the discriminant. $\Box$
+a)) every virtual function of this class is associated with a generic real polynomial of the same class;
+b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$.
+$\Box$
 Remark 3.
-In the statement of Proposition 3 of [15], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four polynomials with non-degenerate principal homogeneous parts.
-However, the entire construction proving this proposition can be performed within the spaces of the versal deformations (1).
-Thus, it also proves Proposition 2.
-These considerations enable us to estimate the number of isotopy classes of generic Morse perturbations associated with any given virtual function.
+In the statement of Proposition 3 of [21], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with nondegenerate principal homogeneous parts.
+However, the entire construction proving this proposition in [21] can be performed within the spaces of the versal deformations (1).
+Thus, it also proves the above Proposition 2.
+These constructions enable us to estimate from above the number of isotopy classes of generic Morse perturbations associated with any given virtual function.
 These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately.
@@
 In the case $X_{9}^{1}$ the analogous group is of order four.
-Then, for each virtual component in the list from [12], we compare all real components represented by the polynomials associated with some its virtual function and forming an orbit of this group, and determine which of these components are identical.
+Then, for each virtual component in the list from [18], we compare all real components associated with this virtual component and determine which of these components are identical.
 This comparison provides the results of the present work concerning the $X_{9}$ singularities.
-The strategy for the $J_{10}$ singularities is exactly the same.
+The strategy for the $J_{10}$ and $P_{8}$ singularities is exactly the same.
 Remark 4.
@@
 Definition 3.
-For any generic polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$, $\mbox{Ind}(f)$ is the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
-It is easy to see that this number is an invariant of the connected components of the discriminant complements.
-2.1. Computations for the $X_{9}^{+}$ singularities
+For any Morse polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
+It is easy to see that this number is an invariant of the connected components of the discriminant complements (and of associated virtual components).
+2.3. Enumeration of components for $X_{9}^{+}$ singularities
 In this subsection, we prove the following statement, conjectured in [18].
@@
 However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$.
-Figure 1 coincides with the well-known list of affine isotopy classes of bounded degree four algebraic plane curves.
-In our case, however, the isotopies are allowed only in a much smaller space $\Phi_{+}$.
+Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves.
+In our case, however, the isotopies are allowed only within a much smaller space $\Phi_{+}$.
 In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
 Clearly, this group is isomorphic to the group of symmetries of the square, in particular is of order eight.
-This group acts on the space $\Phi_{+}$ of polynomials (1) with principal parts of class $X_{9}^{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$).
+This group acts on the space $\Phi_{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$).
 This action preserves the set of generic polynomials.
@@
 Lemma 9 (see [21], Lemma 4).
-A function of the form ((1)) with a positive principal homogeneous part has a critical point of class $E_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{27}{256}$ if and only if it is one of the four polynomials
-(2) $\hskip 82.51282pt{\bf T}(x^{4}\pm x^{3}+y^{4})\qquad\mbox{ or }\qquad{\bf T}(x^{4}+y^{4}\pm y^{3})\ .\hskip 82.51282pt\Box$
+A polynomial of the form ((1)) with a positive principal homogeneous part has a critical point of class $E_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{27}{256}$ if and only if it is one of the four polynomials
+(2) $\hskip 76.82234pt{\bf T}(x^{4}\pm x^{3}+y^{4})\quad\mbox{ or }\quad{\bf T}(x^{4}+y^{4}\pm y^{3})\ .\hskip 76.82234pt\Box$
 Proof of Proposition 8.
@@
 In the next paragraph, we present the outline of this proof; we refer to [21] for all common details.
-Using the Lyashko–Looijenga covering (in the form described in [5]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the associated virtual function.
-Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the discriminant and the set of not strictly Morse functions).
+Using the Lyashko–Looijenga covering (in the form described in [10]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function.
+Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions).
 Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$.
-In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
+(For this existence statement, the parabolicity of $X+9$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
 First, suppose that it is the same point $I(1)$.
-A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that commutes with the Lyashko–Looijenga map: the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$.
-Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this reflection.
-In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this reflection, which belongs to the group $G_{0}$.
+A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map: this is the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$.
+Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this automorphism.
+In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$.
 If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$).
-This leads to the previous situation.
+This leads us to the situation considered previously.
 $\Box$
+Proof of Theorem 5.
 By Proposition 8, any virtual component mentioned in Proposition 7 is associated with at most eight real connected components.
@@
 1 is associated with a single real connected component.
-Pictures $(a)$, $(b)$, $(c)$, $(f)$ of Fig. 1 are realized, respectively, by four polynomials
+Pictures $(a)$, $(c)$, $(b)$, $(f)$ of Fig. 1 are realized, respectively, by polynomials
 $x^{4}+y^{4}+t,$
-$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\hskip 23.49976pt\alpha\in(0,1/4),$
 $x^{4}+y^{4}-t,$
-$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(1/4,1/2),$
+$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/4),$
+$x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\quad\alpha\in(1/4,1/2),$
 with arbitrary $t>0$. All of these polynomials are invariant under the entire group $G_{0}$.
@@
 with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
-(4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2$
+(4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2\ .$
 The polynomials (3) and (4) are connected by the family of polynomials
 (5) $x^{4}+2Ax^{2}y^{2}+y^{4}+2(x\cos\varphi+y\sin\varphi)^{2}-2(x\sin\varphi-y\cos\varphi)^{2}+1/2,$
-depending on the parameter $\varphi\in[0,\pi/2]$.
+that depend on the parameter $\varphi\in[0,\pi/2]$.
 Lemma 10.
-If $A<1$ is sufficiently close to $1$, then all polynomials of this family are non-discriminant, in particular their zero-level sets are all smooth and isotopic to each other.
+[…]
 Proof.
@@
 Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle.
-If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $x^{2}+y^{2}\leq 2$ and maintains its topological structure in this domain for any $\varphi$.
+If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$.
 It also remains empty outside this domain.
@@
 The topological type of the curve shown in picture $(e)$ can be realized by any of the four polynomials
-(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa)$
-and
+(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa),$
 (7) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(y^{3}-3x^{2}y)+\varkappa)$
@@
 These four polynomials form an orbit of the group $G_{0}$.
-All of them belong to the one-parameter family of polynomials
+All of them belong to the one-parametric family of polynomials
 (8) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}+\mbox{Re}(e^{i\tau}(x+iy)^{3}))+\varkappa),$
@@
 For $A=1$, all these polynomials are obtained from each other by the rotations of ${\mathbb{R}}^{2}$.
-They are bounded from below, have a critical point of class $D_{4}^{-}$ with zero critical value at the origin and three real minimum points with critical value $\frac{27}{256}+\varkappa$, and no other critical points.
-Therefore, for $\varkappa<\frac{27}{256}$ all of them realize the topological picture $(e)$.
-Analogously to Lemma 10, taking $A<1$ very close to 1, we will not change these topological pictures for any $\tau$.
+They are bounded from below, have a critical point of class $D_{4}^{-}$ with critical value $\varkappa$ at the origin, three real minimum points with critical value $\frac{27}{256}+\varkappa$, and no other critical points.
+Therefore, for $\varkappa\in\left(0,\frac{27}{256}\right)$ all these polynomials realize the topological picture $(e)$.
+Analogously to the proof of Lemma 10, taking $A<1$ very close to 1 will not change these topological pictures for any $\tau$.
 So, the connected component of the discriminant complement containing these polynomials is invariant under the action of the group $G_{0}$.
 The picture $(g)$ can be realized by either of the four polynomials
-(9) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10x(x^{2}-3y^{2})+15(x^{2}+y^{2})-5)$
-and
-(10) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10y(y^{2}-3x^{2})+15(x^{2}+y^{2})-5).$
-with sufficiently small $\varepsilon>0$. These four polynomials also form an orbit of the group $G_{0}$.
+(9) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10x(x^{2}-3y^{2})+15(x^{2}+y^{2})-5)\ ,$
+(10) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}\pm 10y(y^{2}-3x^{2})+15(x^{2}+y^{2})-5)$
+with sufficiently small $\varepsilon>0$. These four polynomials form an orbit of the group $G_{0}$.
 The family of polynomials
-(11) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}+10\mbox{Re}\left(e^{i\tau}(x+iy))^{3}\right)+15(x^{2}+y^{2})-5),$
+(11) ${\bf T}(x^{4}+(2-\varepsilon)x^{2}y^{2}+y^{4}+10\mbox{Re}\left(e^{i\tau}(x+iy)^{3}\right)+15(x^{2}+y^{2})-5),$
 $\tau\in[0,2\pi],$ contains all four polynomials (9), (10) for $\tau=0,$ $\frac{\pi}{2}$, $\pi$, and $\frac{3\pi}{2}$.
@@
 To realize the pictures ($d$), ($e$), and ($g$) by arbitrarily small perturbations of the $X_{9}^{+}$ singularities, we can multiply by $t^{4-k}$ all monomials of degree $k$ in the formulas (3), (6), and (9).
-The 1-cohomology of discriminant complements.
+2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
 Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane.
 For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$.
-Let $C$ be a connected component of the discriminant complement in the space $\Phi_{+}$ consisting of polynomials with $k\geq 2$ non-nested ovals.
-Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(C)$ is well-defined as follows.
-For any such polynomial $f$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals.
-These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $C$.
-The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $C\to B({\mathbb{R}}^{2},k)$.
+The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
+Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are $k\geq 2$ non-nested ovals.
+Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows.
+For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals.
+These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$.
+The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$.
 Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
-Define the class $W\in H^{1}(C)$ as the class obtained by this map from the generator of the group $H^{1}(B({\mathbb{R}}^{2},k))$.
+Define the class $W\in H^{1}(U)$ as the class obtained by this map from the generator of the group $H^{1}(B({\mathbb{R}}^{2},k))$.
 Theorem 12 (cf. [20], § 1.8.3).
-The class $W$ takes all integer values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
+The class $W$ takes unit values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
 Proof. These cycles are the families (5), (8), and (11). $\Box$
@@
 Clearly, this group is isomorphic to ${\mathbb{Z}}_{2}^{2}$.
-It acts on the space of polynomials $\Phi_{1}$ and preserves the set of generic polynomials.
-Polynomials obtained from each other by this action have equal associated virtual functions and so are not separated by the set-valued invariant.
+It acts on the space $\Phi_{1}$ of polynomials (1) with principal part
+(12) $xy(x^{2}+2Axy+y^{2}),\quad|A|<1,$
+and preserves the set of generic polynomials. Polynomials obtained from each other by this action have equal associated virtual functions and so are not separated by the set-valued invariant.
 Additionally, the involution
 (13) $f(x,y)\mapsto-f(x,-y)$
-acts on the space of polynomials (1) of class $\Phi_{1}$.
-It preserves the set of discriminant polynomials and hence permutes the connected components of its complement.
-This permutation has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components related by this involution is always 1-1.
-This involution naturally extends to the set of associated virtual components: each virtual component associated with the polynomials from the connected component containing the polynomial $f(x,y)$ goes to a virtual component associated with the polynomials isotopic to f(x,y)-f(x,-y).
-The resulting involution on the set of virtual components of class $\Phi_{1}$ also has no fixed elements.
-For any polynomial $f$ with a principal homogeneous part of the form
-(13) $xy(x^{2}+2Axy+y^{2}),\qquad|A|<1,$
-its zero-level set goes to the infinity along four branches: two vertical and two horizontal.
+acts on the space $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement.
+This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y).
+These involutions have no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
+For any polynomial $f$ with a principal homogeneous part of the form (12), its zero-level set goes to the infinity along four branches: two vertical and two horizontal.
 If the polynomial is non-discriminant, then each vertical branch is matched with the horizontal branch belonging to the same component of the zero-level set.
@@
 Definition 6.
-For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with principal homogeneous part (13), denote by ${\bf\tilde{T}}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of the plane.
+For any polynomial $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of degree four with the principal homogeneous part (12), denote by ${\bf\tilde{T}}(f)$ the unique polynomial of the form (1) with the same principal part, which is obtained from $f$ by a parallel translation of the plane.
 (This is the translation by the vector $(\alpha,\beta),$ where $\alpha$ and $\beta$ are the coefficients of $f$ at the monomials $y^{3}$ and $x^{3}$).
 Lemma 16.
-A polynomial of the form ((1)) with the principal part ((13)) has a critical point of class $E_{7}$ with zero critical value and a critical point of class $A_{2}$ with critical value $3^{\frac{15}{2}}2^{-8}$ if and only if it is one of the following four polynomials:
-(14) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm x^{3})$
-or
-(15) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm y^{3})\ .$
+A polynomial of the form ((1)) with the principal part ((12)) has a critical point of class $E_{7}$ with zero critical value and a critical point of class $A_{2}$ with critical value $3^{\frac{15}{2}}2^{-8}$ if and only if it is one of four polynomials
+(14) ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm x^{3})\ \mbox{ \rm or }\ {\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}\pm y^{3})\ .$
 Proof.
@@
 By Lemma 5 of [21], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12).
-In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $ax^{3}$ or $ay^{3}$, $a\neq 0$.
+In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$.
 The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$.
@@
 Proof of Proposition 15.
-The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with one simplification.
+The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with the following simplification.
 In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [21], § 4.5).
-Therefore, the statement follows immediately from the fact that all points (14) and (15) belong to the same orbit of the group $G_{1}$.
+Therefore, the statement of Proposition 15 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$.
 $\Box$
 Proof of Theorem 13.
-Proposition 15 implies that any virtual component of the formal graph of the $X_{9}^{1}$ singularity is associated with at most four real connected components mapped to each other by the group $G_{1}$.
-Below we prove that six of the virtual components described in Proposition 14 are associated with single real connected component each, while the four remaining virtual components are each associated with a pair of real connected components that are mapped into each other by the central symmetry in ${\mathbb{R}}^{2}$.
+Proposition 15 implies that any virtual component of the formal graph of type $X_{9}^{1}$ is associated with at most four real connected components mapped to each other by the group $G_{1}$.
+Below we prove that six of the virtual components described in Proposition 14 are associated with single real connected component each, while each of the four remaining virtual components is associated with a pair of real connected components that are mapped into each other by the central symmetry in ${\mathbb{R}}^{2}$.
 Pictures $(a)$, $(b)$, and $(c)$ of Fig. 2 are realized by the polynomials
 $xy(x^{2}+y^{2})-t,$
-$xy(x^{2}+y^{2})-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/2),$
+$xy(x^{2}+y^{2})-t(x^{2}+y^{2})+\alpha t^{2},\quad\alpha\in(0,1/2),$
 and
-$xy(x^{2}+y^{2}-t)-\alpha t^{2},\qquad\alpha\in(0,1/8)$
+$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8)$
 for any $t>0$.
@@
 These two pictures are topologically equivalent.
-However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement, see [15].
+However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement in the space $\Phi_{1},$ see [21].
 Each polynomial (15) is invariant under the transformation $(x,y)\mapsto(y,x)$.
-The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the polynomial with the opposite value of parameter $t$.
+The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the analogous polynomial with the opposite value of parameter $t$.
 In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
@@
 with positive and negative values of $t$, respectively.
-Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$; the transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (17) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
+Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$.
+The transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (16) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
 The same considerations hold for components obtained from these by the involution (13). This concludes the proof of Theorem 13. $\Box$
 2.6. Components for the $X_{9}^{2}$ singularities
-The group $G_{0}$ (see § 2.1) acts on the space $\Phi_{2}$ of all polynomials (1) with A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way.
-All components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
+The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way.
+All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
 In addition, the operation
 (17) $f(x,y)\mapsto-f\left(\frac{x+y}{\sqrt{2}},\frac{x-y}{\sqrt{2}}\right)$
-acts on the space $\Phi_{2}$ (and on the entire space of polynomials of degree four).
+acts on the space $\Phi_{2}$.
 This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$.
@@
 So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
-The cardinalities of all these orbits are divisors of 8.
 ----++$(a)$
@@
 The shapes of the zero-level sets of the polynomials representing nine of these virtual components are shown in Fig.
-3 (see also Fig.
-33 in [18]), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((18)).
+3 ((see also Fig.
+33 in [24])), and the polynomials representing the remaining nine virtual components are obtained from these by the operation ((17)).
 Theorem 18.
@@
 Theorem 18 was conjectured in [18], and the difference of all the connected components mentioned there was proven.
-It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated by this theorem.
+It remains to prove that there are no additional components, that is, the virtual components listed in Proposition 17 are associated with no more real components than indicated in this theorem.
 Proposition 20.
@@
 Lemma 22.
-Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by a linear transformation to the form
-(19) $\pm\left[t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right]\ ,\hskip 23.49976ptt>0\ .$
+Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
+(18) $\pm\left(t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right)\ ,\quad t>0\ .$
 In particular, the j-invariant of its principal homogeneous part is equal to 5.
@@
 in some affine coordinates.
-Proof. Of all the polynomials (19), only this one has critical value $\frac{-1}{324}$. $\Box$
+Proof. Of all the polynomials (18), only the polynomial (19) has critical value $\frac{-1}{324}$. $\Box$
 Corollary 24.
-If a polynomial $f(x,y)$ of degree four has critical points of classes $D_{6}$ and $A_{3}$ and principal homogeneous part
-(21) $x^{4}+2Ax^{2}y^{2}+y^{4}\ ,$
-then $A$ is a root of the equation
-(22) $\hskip 136.5733pt\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.\hskip 136.5733pt\Box$
+If a polynomial $f(x,y)$ of degree four with principal homogeneous part
+(20) $x^{4}+2Ax^{2}y^{2}+y^{4}$
+has critical points of classes $D_{6}$ and $A_{3}$, then $A$ is a root of the equation
+(21) $\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.$
+Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [10].
 […]
@@
 (24) $(x^{2}-3t)^{2}-6(x^{2}-3t)(y^{2}-t/2)+(y^{2}-t/2)^{2}+t^{4}$
-(26) $x^{4}-6x^{2}(y^{2}+t)+(y^{2}+t)^{2}-t^{2}/2.$
+(25) $x^{4}-6x^{2}(y^{2}+t)+(y^{2}+t)^{2}-t^{2}/2$
 These polynomials are invariant under the reflections $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y),$ which generate a subgroup of index two in $G_{0}$.
@@
 $x^{4}-6x^{2}y^{2}+y^{4}+x^{3},$
-whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at the point closest to the center of the picture.
+whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at its center.
 Then, we perturb this singular point by adding very small terms of lower degrees.
 These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [24].
-All the resulting polynomials are invariant under the reflection in the axis $y=0$.
-Thus, in each case, the virtual component is associated with no more than four real connected components of the discriminant complement.
-These four components differ for the topological reasons, see [12].
-For the remaining nine virtual components, the realizations of the corresponding real connected components that prove Theorem 18 in these cases can be obtained from the polynomials (23)–(31) via the operation (18).
-$\Box$
-3. Components for $J_{10}$
+All the resulting polynomials (2630) are invariant under the reflection in the axis $y=0$.
+Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement.
+These four components differ for the topological reasons.
+For the remaining nine virtual components, the convenient realizations of the corresponding real connected components can be obtained from the polynomials (22)–(30) via the operation (17). $\Box$
+3. Components for $J_{10}$ singularities
 3.1.
@@
 (32) $(x-Ay^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad A\in(-1,1),$
-with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, see [5].
+with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, cf. [10].
+Definition 7.
+A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (31) or (32) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal quasihomogeneous part of $f$ (of degree six with weights $\deg x=2,\deg y=1$) consists of three different complex parabolas or two parabolas and one line in ${\mathbb{C}}^{2}$.
 Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively.
+The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consists of 10 elements.
 Proposition 25 (see [23]).
-If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials are either the same or mapped to each other by the involution
+If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
 (33) $f(x,y)\leftrightarrow f(x,-y)\ .$
@@
 Proposition 27 (see [18]).
-There are exactly ten virtual components of the formal graph of $J_{10}^{1}$ singularities.
+There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$.
 These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig.
@@
 4 is associated with a single real connected component.
-Each of the virtual components/topological types shown in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((34)).
-Proof of Theorem 28. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
+Each of the virtual components/topological types represented in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((33)).
+Proof. The topological types shown in pictures ($a$), ($b$), ($c$), and ($d$) are represented respectively by the following polynomials:
 $x(x^{2}+y^{4}+t),$
@@
 Therefore, the corresponding virtual components are associated with single connected components of the discriminant complement.
-The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from the realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
+The realizations of pictures (b~)($\tilde${b}), (c~)($\tilde${c}), and (d~)($\tilde${d}) can be obtained from these realizations of pictures $(b),$ $(c)$, and $(d)$ by the involution
 (34) $f(x,y)\mapsto-f(-x,y).$
@@
 No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (33).
-Indeed, the polynomials (32) are of degree three in the variable $x$.
-Therefore, any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times.
+Indeed, the polynomials (31) are of degree three in the variable $x$, so any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times.
+Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other.
+On the other hand, the involution (34) moves the component of picture (f) to the same component as (13).
 An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows.
@@
 Let us take three arbitrary points inside the rightmost three ovals.
-There is a single parabola of the form $x=ay^{2}+by+c$ that passes through these three points.
+There is a single parabola (or line) of the form $x=ay^{2}+by+c$ that passes through these three points.
 The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$.
@@
 This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate.
-The position of the parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement.
-However, these relative positions are different for any polynomial with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y).
+The position of this parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement.
+However, these relative positions are different for any polynomial $f(x,y)$ with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y).
 The same is true for polynomials that realize the picture (e~)($\tilde${e}).
@@
 Proposition 29 (see [18]).
-There are exactly 23 virtual components of class $\Omega_{3}$.
-The shapes of the zero-level sets of polynomials $f(x,y)$ representing some thirteen of them are shown in Fig.
-5 (see also Fig.
-29 in [18]).
+There are exactly 23 virtual components of the formal graph of type $J_{10}^{3}$.
+The shapes of the zero-level sets of some polynomials $f(x,y)$ representing thirteen of these components are shown in Fig.
+5.
 In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components.
-We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),…by the involution induced by the transformation (35), that is, by the reflection in a horizontal line and simultaneous change of all signs.
+$\Box$
 Remark 30.
-Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however it follows from the Bezout’s theorem that they represent different connected components.
+Twelve topological types shown in these pictures are given in Fig. 29 of [24]. Pictures $(d)$ and $(e)$ of Fig. 5 are topologically equivalent, however they represent different virtual components.
+We will denote by (a~)($\tilde${a}), (b~)($\tilde${b}),…the pictures obtained from the corresponding pictures $(a)$, (b)(b),… of Fig.
+5 by the transformation (34), that is, by the reflection in a horizontal line and simultaneous change of all signs.
 Theorem 31.
-There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((33)). Each virtual component represented by a picture of Fig. 5 in the list
+There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((32)). Each virtual component represented by one of ten pictures
 (35) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
@@
 Proof.
-The difference of all these 33 connected components easily follows from the topological considerations and Bezout’s theorem, see [12].
+The difference of all these 33 connected components follows from the topological considerations and Bezout’s theorem, see [18].
 By Proposition 25, it remains to be shown that, in all cases not covered by the list (35), the involution (33) does not change the real connected component.
+The following polynomials (36), (37), (38), (39), are all invariant under this involution, so their virtual components are represented by single connected components of the complements of the discriminants.
+The polynomial
+(36) $(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
+with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$.
+Picture $(m)$ is represented by the polynomials
+(37) $x(x-y^{2}-t)(x+y^{2}+t),\quad t>0.$
 The polynomials
-$(x-t)^{3}-(x-t)y^{4}+3(x-t)^{2}+5(x-t)y^{2}-4(x-t)-Ct^{3}$
-with arbitrary $t>0$ realizes the picture $(a)$ for $C<6-\frac{14}{3}\sqrt{\frac{7}{3}}$, the picture $(b)$ for $C\in\left(6-\frac{14}{3}\sqrt{\frac{7}{3}},-\frac{1}{2}\right)$, and the picture $(d)$ for $C\in\left(-\frac{1}{2},0\right)$.
-All of these polynomials are even with respect to the variable $y$, so their virtual components are represented by single connected components of the complements of the discriminants.
-Picture $(m)$ is represented by the polynomials
-$x(x-y^{2}-t)(x+y^{2}+t),\qquad t>0,$
-which are also self-symmetric.
+(38) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
+realize picture $(c)$ for any $C\in(0,1/2)$.
 The polynomials
-$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+3t(x-t)y^{2}+Ct^{3}$
-realize picture $(c)$ for any $C\in(0,1/2))$. They are also invariant under the involution (34).
-The polynomials
-$(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$
+(39) $(x-t)^{3}-(x-t)y^{4}+3t(x-t)^{2}+5t(x-t)y^{2}-4t^{2}(x-t)+\frac{t^{2}}{10}y^{2}-0.35t^{3}$
 realize picture $(f)$.
@@
 To realize the picture $(e)$, we start with the polynomial
-(37) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2)$
+(40) $x(x+(y-2)^{2}-2)(x+(y+2)^{2}-2),$
 whose zero-level set is shown in Fig.
@@
 Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal quasihomogeneous part will be of class $J_{10}^{3}$.
-The polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values.
+The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values.
 It will also have two critical points in the complex domain.
 Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain.
-The topological picture of the zero-level set of the resulting polynomial will be the desired one.
-Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see Table 2.
-Finally, we apply the operator $\bf T$} to reduce the polynomial to the form (1).
-The resulting polynomial realizes the picture $(e)$, is of class $\Omega_{3}$, and is invariant under the involution (34).
+The zero-level set of the resulting polynomial will have the desired shape shown in picture (e) of Fig.
+5.
+Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see (32).
+Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$.
+The resulting polynomial realizes the picture $(e)$, and is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$.
+$\Box$
+4. The cases $P_{8}$
+4.1.
+Two classes of function singularities $P_{8}^{1}$ and $P_{8}^{2}$ in three variables have the canonical normal form
+(41) $x^{3}+y^{3}+z^{3}-3Axyz,\quad A\neq 1,$
+with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$.
+The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected components in ${\mathbb{R}P}^{2}$.
+Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to the plane with a single singular point at the origin.
+The zero set of a polynomial of class $P_{8}^{2}$ consists of such a surface and a surface that is homeomorphic to the standard quadratic cone.
+It follows easily from Morse theory that the closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
+The standard versal deformations of $P_{8}^{1}$ and $P_{8}^{2}$ singularities consist of the polynomials
+(42) $f_{\lambda,A}=f_{A}+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}z+\lambda_{5}xy+\lambda_{6}xz+\lambda_{7}yz,$
+where $f_{A}$ is the corresponding normal form (41).
+Definition 8.
+A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
+Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$.
+4.2. The homological invariant of real and virtual components
+Let $f_{0}$ be a function singularity of class $P_{8}^{1}$ or $P_{8}^{2}$ having the form (41).
+Let $D$ be a small neighborhood of the point $\{f_{0}\}$ in the space $\Xi_{1}$ or $\Xi_{2}$.
+Let $B$ be a sufficiently large (compared to $D$) closed ball in ${\mathbb{R}}^{3}$, and let $f_{\lambda}$, $\lambda\in D$, be a small perturbation of $f_{0}$.
+Define $T_{-}(\lambda)$ as a real number slightly lower than the minimum of 0 and the lowest real critical value of $f_{\lambda}$.
+Similarly, define $T_{+}$ as a number slightly higher than the maximum of 0 and the highest real critical value of $f_{\lambda}$.
+For any $\lambda\in D$, define three sets: $W_{\lambda,0}=f_{\lambda}^{-1}((-\infty,0])\cap B$, $W_{\lambda,-}=f_{\lambda}^{-1}((-\infty,T_{-}])\cap B,$ and $W_{\lambda,+}=f_{\lambda}^{-1}((-\infty,T_{+}])\cap B$.
+The sets $W_{\lambda,-}$ and $B\setminus W_{\lambda,+}$ form then a trivializable fiber bundle over $D$.
+These sets contain their intersections with $\partial B$ as deformation retracts.
+In the case $P_{8}^{1}$ these intersections are just the discs.
+In the case $P_{8}^{2}$ each of them homeomorphic to the disjoint union of a disc and an annulus.
+In particular, the reduced homology groups $\tilde{H}_{i}(W_{\lambda,-})$ and $\tilde{H}_{i}(B\setminus W_{\lambda,+})$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
+Definition 9.
+The homology index $\mbox{HI}(f_{\lambda})$ of a non-discriminant function $f_{\lambda}$, $\lambda\in D$ is the triple of Betti numbers $b_{0},b_{1},b_{2}$ of the relative homology group $H_{*}(W_{\lambda,0},W_{\lambda,-})$.
+It is convenient to consider this homology group in the context of the exact homological sequence of the triple
+(43) $(W_{\lambda,+},W_{\lambda,0},W_{\lambda,-}).$
+[…]
+Clearly, the invariant Ind is equal to the alternating sum of three elements of the homology index.
+4.3. The symmetry group and the relation between the real and virtual components
+The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following exceptions: the frames of vanishing cycles in the 2-homology groups of zero-level sets of polynomials consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values.
+We also have the following analog of Proposition 2
+Proposition 32 (see [22], Corollary 8).
+1. Any virtual Morse function of the type $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
+2.
+If two virtual Morse functions of type $P_{8}^{1}$ (respectively, $P_{8}^{2}$) belong to the same virtual component, then they are associated with two generic Morse polynomials of class $\Xi_{1}$ (respectively, $\Xi_{2}$) that are isotopic to each other within the set of Morse polynomials.
+$\Box$
+The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ preserves the normal form (41) and acts on the spaces $\Xi_{1}$ and $\Xi_{2}$.
+This action is generated by the following transformations:
+(44) $f(x,y,z)\mapsto f(y,x,z)\quad\mbox{and}\quad f(x,y,z)\mapsto f(y,z,x).$
+The first transformation is an involution, and the second is of order three.
+In particular, this group can act on the set of the connected components of the discriminant complement in the spaces $\Xi_{i}$.
+The transformations (44) preserve all topological characteristics of the functions.
+Thus, all connected components in the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
+Additionally, the involution
+(45) $f(x,y,z)\mapsto-f(-x,-y,-z)$
+also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It preserves the invariant Ind and defines involutions on both sets of real and virtual components of these classes.
+Proposition 33.
+If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class that contain $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$.
+Definition 10.
+For any polynomial $f:{\mathbb{R}}^{3}\to{\mathbb{R}}$ of degree three with principal homogeneous part of the form (41), ${\bf T}(f)$ is the unique polynomial of the form (42) with the same principal homogeneous part, which is obtained from $f$ by a parallel translation in ${\mathbb{R}}^{3}$.
+Namely, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
+Lemma 34.
+[…]
+(46) ${\bf T}\left(x^{3}+y^{3}+z^{3}+6xyz+(x+y-2z)^{2}\right)$
+and two other polynomials obtained from it by permuting the coordinates.
+Proof.
+According to Lemma 35 of [22], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$.
+Three of these polynomials (mentioned in formula (26) of [22]) are locally equivalent to $\xi^{3}-\eta^{2}-\zeta^{2}+\frac{64}{3}$ at the $A_{2}$ point.
+The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$.
+$\Box$
+Lemma 35.
+The set of all polynomials of class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
+(47) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(1+\sqrt{3})xyz+(x+y+(1+\sqrt{3})z)^{2}\right),\$
+[…]
+Proof. This follows from statement 2 of Lemma 44 of [22] (the coefficients $\varkappa$ and $c$ in formula (45) of [22] are fixed by our conditions on the critical values). $\Box$
+Proof of Proposition 33. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 34 (respectively, 35) repeats the deduction of Proposition 8 from Lemma 9. $\Box$
+4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
+Proposition 36 (see [18], Proposition 7).
+1. There are exactly 6503 virtual functions of the class $P_{8}^{1}$.
+2.
+There are exactly seven virtual components of the formal graph of type $P_{8}^{1}$: two with $\mbox{\rm Ind}=1$, two with $\mbox{\rm Ind}=-3$, and one virtual component with each value 0, 1-1, and 2-2 of the Ind invariant.
+[…]
+Theorem 37.
+Each virtual component of type $P_{8}^{1}$ is associated with only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$.
+So, there are exactly seven such connected components.
+Proof. The invariant Ind of the polynomials
+(48) $x^{3}+y^{3}+z^{3}-3(x+y+z)+C$
+is equal to 0 for C<6C<-6 and $C>6$, to $1$ for $C\in(-6,-2)$ and $C\in(2,6)$, and to 2-2 for $C\in(-2,2)$.
+These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates.
+Therefore, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
+Two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2.
+For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$.
+According to the previous paragraph, the associated virtual components are also different.
+The polynomials
+(49) ${\bf T}\left(x^{3}+y^{3}+z^{3}+3(x+y+z)^{2}-(x^{2}+y^{2}+z^{2})/2+\frac{7}{36}\right)$
+and
+(50) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(x+y+z)^{2}+(x^{2}+y^{2}+z^{2})/2-\frac{7}{36}\right)$
+have $\mbox{Ind}=-3$.
+They are related to each other by the involution (45).
+Hence, by the last statement of Proposition 36, they realize two different virtual components.
+Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components.
+[…]
+Lemma 38.
+For each linear function $l\in L$,
+1)) the polynomial
+(51) $x^{3}+y^{3}+z^{3}-l^{2}$
+has only one real critical point, namely a critical point of class $D_{4}^{+}$ at the origin.
+In particular, this polynomial has four critical points ((counted with multiplicities)) in the complex domain.
+2)) For any sufficiently small $\varepsilon>0$, the polynomial
+(52) $F(l)\equiv x^{3}+y^{3}+z^{3}-l^{2}-\varepsilon p_{l}^{2}$
+has only two critical points in ${\mathbb{R}}^{3}$: a critical point of type $A_{3}$ at the origin, at which this polynomial is equal to
+(53) $\xi^{4}-\eta^{2}-\zeta^{2}$
+in some local coordinates, and a Morse critical point with Morse index $1$ and negative critical value.
+Proof.
+1.
+At each critical point of the polynomial (51), the sum of its three partial derivatives is equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain.
+In coordinates $l,p_{l},q_{l}$, this polynomial has the form
+(54) $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}+l\psi_{2}(p_{l},q_{l})+l^{2}\psi_{1}(p_{l},q_{l},l),$
+where $\psi_{2}$ and $\psi_{1}$ are some homogeneous polynomials of degrees 2 and 1, respectively.
+The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [2], §12, this polynomial has the normal form $Z^{2}+X^{3}+XY^{2}$ of $D_{4}^{+}$ class in some local coordinates.
+2.
+Let $\alpha$ be the coefficient of the polynomial $\psi_{2}$ at the monomial $q_{l}^{2}$.
+Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
+The substitution $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2},P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial to the polynomial with the lowest quasihomogeneous part
+$\tilde{l}^{2}-\varepsilon P_{l}^{2}+\left(\frac{\alpha^{2}}{4}+\frac{1}{4\varepsilon}\right)q_{l}^{4}\ .$
+Again, using the techniques of §12 of [2], this function can be reduced to the standard form (53) of $A_{3}$ singularities via a local change of coordinates.
+If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not bring four non-real critical points of (51) back to the real domain.
+On the other hand, by the index considerations the obtained polynomial $F(l)$ has one Morse critical point in addition to the $A_{3}$ point at the origin.
+This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l$.
+It is easy to calculate that for $l\equiv x$ it is equal to 1.
+Our critical point of class $A_{3}$ is topologically equivalent to a Morse critical point with Morse index 2 (and can be perturbed to such a point by an arbitrarily small change of the polynomial $F(l)$).
+The Morse-like complex of the polynomial $F(l)$ is acyclic, therefore the critical value of this $A_{3}$ type critical point should be higher than that of the only other critical point of Morse index 1.
+$\Box$
+Denote by $C(l)$ the critical value of the polynomial $F(l)$ at its Morse critical point. Then the polynomial
+$\tilde{F}(l)\equiv F(l)-\frac{C(l)}{2}$
+is non-discriminant and has $\mbox{Ind}=-1$.
+The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it.
+The set of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$.
+Thus the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement.
+$\Box$
+4.5. One-dimensional cohomology of discriminant complements of $P_{8}^{1}$ singularities
+Theorem 39.
+The first cohomology group of the component of the discriminant complement of the $P_{8}^{1}$ singularities that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
+Proof.
+The zero-level set of the polynomials in this component is homeomorphic to the plane with a handle.
+Indeed, the topology of these sets remains constant along the component.
+For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to the plane) by passing through a Morse critical point of Morse index 1 once.
+In particular, the first homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$.
+These homology groups form a local system over the component, in particular the fundamental group of the component acts on them by monodromy.
+For each $l\in L$, one basic homology class of the zero-level set of $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$.
+When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement.
+The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation.
+$\Box$
+4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularity
+Definition 11.
+The invariant Card of a virtual component is the number of its vertices.
+The involution (45) preserves this invariant.
+Table 3. Statistics for $P_{8}^{2}$ singularity
+No. 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
+Ind 3-3 2-2 1-1 0 1
+Card 258 156 60 60 1216 336 336 1318 844 844 1648 1648 262 94 94
+Proposition 40 (see [18], Proposition 8).
+1. There are exactly 9174 distinct virtual functions of $P_{8}^{2}$ singularity class.
+2. These virtual functions are distributed over exactly fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and third rows of Table 3.
+3. All pairs of different virtual components of $P_{8}^{2}$ singularity with the same invariant Card value can be realized by generic polynomials connected by involution ((45)).
+Theorem 41.
+Each of the ten virtual components of the formal graph of type $P_{8}^{2}$ that are mentioned in columns 1, 3, 4, 6, 7, 11, 12, 13, 14, and 15 of Table 3 is associated with only one real connected component of the discriminant complement.
+Proof of this theorem takes the rest of this section.
+All of the polynomials given in the next formulas (55), (56), (57), and (58) are invariant under the $S(3)$ action.
+Thus, for any polynomial $f$ of this list and any noncritical value $C$ of $f$, the set-theoretical invariant of the polynomial fCf-C is a virtual component associated with only one real component of the discriminant complement.
+We will show that these virtual components, or those obtained from them via the involution (45), include all virtual components with Card equal to 258, 336, 1648, 262, and 94.
+The proof of Theorem 41 for the components with $\mbox{Card}=60$ is more implicit, see § 43.
+Components with $\mbox{\rm Ind}=0$ and $\mbox{\rm Card}=1648$.
+For each polynomial $f$ of class $\Xi_{2}$, the polynomials $f+C$ and fCf-C with sufficiently large $C$ realize two different virtual components with $\mbox{Ind}=0$.
+Indeed, the homology index of one of these components is $(000)$, and the other’s is $(011)$.
+In particular, the $S(3)$-invariant polynomials
+(55) $x^{3}+y^{3}+z^{3}-4xyz+C\qquad\mbox{and}\qquad x^{3}+y^{3}+z^{3}-4xyz-C$
+represent these two components for any $C>0$. These polynomials are not Morse, but any their sufficiently small generic perturbations are associated with virtual functions with $\mbox{Card}=1648$.
+Components with $\mbox{\rm Card}=262$. The polynomial
+(56) $x^{3}+y^{3}+z^{3}-4xyz+(x+y+z)$
+has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2.
+Thus, it has invariant values $\mbox{Ind}=1$ and $\mbox{HI}=(00\bar{1})$.
+This polynomial is invariant under the involution (45).
+According to the last statement of Proposition 40, it cannot represent a component of the discriminant complement with $\mbox{Card}=94$.
+Therefore, it belongs to the component with $\mbox{Card}=262$.
+Components with $\mbox{\rm Card}=94$ and $336$. The polynomial
+(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz-\frac{3}{2}(x^{2}+y^{2}+z^{2})$
+has eight real Morse critical points: one local maximum with critical value 0, three points of Morse index $2$ with value $\frac{11}{49}$, three points of index $1$ with value $\frac{1}{2}$, and one more point of index $2$ with value $\frac{3}{2}$.
+Therefore, the index $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{3}{2}$, to 1 for $C\in\left(-\frac{3}{2},-\frac{1}{2}\right),$ to 2-2 for $C\in\left(-\frac{1}{2},-\frac{11}{49}\right),$ to $1$ for $C\in\left(-\frac{11}{49},0\right),$ and to 0 for $C>0$.
+The group $H_{*}(W_{\lambda,+},W_{\lambda,0})$ corresponding to the polynomials fCf-C with $C\in\left(-\frac{11}{49},0\right)$ is isomorphic to $\mathbb{Z}$} in dimension three and is trivial in all other dimensions.
+Therefore, by the exact sequence of the triple (43), the homology index $\mbox{HI}(f+\frac{1}{7})$ is equal to $(0,1,2)$.
+The involution (45) moves this polynomial to that with $\mbox{HI}=(1,0,0)$.
+In particular, these two polynomials represent two different components, and the associated virtual components are also different.
+According to Table 3, these are the components with $\mbox{Card}=94$.
+The homology index $\mbox{HI}(f-C)$ with $C\in\left(-\frac{3}{2},-\frac{1}{2}\right)$ is $(0,0,1),$ because the corresponding group $H_{*}(W_{\lambda,0},W_{\lambda,-})$ is generated by the class of one critical point of Morse index $2$.
+Therefore, it is again the component with $\mbox{Card}=262$.
+[…]
+Components with $\mbox{\rm Card}=258,$ $336,$ and $94$. The polynomial
+(58) $f=x^{3}+y^{3}+z^{3}-12xyz-\frac{1}{2}(x^{2}+y^{2}+z^{2})$
+has eight real Morse critical points: a local maximum with critical value 0, three points with Morse index 2 and critical value $\frac{11}{3969}$, one point more with Morse index 2 and critical value $\frac{1}{162}$, and three points with index 1 and critical value $\frac{1}{54}$.
+Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{1}{54}$, to 3-3 for $C\in\left(-\frac{1}{54},-\frac{1}{162}\right)$, to 2-2 for $C\in\left(-\frac{1}{162},-\frac{11}{3969}\right)$, to $1$ for $C\in\left(-\frac{11}{3969},0\right)$, and to 0 for $C>0$.
+Lemma 42.
+[…]
+Proof.
+By Proposition 40 of [22], the intersection matrix of vanishing cycles of the canonical basis for all generic polynomials of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Dynkin diagram shown in
+Figure 7.
+Fig.
+7.
+This diagram does not contain edges of multiplicity 2-2.
+Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2.
+Therefore, the virtual component associated with our polynomial is the unique remaining virtual component with $\mbox{Ind}=-3$.
+$\Box$
+Remark 43.
+It is easy to see that the polynomials (58) with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
+Components with $\mbox{Card}=60$.
+[…]
+Theorem 44.
+There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ realizing a virtual component with $\mbox{\rm Card}=60.$
+Proof. One of the two virtual components, each of which has $\mbox{Card}=60$, contains a virtual function that models a generic polynomial $f_{0}$ of class $\Xi_{2}$ such that
+1) all of its critical points are real,
+2) the three highest critical values are positive and the other five values are negative,
+3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is shown in (59),
+4) the Morse indices of the points 1, 2, 3, and 5 are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
+According to Proposition 32, there is a real polynomial $f_{0}\in\Xi_{2}$ with these data.
+According to Proposition 11 of [22], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), all odd Morse indices are equal to 1, and all even Morse indices are equal to 2.
+Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by their increase, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
+Using the Lyashko–Looijenga map (in the form of [10]) we can lift the path $i$ uniquely into a path $I:[0,1]\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$.
+As in the proof of Theorem 7 of [22], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined.
+Indeed, according to Proposition 2 of [10], this is guaranteed by the fact that all quadratic forms corresponding to the limit critical values are elliptic.
+The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with value $c_{-}$ and three critical points with value $c_{+}$.
+Lemma 45.
+This polynomial $I(1)$ is $S(3)$-invariant.
+Proof.
+The critical points of the polynomial $I(1)$ are ordered according to the previous ordering of the critical points of the polynomial $f_{0}$, from which they were obtained via continuous deformation along the path $I$.
+This ordering can be uniquely extended to a continuous system of orderings of the critical points of each polynomial $\tilde{f}$ from a sufficiently small neighborhood $U$ of the point $I(1)$ in $\Xi_{2}$.
+Define the map $\Lambda:U\to{\mathbb{R}}^{8}$, mapping each point $\tilde{f}\in U$ to the sequence of its correspondingly ordered critical values.
+By the versality property, this map is a local diffeomorphism.
+Let $\varepsilon>0$ be a very small number.
+For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$, respectively.
+Different permutations $\sigma$ of critical points define six different points $\tilde{f}_{\sigma}\in U$.
+All of these points are associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
+As in the proof of Proposition 33 (repeating the proof of Proposition 8), the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$.
+Therefore, each such element maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point from the same neighborhood.
+By continuity, it also maps the polynomial $I(1)$ to itself.
+$\Box$
+By the construction, the Card invariant of the virtual function associated with arbitrarily small perturbations of the polynomial $I(1)$ is equal to 60. This proves Theorem 44. $\Box$
+The involution (45) moves the polynomial $I(1)$ to a polynomial in the other component of the discriminant complement with $\mbox{Card}=60$.
+Thus, this component is also represented by an $S(3)$-invariant polynomial.
+This concludes the proof of Theorem 41.
 $\Box$
 5. Petrovskii lacunas
-The above enumeration of the connected components of the discriminant complements at the $X_{9}$ and $J_{10}$ singularities proves all of the conjectures in [11] concerning the number of local Petrovskii lacunas at the corresponding singular points of wavefronts.
-In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity when the dimension $N$ and the index $i_{+}$ are both even is component $(a)$ of Fig.
+The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs.
+In this theory, generating families of functions play a crucial role.
+Wave functions are defined on the parameter spaces of these families via integrals along certain cycles in the zero-level sets of the corresponding functions.
+Wavefronts are the discriminant sets of these families.
+Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, and the wave function is regular as well.
+However, over the discriminant sets of parameters, the level sets containing the integration cycles degenerate and the wave function may be singular.
+The singularity of a wave function in a neighborhood of a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point.
+The type of this asymptotic behavior depends primarily on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle.
+The homology classes depend on the wavefront complement components from which we approach the wavefront point.
+Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
+If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [16]).
+Versal deformations of stably equivalent function singularities can be naturally identified with each other.
+This identification establishes a one-to-one correspondence between the components of their discriminant complements.
+The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number of variables, and on the parities of the inertia indices of the quadratic part of the singularity.
+For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas.
+On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or the stably equivalent functions in n+2k1n+2k-1 variables generally has a different set.
+Theorem 46.
+The numbers of local lacunas near parabolic singularities of wavefronts in versal generating families are given in Table 4.
+Table 4. Numbers of local lacunas at parabolic singularities
+Singularity NN even NN even NN odd NN odd
+class $i_{+}$ even $i_{+}$ odd $i_{+}$ even $i_{+}$ odd
+$P_{8}^{1}$ 0 0 22 0
+$P_{8}^{2}$ 0 0 33 0
+$X_{9}^{+}$ 11 0 22 0
+$X_{9}^{-}$ 11 0 0 22
+$X_{9}^{1}$ 0 0 0 0
+$X_{9}^{2}$ 0 44 0 0
+$J_{10}^{3}$ 0 11 0 0
+$J_{10}^{1}$ 0 0 0 0
+In comparison with the results of [17], this table replaces the lower estimate $\geq 2$ for $P_{8}^{2}$ with the exact number 3, and replaces some five estimates of the form $\geq k$ with the same numbers $k$.
+In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig.
 1.
 The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure.
-The four lacunas at the $X_{9}^{2}$ singularity, with even $N$ and odd $i_{+}$, are shown in picture $(i)$ of Fig.
+The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig.
 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (17) to these two.
-The unique lacuna at the $J_{10}^{3}$ singularity, with even $N$ and odd $i_{+}$, is shown in picture $(m)$ of Fig.
+The unique lacuna at the $J_{10}^{3}$ singularity with even $N$ and odd $i_{+}$ is shown in picture $(m)$ of Fig.
 5.
+The two local lacunas of the $P_{8}^{1}$ singularity at odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$.
+Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$.
+The only one new local lacuna found now is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.

### v3

@@
 Abstract.
-We provide a complete list of the connected components of the spaces of non-discriminant functions within standard versal deformations of function singularities of classes $X_{9}$, $J_{10}$ and $P_{8}^{1}$ (as well as a partial list for the remaining class, $P_{8}^{2}$).
+We list all connected components of sets of non-discriminant functions within versal deformations of all parabolic function singularities.
+These singularities are the second most common family of singularity classes of smooth functions after simple singularities.
 Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [24] with the same title.
 As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs.
-In particular, we discover a new local lacuna at the $P_{8}^{2}$ singularities.
-We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, unlike all simple singularities.
+We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
+These results are applications of a general method for investigating non-singular perturbations of real function singularities.
+An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
 Key words and phrases:
@@
 $1.$
-Discriminant varieties of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of these functions are singular.
+Discriminant varieties of families of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of the corresponding functions are singular.
 These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g.
-[1], [2], [3], [8], [15][16], [17].
+[2], [3], [4], [9], [14], [20], [22], [23].
 Many important functions in these theories arise as integral transforms.
-These functions are defined on the parameter spaces of certain function deformations (called generating families) and are equal to the integrals along the cycles in the zero-level sets of the corresponding functions.
-These integrals are regular outside the discriminant varieties of the deformations.
-Their analytic behavior near the discriminants depends on the singularity type of the deformed discriminant function and on the connected component of the complement of the discriminant.
-Thus, enumerating these components is one of the first steps in studying the analytical properties of these integral transforms.
+Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals along the cycles in the zero-level sets of the corresponding functions.
+These integral functions are regular outside the discriminant varieties of the families.
+The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding discriminant function and on the connected component of the complement of the discriminant, from which we approach this point.
+Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
 Definition 1 (see [3]).
-A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},0)\to({\mathbb{R}},0)$ at its critical point.
-Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$.
+A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},a)\to({\mathbb{R}},0)$ at its critical point.
+Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$ (i.e., they have the same expression in appropriate local coordinate systems centered at $a$ and $b$).
 Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
 $f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
-An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin in ${\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic.
-The discriminant of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value (or, equivalently, the zero-level set $f_{\lambda}^{-1}(0)$ is singular near real space).
+An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic.
+The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level set of the analytic continuation of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ is singular at this point).
 For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation.
-The discriminant sets of all versal deformations of stably equivalent singularities are topologically equivalent, and the components of their complements are in a one-to-one correspondence.
-Therefore, for any stable equivalence class of singularities it is sufficient to enumerate the components for only one versal deformation of one function of this class.
+The discriminant varieties of all versal deformations of stably equivalent singularities are topologically equivalent, and the connected components of their complements in the parameter spaces are in a one-to-one correspondence.
+Therefore, for any stable equivalence class of singularities it is sufficient to enumerate these components for only one versal deformation of one function of this class.
 The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V.
 Arnold, see [3].
-All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of functions in two variables.
-The components of discriminant complements for these singularities were listed earlier, see § 1.3.
-Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and $P_{8}$ singularities in three variables
+All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of such functions.
+The components of discriminant complements for simple singularities were listed earlier, see § 1.3.
+Table 1. Normal forms of simple, $X_{9}$ and $J_{10}$ singularities in two variables and of $P_{8}$ singularities in three variables
 Notation Normal form Restriction
@@
 $A_{2k}$ $x^{2k+1}\pm y^{2}$ $k\geq 1$
-$D_{2k}^{\pm}$ $x^{2}y\/\pm y^{2k-1}$ $k\geq 2$
-$D_{2k-1}$ $\pm(x^{2}y\/+y^{2k-2})$ $k\geq 3$
+$D_{k}^{\pm}$ $x^{2}y\/\pm y^{k-1}$ $k\geq 4$
 $E_{6}^{\pm}$ $x^{3}\pm y^{4}$
@@
 $E_{8}$ $x^{3}+y^{5}$
-$X_{9}^{\pm+}$ $\pm(x^{4}+2Ax^{2}y^{2}+y^{4})$ $|A|<1$
+$X_{9}^{\pm}$ $\pm(x^{4}+2Ax^{2}y^{2}+y^{4})$ $|A|<1$
 $X_{9}^{1}$ $xy(x^{2}+2Axy+y^{2})$ $|A|<1$
@@
 $P_{8}^{2}$ $x^{3}+y^{3}+z^{3}-3Axyz$ $A>1$
-In this work we list analogous components for the next natural class in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V.
+In this work we list all analogous components for the next natural collection of equivalence classes in the hierarchy of function singularities, the parabolic singularities, which were also distinguished by V.
 Arnold.
-The classification of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$.
+The list of complex parabolic singularities up to the stable equivalence consists of three classes $P_{8}$, $X_{9}$, and $J_{10}$.
 The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
-Functions of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables.
-Any function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any function in three variables of one of classes $P_{8}^{*}$ can be reduced to a normal form indicated in Table 1 by an appropriate choice of local coordinates centered at the critical point.
+Singularities of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables.
+Any real function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any singularity in three variables of one of classes $P_{8}^{*}$, has the normal form indicated in Table 1 in appropriate local coordinates centered at the critical point.
 Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
-We give complete lists of components for all parabolic singularity classes except $P_{8}^{2}$, for which we have only partial results.
 Table 2. Numbers of rigid isotopy and topological classes of non-discriminant perturbations of parabolic singularities
 Singularity class $X_{9}^{\pm}$ $X_{9}^{1}$ $X_{9}^{2}$ $J_{10}^{1}$ $J_{10}^{3}$ $P_{8}^{1}$ $P_{8}^{2}$
-Isotopy 7 14 52 13 33 7 $\geq 15$
+Rigid isotopy 7 14 52 13 33 7 15
 Topology 6 12 50 10 31
@@
 By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood.
-The topological type is an invariant of the components of the discriminant complements.
-The “rigid isotopy” classification of parabolic singularity perturbations given in the present work is only slightly stricter than the topological classification, see Table 2.
+Clearly, the topological type is an invariant of the connected components of the discriminant complements.
+Our “rigid isotopy” classification of non-discriminant perturbations separating all these components is only slightly stricter than the topological classification, see Table 2.
 Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [8].
@@
 However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [24] consist of pairs of different connected components.
-Also, we present below partial results for the case $P_{8}^{2}$, which prove that at least 10 out of 15 classes listed in [18] are indeed represented by only one connected component each.
-Thus, the number 11 above is replaced in Table 2 by 13, and the number 15 is still under suspicion.
-These results are summarized in Theorems 5, 13, 18, 28, 31, 37, and 41 below.
+Thus, the number 11 above is replaced in Table 2 by 13.
+These results are summarized in Theorems 6, 14, 19, 29, 32, 38, and 42 below.
 1.3. On the discriminant complements for simple singularities
@@
 E.
-Looijenga [13] proved that all local components of discriminant complements of real simple singularities are topologically trivial.
-Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 12 and 39.
+Looijenga [16] proved that all local components of discriminant complements of real simple singularities are homotopically trivial.
+Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 40.
 1.5. An application to hyperbolic PDEs.
-In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [1], [8], [16]).
-Near the non-singular points of wavefronts (corresponding to the Morse functions) the lacunas were enumerated in [6] and [5].
+In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [17], [22]).
+Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6].
 Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [20].
@@
 We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
-For the singularities of classes $P_{8}^{1}$, $X_{9}^{*}$ and $J_{10}^{*}$, these results follow directly from our complete lists of components of discriminant complements.
-Our partial results on enumerating components for the $P_{8}^{2}$ singularities are sufficient to prove that our new list of local lacunas is also complete in that case.
 2. Results for the $X_{9}^{*}$ classes
@@
 Real parabolic function singularities of type $X_{9}$ form four stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
-The standard versal deformations of functions in all these classes have the same form: they are families of polynomials
+The standard versal deformations of these singularities in all four cases have the same form: they are families of polynomials
 (1) $f_{A}(x)+\lambda_{1}+\lambda_{2}x+\lambda_{3}y+\lambda_{4}x^{2}+\lambda_{5}xy+\lambda_{6}y^{2}+\lambda_{7}x^{2}y+\lambda_{8}xy^{2},$
 depending on nine parameters $\lambda_{1},\dots,\lambda_{8},$ and $A$, where $f_{A}$ is the corresponding normal form from Table 1.
-According to versality theory (see, for example, [2]), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families.
+According to versality theory (see [3], § 8), studying all other deformations of singularities of stable equivalence classes $X_{9}^{*}$ can be reduced to studying these families.
+The entire study of $X_{9}^{-}$ singularities and their deformations can be reduced to that of $X_{9}^{+}$ via the multiplication by 1-1.
 Definition 2.
@@
 These varieties divide the spaces $\Phi_{*}$ into connected components.
-Enumerating of these components is our goal in this section.
+Enumerating these components is our goal in this section.
 2.2. Virtual functions and virtual components
 In §2 of [24] we introduced the set-valued invariant of connected components of the discriminant complements. Let us briefly review this construction in the case of $X_{9}$ singularities.
-The virtual function associated with a generic real polynomial is a collection of topological characteristics of the complex zero-level set of this polynomial, depending also on the collections of paths in ${\mathbb{C}}^{1}$ defining the bases of vanishing cycles in the homology of this level set.
+A virtual function associated with a generic real polynomial $f$ is a collection of topological characteristics of the complex zero-level set of this polynomial.
 Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
+  1. (1)
     the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
+  2. (2)
     the string of intersection indices of these nine cycles with the set of real points,
+  3. (3)
     the Morse indices of real critical points of $f$, and
+  4. (4)
     the numbers of positive and negative critical values of $f$.
-The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model surgeries of real functions.
+Remark 1.
+Besides the function $f$, a virtual function depends on the collection of paths in ${\mathbb{C}}^{1}$ defining the base of vanishing cycles in the homology of the level set.
+If the function has no more than one pair of non-real critical values, then there is only one canonical system of paths, and hence only one associated virtual function.
+If all critical points of a polynomial are real, then the intersection indices from item (2) are easily determined by the remaining elements (1), (3), and (4) of a virtual function and hence may not be included in its description, see Theorem V.1.3 in [22].
+The (abstract) virtual functions of class $\Phi_{*}$ include all virtual functions associated with generic real polynomials of this class, as well as all analogous collections of data that can be obtained from them via standard flips that model elementary surgeries of real functions.
 The formal graph of any type $X_{9}^{*}$ is the graph, whose vertices correspond to all the virtual functions of the corresponding class $\Phi_{*}$, and whose edges correspond to the standard flips.
@@
 The formal graphs of all simple and parabolic singularities are finite.
-Enumerating the virtual components of such a graph is a combinatorial problem that can be solved using a computer program.
+Enumerating the virtual components of such a graph is a combinatorial problem.
+An (available) computer program solving it is described in [22] and [24].
 The work [24] provides the results of these computations for all parabolic singularities, as well as the immediate consequences of these computations in the terms of real functions.
-The remaining problem is to find all real connected components with this value of the set-valued invariant for each virtual component.
+The remaining problem is to find all real connected components associated with each virtual component.
 The results of [12] imply that in the case of parabolic singularities all standard flips of virtual functions can be realized by real surgeries of generic real functions.
@@
 a)) every virtual function of this class is associated with a generic real polynomial of the same class;
-b)) every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$.
+b)) for any generic polynomial $f$ of this class, every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$.
 $\Box$
 Remark 4.
-In the statement of Proposition 3 of [21], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with nondegenerate principal homogeneous parts.
+In the statement of Proposition 3 of [27], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with non-degenerate principal homogeneous parts.
 However, the entire construction proving this proposition in [27] can be performed within the spaces of the versal deformations (1).
 Thus, it also proves the above Proposition 3.
-These constructions enable us to estimate from above the number of isotopy classes of generic Morse perturbations associated with any given virtual function.
-These estimates are given in Propositions 8, 15, and 20 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately.
-Consequently, we obtain estimates of the number of connected components of the complement of a real discriminant which are associated with each virtual component.
+These constructions provide upper bounds of the numbers of isotopy classes of generic Morse perturbations associated with virtual functions.
+These bounds are given in Propositions 9, 16, and 21 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately.
+Consequently, we obtain estimates of the numbers of connected components of the real discriminant complement which are associated with each virtual component.
 Specifically, in the cases $X_{9}^{+}$ and $X_{9}^{2}$, all real connected components of discriminant complements associated with the same virtual component can be obtained from each other by the action of a transformation group of order eight.
@@
 In [27], [28], and [29], we use the same strategy with a different definition of virtual components.
-In these works, we classify the Morse perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions.
+In these works, we classify the generic perturbations of function singularities up to isotopy within the space of Morse functions instead of the space of non-discriminant functions.
 Thus, the virtual components of the formal graph are determined there by removing the edges that change the Morse type.
 Definition 3.
-For any Morse polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
-It is easy to see that this number is an invariant of the connected components of the discriminant complements (and of associated virtual components).
+For any generic polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
+It is easy to see that this number is an invariant of the connected components of the discriminant complements, and also of associated virtual components.
 2.3. Enumeration of components for $X_{9}^{+}$ singularities
@@
 Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves.
-In our case, however, the isotopies are allowed only within a much smaller space $\Phi_{+}$.
+In our case, however, the isotopies are allowed only within the space $\Phi_{+}$ which is much smaller than the space of all degree four polynomials.
 In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
 Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$.
-(For this existence statement, the parabolicity of $X+9$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
+(For this existence statement, the parabolicity of $X_{9}$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
 First, suppose that it is the same point $I(1)$.
@@
 By Proposition 9, any virtual component mentioned in Proposition 8 is associated with at most eight real connected components.
-Let us prove that in fact each virtual component represented by a polynomial shown in Fig.
-1 is associated with a single real connected component.
+Let us prove that in fact each of these seven virtual components is associated with a single real connected component (that is, these components are $G_{0}$-invariant).
 Pictures $(a)$, $(c)$, $(b)$, $(f)$ of Fig. 1 are realized, respectively, by polynomials
@@
 (3) $x^{4}+2Ax^{2}y^{2}+y^{4}+2x^{2}-2y^{2}+1/2$
-with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
+with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of it and the polynomial
 (4) $x^{4}+2Ax^{2}y^{2}+y^{4}-2x^{2}+2y^{2}+1/2\ .$
@@
 Proof.
-If we replace $A$ by 1 in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$.
+If we take $A=1$ in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$.
 Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle.
-If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$.
+After replacing $A=1$ by $1-\varepsilon$ with sufficiently small $\varepsilon$, the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$.
 It also remains empty outside this domain.
@@
 The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
-Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are $k\geq 2$ non-nested ovals.
+Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals.
 Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows.
@@
 (13) $f(x,y)\mapsto-f(x,-y)$
-acts on the space $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement.
+acts on the space $\Phi_{1}$.
+It preserves the discriminant variety and hence permutes the connected components of its complement.
 This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y).
@@
 There are exactly ten virtual components of type $X_{9}^{1}$.
-They split into five pairs of virtual components related to each other by the involution induced by ((13)).
-The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d), and ((f)) of Fig.
+They split into five pairs of virtual components related to each other by the involution ((13)).
+The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d)), and ((f)) of Fig.
 2.
@@
 By Lemma 5 of [27], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12).
-In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$.
-The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$.
+We can assume that the coordinates $(x,y)$ are centered at the $E_{7}$ point.
+In these coordinates, the monomials of degree $\leq 2$ of the polynomial $f$ vanish, and the homogeneous part of degree three is proportional to the third degree of a linear function that also divides the degree four homogeneous part of $f$.
+Thus, the degree three homogeneous part has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$.
+The critical value of the polynomial at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$.
 This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1.
@@
 The deduction of this proof from Lemma 17 essentially repeats the proof of Proposition 9 with the following simplification.
-In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [21], § 4.5).
+In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [27], § 4.5).
 Therefore, the statement of Proposition 16 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$.
@@
 The polynomials obtained from them by the involution (13) and representing other three connected components of the discriminant complement are also invariant under this action.
-Picture $(d)$ (respectively, $(e)$) is realized by the polynomial
-(15) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),$
-$\alpha\in(0,0.003)$, with positive (respectively, negative) $t$.
+Picture $(d)$ (respectively, $(e)$) is realized by the polynomials
+(15) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),\quad\alpha\in(0,0.003),$
+with positive (respectively, negative) $t$.
 These two pictures are topologically equivalent.
@@
 There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$.
-These eight polynomials form an orbit of the natural action of the group $G_{0}$.
+These eight polynomials form an orbit of the natural action of the group $G_{0}$ on $\Phi_{2}$.
 Lemma 23.
@@
 […]
-The rest of the proof of Proposition 20 repeats the proof of Proposition 3 of [21] in § 5.4 of [21]. $\Box$
+The rest of the proof of Proposition 21 repeats the proof of Proposition 3 in § 5.4 of [27]. $\Box$
 Proof of Theorem 19. Pictures ($a$) and ($b$) of Fig. 3 are realized, respectively, by the polynomials (22) and (23) with arbitrary $t>0$.
@@
 These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [30].
-All the resulting polynomials (2630) are invariant under the reflection in the axis $y=0$.
+All the resulting polynomials (26)–(30) are invariant under the reflection in the axis $y=0$.
 Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement.
@@
 Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively.
-The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consists of 10 elements.
+The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consist of ten elements.
 Proposition 26 (see [29]).
-If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
+If two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are associated with the same virtual function, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
 (33) $f(x,y)\leftrightarrow f(x,-y)\ .$
-In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components. $\Box$
-In the following pictures, the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
+In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components of the discriminant complements in $\Omega_{*}$. $\Box$
+In the following pictures, it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
 3.2. Components for $J_{10}^{1}$ singularities
@@
 There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$.
-These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig.
+These virtual components are realized by polynomials, the topological shapes of whose zero-level sets are shown on Fig.
 4.
@@
 There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$.
-Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig.
+Each of the virtual components realized by polynomials with topological types shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig.
 4 is associated with a single real connected component.
@@
 ---++++++$(m)$
-Figure 5. Perturbations of $J_{10}^{3}$
+Figure 5. Perturbations of $J_{10}^{3}$ singularities
 Proposition 30 (see [24]).
@@
 There are exactly 33 connected components of the complement of the discriminant variety of the versal deformation ((32)). Each virtual component represented by one of ten pictures
-(35) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
+(35) (g),(h),(h~),(i),(i~),(j),(j~),(k),(k~),(l)(g),\ (h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (l)
 […]
@@
 $\Box$
-4. The cases $P_{8}$
+4. The case of $P_{8}$ singularities
 4.1.
@@
 (41) $x^{3}+y^{3}+z^{3}-3Axyz,\quad A\neq 1,$
-with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$.
-The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected components in ${\mathbb{R}P}^{2}$.
-Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to the plane with a single singular point at the origin.
-The zero set of a polynomial of class $P_{8}^{2}$ consists of such a surface and a surface that is homeomorphic to the standard quadratic cone.
-It follows easily from Morse theory that the closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
+with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5].
+The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$.
+Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin.
+The zero set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
 The standard versal deformations of $P_{8}^{1}$ and $P_{8}^{2}$ singularities consist of the polynomials
@@
 Definition 8.
-A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
-Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$.
+A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
+Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$, respectively.
 4.2. The homological invariant of real and virtual components
-Let $f_{0}$ be a function singularity of class $P_{8}^{1}$ or $P_{8}^{2}$ having the form (41).
-Let $D$ be a small neighborhood of the point $\{f_{0}\}$ in the space $\Xi_{1}$ or $\Xi_{2}$.
-Let $B$ be a sufficiently large (compared to $D$) closed ball in ${\mathbb{R}}^{3}$, and let $f_{\lambda}$, $\lambda\in D$, be a small perturbation of $f_{0}$.
-Define $T_{-}(\lambda)$ as a real number slightly lower than the minimum of 0 and the lowest real critical value of $f_{\lambda}$.
-Similarly, define $T_{+}$ as a number slightly higher than the maximum of 0 and the highest real critical value of $f_{\lambda}$.
-For any $\lambda\in D$, define three sets: $W_{\lambda,0}=f_{\lambda}^{-1}((-\infty,0])\cap B$, $W_{\lambda,-}=f_{\lambda}^{-1}((-\infty,T_{-}])\cap B,$ and $W_{\lambda,+}=f_{\lambda}^{-1}((-\infty,T_{+}])\cap B$.
-The sets $W_{\lambda,-}$ and $B\setminus W_{\lambda,+}$ form then a trivializable fiber bundle over $D$.
-These sets contain their intersections with $\partial B$ as deformation retracts.
-In the case $P_{8}^{1}$ these intersections are just the discs.
-In the case $P_{8}^{2}$ each of them homeomorphic to the disjoint union of a disc and an annulus.
-In particular, the reduced homology groups $\tilde{H}_{i}(W_{\lambda,-})$ and $\tilde{H}_{i}(B\setminus W_{\lambda,+})$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
+Let $f$ be a polynomial in the class $\Xi_{1}$ or $\Xi_{2}$.
+Let $T_{-}$ (respectively, $T_{+}$) be an arbitrary negative (respectively, positive) constant which is lower (respectively, higher) than all critical values of $f$.
+Consider three subsets in ${\mathbb{R}}^{3}$: $W_{0}=f^{-1}((-\infty,0])$, $W_{-}=f^{-1}((-\infty,T_{-}]),$ and $W_{+}=f^{-1}((-\infty,T_{+}])$.
+If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$) then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle).
+In particular, the reduced homology groups $\tilde{H}_{i}(W_{-})$ for $f\in\Xi_{2}$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
 Definition 9.
-The homology index $\mbox{HI}(f_{\lambda})$ of a non-discriminant function $f_{\lambda}$, $\lambda\in D$ is the triple of Betti numbers $b_{0},b_{1},b_{2}$ of the relative homology group $H_{*}(W_{\lambda,0},W_{\lambda,-})$.
+The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f\in\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
 It is convenient to consider this homology group in the context of the exact homological sequence of the triple
-(43) $(W_{\lambda,+},W_{\lambda,0},W_{\lambda,-}).$
-[…]
+(43) $(W_{+},W_{0},W_{-}).$
+The groups $H_{*}(W_{+},W_{-})$ participating in this exact sequence are the same for all $f$ from the same space $\Xi_{i}$, because the pairs of sets $(W_{+},W_{-})$ corresponding to different functions $f$ form a trivializable fiber bundle over these spaces.
+In the case $\Xi_{1}$ these groups are trivial.
+For $\Xi_{2}$, they are equal to $\mathbb{Z}$} for $i=1$ or $2$ and are trivial for all other $i$.
 Clearly, the invariant Ind is equal to the alternating sum of three elements of the homology index.
 4.3. The symmetry group and the relation between the real and virtual components
-The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following exceptions: the frames of vanishing cycles in the 2-homology groups of zero-level sets of polynomials consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values.
-We also have the following analog of Proposition 2
+The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values.
+We also have the following analog of Proposition 3.
 Proposition 33 (see [28], Corollary 8).
-1. Any virtual Morse function of the type $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
+1. Any virtual function of the class $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
 2.
-If two virtual Morse functions of type $P_{8}^{1}$ (respectively, $P_{8}^{2}$) belong to the same virtual component, then they are associated with two generic Morse polynomials of class $\Xi_{1}$ (respectively, $\Xi_{2}$) that are isotopic to each other within the set of Morse polynomials.
+If two virtual functions of the class $\Xi_{1}$ ((respectively, $\Xi_{2})$ belong to the same virtual component, then they are associated with two generic polynomials of class $\Xi_{1}$ ((respectively, $\Xi_{2})$ that belong to the same connected component of the space of generic polynomials.
 $\Box$
@@
 The first transformation is an involution, and the second is of order three.
-In particular, this group can act on the set of the connected components of the discriminant complement in the spaces $\Xi_{i}$.
+In particular, this group can act on the set of the connected components of the discriminant complements in the spaces $\Xi_{i}$.
 The transformations (44) preserve all topological characteristics of the functions.
-Thus, all connected components in the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
+Therefore, all these connected components within the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
 Additionally, the involution
 (45) $f(x,y,z)\mapsto-f(-x,-y,-z)$
-also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It preserves the invariant Ind and defines involutions on both sets of real and virtual components of these classes.
+also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It defines an involution on both sets of real and virtual components. This involution preserves the invariant Ind.
+If $f$ and $\tilde{f}$ are two functions related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
 Proposition 34.
-If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class that contain $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$.
+If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class containing $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$.
+In particular, in this case the connected components of the discriminant complement containing these polynomials are also mapped to each other by an element of the group $S(3)$.
 Definition 10.
 For any polynomial $f:{\mathbb{R}}^{3}\to{\mathbb{R}}$ of degree three with principal homogeneous part of the form (41), ${\bf T}(f)$ is the unique polynomial of the form (42) with the same principal homogeneous part, which is obtained from $f$ by a parallel translation in ${\mathbb{R}}^{3}$.
-Namely, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
+Specifically, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
 Lemma 35.
@@
 (46) ${\bf T}\left(x^{3}+y^{3}+z^{3}+6xyz+(x+y-2z)^{2}\right)$
-and two other polynomials obtained from it by permuting the coordinates.
+and two other polynomials obtained from this one by permuting the coordinates.
 Proof.
 According to Lemma 35 of [28], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$.
-Three of these polynomials (mentioned in formula (26) of [22]) are locally equivalent to $\xi^{3}-\eta^{2}-\zeta^{2}+\frac{64}{3}$ at the $A_{2}$ point.
+Three of these polynomials (mentioned in formula (26) of [28]) have signature (0,2) at the $A_{2}$ point.
 The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$.
@@
 […]
-Proof. This follows from statement 2 of Lemma 44 of [22] (the coefficients $\varkappa$ and $c$ in formula (45) of [22] are fixed by our conditions on the critical values). $\Box$
-Proof of Proposition 33. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 34 (respectively, 35) repeats the deduction of Proposition 8 from Lemma 9. $\Box$
+Proof.
+This lemma follows almost immediately from statement 2 of Lemma 44 in [28] (the coefficients $\varkappa$ and $c$ in formula (45) of [28] are fixed by our conditions on the critical values).
+$\Box$
+Proof of Proposition 34.
+The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 35 (respectively, 36) repeats the deduction of Proposition 9 from Lemma 10.
+$\Box$
 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
@@
 Theorem 38.
-Each virtual component of type $P_{8}^{1}$ is associated with only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$.
-So, there are exactly seven such connected components.
-Proof. The invariant Ind of the polynomials
+Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$.
+In particular, there are exactly seven such connected components.
+Proof. The invariant Ind of the polynomial
 (48) $x^{3}+y^{3}+z^{3}-3(x+y+z)+C$
@@
 These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates.
-Therefore, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
-Two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2.
+According to Proposition 34, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
+The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2.
 For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$.
@@
 They are related to each other by the involution (45).
-Hence, by the last statement of Proposition 36, they realize two different virtual components.
+Therefore, according to the final statement of Proposition 37, they realize two distinct virtual components.
 Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components.
+Thus, Theorem 38 is proved for all virtual components with Ind equal to 1,0,2,1,0,-2, and 3-3.
+Let us prove it for the only remaining virtual component with $\mbox{Ind}=-1$.
 […]
@@
 (53) $\xi^{4}-\eta^{2}-\zeta^{2}$
-in some local coordinates, and a Morse critical point with Morse index $1$ and negative critical value.
+in some local coordinates, and a Morse critical point with Morse index $1$ and a negative critical value.
 Proof.
 1.
-At each critical point of the polynomial (51), the sum of its three partial derivatives is equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain.
+The sum of the three partial derivatives of the polynomial (51) is identically equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain.
 In coordinates $l,p_{l},q_{l}$, this polynomial has the form
@@
 where $\psi_{2}$ and $\psi_{1}$ are some homogeneous polynomials of degrees 2 and 1, respectively.
-The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [2], §12, this polynomial has the normal form $Z^{2}+X^{3}+XY^{2}$ of $D_{4}^{+}$ class in some local coordinates.
+The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [3], §12, this polynomial can be reduced to the normal form $\tilde{z}^{2}+\tilde{x}^{3}+\tilde{x}\tilde{y}^{2}$ of $D_{4}^{+}$ class in some local coordinates $\tilde{x},\tilde{y}$, and $\tilde{z}$.
 2.
@@
 Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
-The substitution $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2},P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial to the polynomial with the lowest quasihomogeneous part
+Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest quasihomogeneous part
 $\tilde{l}^{2}-\varepsilon P_{l}^{2}+\left(\frac{\alpha^{2}}{4}+\frac{1}{4\varepsilon}\right)q_{l}^{4}\ .$
 Again, using the techniques of §12 of [3], this function can be reduced to the standard form (53) of $A_{3}$ singularities via a local change of coordinates.
-If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not bring four non-real critical points of (51) back to the real domain.
-On the other hand, by the index considerations the obtained polynomial $F(l)$ has one Morse critical point in addition to the $A_{3}$ point at the origin.
-This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l$.
-It is easy to calculate that for $l\equiv x$ it is equal to 1.
-Our critical point of class $A_{3}$ is topologically equivalent to a Morse critical point with Morse index 2 (and can be perturbed to such a point by an arbitrarily small change of the polynomial $F(l)$).
-The Morse-like complex of the polynomial $F(l)$ is acyclic, therefore the critical value of this $A_{3}$ type critical point should be higher than that of the only other critical point of Morse index 1.
+If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not return the non-real critical points of (51) to the real domain.
+By the index considerations, the obtained polynomial $F(l)$ has one real Morse critical point in addition to the $A_{3}$ point at the origin.
+This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l\in L$.
+It is easy to calculate that for $l\equiv x-y$ it is equal to 1.
+The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points and the local maximum at the $A_{3}$ point.
+Therefore the critical value at its Morse critical point is negative.
 $\Box$
@@
 The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it.
-The set of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$.
-Thus the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement.
+The union of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$.
+Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement.
 $\Box$
-4.5. One-dimensional cohomology of discriminant complements of $P_{8}^{1}$ singularities
+4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
 Theorem 40.
-The first cohomology group of the component of the discriminant complement of the $P_{8}^{1}$ singularities that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
+The first cohomology group of the connected component of the discriminant complement in $\Xi_{1}$ that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
 Proof.
-The zero-level set of the polynomials in this component is homeomorphic to the plane with a handle.
+The zero-level sets of the polynomials in this component are homeomorphic to the plane with a handle.
 Indeed, the topology of these sets remains constant along the component.
-For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to the plane) by passing through a Morse critical point of Morse index 1 once.
-In particular, the first homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$.
-These homology groups form a local system over the component, in particular the fundamental group of the component acts on them by monodromy.
-For each $l\in L$, one basic homology class of the zero-level set of $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$.
+For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
+Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$.
+These homology groups form a local system over the space of all such polynomials, i.e., over the connected component under consideration.
+In particular, the fundamental group of this component acts on these groups by monodromy.
+For each $l\in L$, one basic homology class of the zero-level set of the polynomial $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$.
 When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement.
 The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation.
+Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component.
 $\Box$
-4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularity
+4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
 Definition 11.
-The invariant Card of a virtual component is the number of its vertices.
-The involution (45) preserves this invariant.
+The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that component.
+The involution (45) of virtual components of classes $\Xi_{1}$ and $\Xi_{2}$ preserves this invariant.
 Table 3. Statistics for $P_{8}^{2}$ singularity
@@
 Proposition 41 (see [24], Proposition 8).
-1. There are exactly 9174 distinct virtual functions of $P_{8}^{2}$ singularity class.
-2. These virtual functions are distributed over exactly fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and third rows of Table 3.
-3. All pairs of different virtual components of $P_{8}^{2}$ singularity with the same invariant Card value can be realized by generic polynomials connected by involution ((45)).
+1. There are exactly 9174 distinct virtual functions of class $\Xi_{2}$.
+2. These virtual functions are distributed among fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and the third rows of Table 3.
+3.
+All pairs of different virtual components of class $\Xi_{2}$ with the same Card value ((i.e., the pairs of components 3 and 4, 6 and 7, 9 and 10, 11 and 12, and 14 and 15 in Table 3)) can be realized by generic polynomials connected by involution ((45)).
 Theorem 42.
-Each of the ten virtual components of the formal graph of type $P_{8}^{2}$ that are mentioned in columns 1, 3, 4, 6, 7, 11, 12, 13, 14, and 15 of Table 3 is associated with only one real connected component of the discriminant complement.
+Each virtual component of the formal graph of type $P_{8}^{2}$ is associated with only one real connected component of the discriminant complement.
+In particular, there are exactly fifteen such connected components.
 Proof of this theorem takes the rest of this section.
-All of the polynomials given in the next formulas (55), (56), (57), and (58) are invariant under the $S(3)$ action.
-Thus, for any polynomial $f$ of this list and any noncritical value $C$ of $f$, the set-theoretical invariant of the polynomial fCf-C is a virtual component associated with only one real component of the discriminant complement.
-We will show that these virtual components, or those obtained from them via the involution (45), include all virtual components with Card equal to 258, 336, 1648, 262, and 94.
-The proof of Theorem 41 for the components with $\mbox{Card}=60$ is more implicit, see § 43.
-Components with $\mbox{\rm Ind}=0$ and $\mbox{\rm Card}=1648$.
-For each polynomial $f$ of class $\Xi_{2}$, the polynomials $f+C$ and fCf-C with sufficiently large $C$ realize two different virtual components with $\mbox{Ind}=0$.
-Indeed, the homology index of one of these components is $(000)$, and the other’s is $(011)$.
-In particular, the $S(3)$-invariant polynomials
-(55) $x^{3}+y^{3}+z^{3}-4xyz+C\qquad\mbox{and}\qquad x^{3}+y^{3}+z^{3}-4xyz-C$
-represent these two components for any $C>0$. These polynomials are not Morse, but any their sufficiently small generic perturbations are associated with virtual functions with $\mbox{Card}=1648$.
-Components with $\mbox{\rm Card}=262$. The polynomial
+All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$ action.
+Thus, if $f$ is any of these polynomials and $C$ is a noncritical value of $f$, then Proposition 34 implies that the virtual component associated with the polynomial fCf-C is associated with only one real component of the discriminant complement.
+We will show that all virtual components with Card equal to 258, 336, 1648, 262, or 94 can be realized in this way.
+Hence, the assertion of Theorem 42 is true for these components.
+Components with $\mbox{\rm Card}=1648$.
+For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$.
+The homology indices of these components are equal to $(000)$ and $(011)$.
+For example let $f$ be the $S(3)$-invariant polynomial
+(55) $x^{3}+y^{3}+z^{3}-4xyz.$
+The polynomial fCf-C with arbitrary $C\neq 0$ then represents one or the other of these two components depending on the sign of $C$.
+These polynomials are not Morse, but all their sufficiently small generic perturbations have $\mbox{Card}=1648$.
+Component with $\mbox{\rm Card}=262$. The polynomial
 (56) $x^{3}+y^{3}+z^{3}-4xyz+(x+y+z)$
 has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2.
-Thus, it has invariant values $\mbox{Ind}=1$ and $\mbox{HI}=(00\bar{1})$.
+Thus, it has $\mbox{Ind}=1$ and homology index $(001)$.
 This polynomial is invariant under the involution (45).
-According to the last statement of Proposition 40, it cannot represent a component of the discriminant complement with $\mbox{Card}=94$.
-Therefore, it belongs to the component with $\mbox{Card}=262$.
+Therefore, according to the last statement of Proposition 41, the associated virtual component cannot have $\mbox{Card}=94$.
+Thus, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
 Components with $\mbox{\rm Card}=94$ and $336$. The polynomial
-(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz-\frac{3}{2}(x^{2}+y^{2}+z^{2})$
-has eight real Morse critical points: one local maximum with critical value 0, three points of Morse index $2$ with value $\frac{11}{49}$, three points of index $1$ with value $\frac{1}{2}$, and one more point of index $2$ with value $\frac{3}{2}$.
-Therefore, the index $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{3}{2}$, to 1 for $C\in\left(-\frac{3}{2},-\frac{1}{2}\right),$ to 2-2 for $C\in\left(-\frac{1}{2},-\frac{11}{49}\right),$ to $1$ for $C\in\left(-\frac{11}{49},0\right),$ and to 0 for $C>0$.
-The group $H_{*}(W_{\lambda,+},W_{\lambda,0})$ corresponding to the polynomials fCf-C with $C\in\left(-\frac{11}{49},0\right)$ is isomorphic to $\mathbb{Z}$} in dimension three and is trivial in all other dimensions.
-Therefore, by the exact sequence of the triple (43), the homology index $\mbox{HI}(f+\frac{1}{7})$ is equal to $(0,1,2)$.
-The involution (45) moves this polynomial to that with $\mbox{HI}=(1,0,0)$.
-In particular, these two polynomials represent two different components, and the associated virtual components are also different.
-According to Table 3, these are the components with $\mbox{Card}=94$.
-The homology index $\mbox{HI}(f-C)$ with $C\in\left(-\frac{3}{2},-\frac{1}{2}\right)$ is $(0,0,1),$ because the corresponding group $H_{*}(W_{\lambda,0},W_{\lambda,-})$ is generated by the class of one critical point of Morse index $2$.
-Therefore, it is again the component with $\mbox{Card}=262$.
-[…]
-Components with $\mbox{\rm Card}=258,$ $336,$ and $94$. The polynomial
-(58) $f=x^{3}+y^{3}+z^{3}-12xyz-\frac{1}{2}(x^{2}+y^{2}+z^{2})$
-has eight real Morse critical points: a local maximum with critical value 0, three points with Morse index 2 and critical value $\frac{11}{3969}$, one point more with Morse index 2 and critical value $\frac{1}{162}$, and three points with index 1 and critical value $\frac{1}{54}$.
-Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{1}{54}$, to 3-3 for $C\in\left(-\frac{1}{54},-\frac{1}{162}\right)$, to 2-2 for $C\in\left(-\frac{1}{162},-\frac{11}{3969}\right)$, to $1$ for $C\in\left(-\frac{11}{3969},0\right)$, and to 0 for $C>0$.
+(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz+3(x^{2}+y^{2}+z^{2})$
+has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one point with Morse index $1$ and critical value $12$.
+Therefore, the invariant $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{49}\right),$ to 2-2 for $C\in\left(\frac{88}{49},4\right),$ to 1 for $C\in\left(4,12\right),$ and to 0 for $C>12$.
+The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions.
+Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$.
+The involution (45) moves this polynomial to that with $\mbox{HI}=(012)$.
+In particular, these two polynomials represent two different components.
+They are $S(3)$-invariant, therefore by Proposition 34, the associated virtual components are also different.
+According to Proposition 41 and Table 3, these are two components with $\mbox{Card}=94$.
+According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has $\mbox{HI}=(031)$.
+Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants.
+The only possibility for this is two components with $\mbox{Card}=336.$
+Component with $\mbox{\rm Card}=258$. The polynomial
+(58) $f=x^{3}+y^{3}+z^{3}-12xyz+3(x^{2}+y^{2}+z^{2})$
+has eight real Morse critical points: one local minimum with critical value 0; three critical points with Morse index 1 and critical value $\frac{88}{147}$; one critical point with Morse index 1 and critical value $\frac{4}{3}$; and three critical points with Morse index 2 and critical value $4$.
+Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{147}\right)$, to 2-2 for $C\in\left(\frac{88}{147},\frac{4}{3}\right)$, to 3-3 for $C\in\left(\frac{4}{3},4\right)$, and to 0 for $C>4$.
 Lemma 43.
@@
 Proof.
-By Proposition 40 of [22], the intersection matrix of vanishing cycles of the canonical basis for all generic polynomials of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Dynkin diagram shown in
+By Proposition 40 of [28], the intersection matrix of the canonical basis of vanishing cycles for any generic polynomial of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Coxeter-Dynkin graph shown in
 Figure 7.
@@
 7.
-This diagram does not contain edges of multiplicity 2-2.
+This graph does not contain edges of multiplicities other than 1 and 1-1.
 Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2.
-Therefore, the virtual component associated with our polynomial is the unique remaining virtual component with $\mbox{Ind}=-3$.
+Therefore, the virtual component associated with our polynomial is the only remaining virtual component with $\mbox{Ind}=-3$.
 $\Box$
 Remark 44.
-It is easy to see that the polynomials (58) with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
+It is easy to check that the polynomials fCf-C with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of some components with $\mbox{Card}=336$ and $94$, respectively.
 Components with $\mbox{Card}=60$.
+Proposition 45.
+There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ that realizes a virtual component with $\mbox{\rm Card}=60.$
+Proof.
+According to Proposition 41, two virtual components exist with $\mbox{Card}=60$.
+Our program shows that one of these components contains a virtual function that models a generic polynomial of class $\Xi_{2}$ such that
+1) all of its critical points are real,
+2) the three highest critical values are positive and the other five critical values are negative,
+12536748
+Figure 8. Coxeter-Dynkin graph for (59
+3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is as follows:
 […]
-Theorem 44.
-There exists a $S(3)$-invariant polynomial $f\in\Xi_{2}$ realizing a virtual component with $\mbox{\rm Card}=60.$
-Proof. One of the two virtual components, each of which has $\mbox{Card}=60$, contains a virtual function that models a generic polynomial $f_{0}$ of class $\Xi_{2}$ such that
-1) all of its critical points are real,
-2) the three highest critical values are positive and the other five values are negative,
-3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is shown in (59),
-4) the Morse indices of the points 1, 2, 3, and 5 are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
-According to Proposition 32, there is a real polynomial $f_{0}\in\Xi_{2}$ with these data.
-According to Proposition 11 of [22], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), all odd Morse indices are equal to 1, and all even Morse indices are equal to 2.
-Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by their increase, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
-Using the Lyashko–Looijenga map (in the form of [10]) we can lift the path $i$ uniquely into a path $I:[0,1]\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$.
+and hence its Coxeter-Dynkin graph is shown in Fig. 8,
+4) the Morse indices of the points number 1, 2, 3, and 5 with respect to this order are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
+According to Proposition 33, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data.
+According to Proposition 41 of [28], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2.
+Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
+Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$.
 As in the proof of Theorem 7 of [28], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined.
-Indeed, according to Proposition 2 of [10], this is guaranteed by the fact that all quadratic forms corresponding to the limit critical values are elliptic.
-The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with value $c_{-}$ and three critical points with value $c_{+}$.
+Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic.
+(These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
 Lemma 46.
@@
 Let $\varepsilon>0$ be a very small number.
-For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$, respectively.
-Different permutations $\sigma$ of critical points define six different points $\tilde{f}_{\sigma}\in U$.
-All of these points are associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
-As in the proof of Proposition 33 (repeating the proof of Proposition 8), the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$.
-Therefore, each such element maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point from the same neighborhood.
+For any permutation $\sigma\equiv\binom{1\ 2\ 3}{i_{1}\,i_{2}\,i_{3}}$ of critical points of $I(1)$ with critical value $c_{-}$, there is a unique point $\tilde{f}_{\sigma}\in U$, whose critical values at the critical points with numbers $i_{1},4,5,$ and $5+i_{1}$ are the same as those of the polynomial $I(1)$, the values at the points $i_{2}$ and $5+i_{2}$ are equal respectively to $c_{-}+\varepsilon$ and $c_{+}+\varepsilon$, and the values at the points $i_{3}$ and $5+i_{3}$ are equal respectively to $c_{-}-\varepsilon$ and $c_{+}-\varepsilon$.
+Different permutations $\sigma$ define six different points $\tilde{f}_{\sigma}\in U$.
+These points are all associated with the same virtual function, because any permutation $\sigma$ of the first three rows and columns in (59) and the simultaneous permutation of the last three rows and columns by the same rule $\sigma$ preserves the matrix.
+As in the proof of Proposition 34, the six points $\tilde{f}_{\sigma}\in U$ are all mapped to each other by the action of elements of the group $S(3)$.
+The order of this group is also six, therefore each element of the group only permutes these six points.
+In particular, it maps a point from an arbitrarily small neighborhood $U$ of the polynomial $I(1)$ to another point in the same neighborhood.
 By continuity, it also maps the polynomial $I(1)$ to itself.
 $\Box$
-By the construction, the Card invariant of the virtual function associated with arbitrarily small perturbations of the polynomial $I(1)$ is equal to 60. This proves Theorem 44. $\Box$
+Due to the construction of the polynomial $I(1)$, the Card invariant of arbitrarily small generic perturbations of this polynomial equal to that of the initial polynomial $f_{0}$, that is, it is equal to 60.
+This proves Proposition 45.
+$\Box$
 The involution (45) moves the polynomial $I(1)$ to a polynomial in the other component of the discriminant complement with $\mbox{Card}=60$.
 Thus, this component is also represented by an $S(3)$-invariant polynomial.
-This concludes the proof of Theorem 41.
+This proves the assertion of Theorem 42 for both virtual components with $\mbox{Card}=60$.
+Component with $\mbox{Card}=1216$. Denote by $\hat{f}$ the polynomial
+(60) $x^{3}+y^{3}+z^{3}-6xyz+3(x-y)^{2}$
+This polynomial has a critical point of class $D_{4}^{-}$ with critical value 0 at the origin, two Morse critical points with Morse index 2 and a common critical value $\omega_{1}\approx 3.14$, and two non-real critical points with a common critical value $\omega_{2}\approx-11.63.$
+Lemma 47.
+$\mbox{Card}(\hat{f}-1)=1216$.
+Proof.
+Due to the versality of the space $\Xi_{2}$, for any sufficiently small $\varepsilon>0$ we can slightly perturb the polynomial $\hat{f}$ within the space $\Xi_{2}$ in such a way that the resulting polynomial $f_{\varepsilon}$ satisfies the following conditions:
+  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $2\varepsilon^{2}$,
+  • the critical values at its two critical points with Morse index 2 are equal to $\omega_{1}$ and $\omega_{1}+\varepsilon^{2}$,
+  • the critical values at the non-real critical points are equal to $\omega_{2}+i\varepsilon^{2}$ and $\omega_{2}-i\varepsilon^{2}$.
+Denote the polynomial $f_{\varepsilon}-1$ by $\bar{f}$.
+It belongs to the same component of the discriminant complement as the polynomial $\hat{f}-1$.
+The canonical system of paths in ${\mathbb{C}}^{1}$ defining the frame of vanishing cycles in $H_{2}(\bar{f}_{\varepsilon}^{-1}(0))$ then looks as shown in Fig.
+9 (left), where the numbers 1–8 denote the order of basic vanishing cycles corresponding to the critical values.
+[…]
+012341-1785116
+Figure 9. System of paths for $\bar{f}$ (left) and $f!$ (right)
+The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program.
+Based on the information on the polynomial $\bar{f}$, let us try to find it in this list.
+According to § 5.2 of [28] there are exactly 1897 virtual functions of class $\Xi_{2}$ with six real critical points.
+Only 140 of these virtual functions are associated with real polynomials that have a minimum point.
+Each of these haves, in addition to the minimum, three critical points with Morse index 1 and two points with Morse index 2.
+Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points.
+Of these 20 virtual functions, only 12 have these positive critical values at the critical points with Morse index 2.
+Thus, the virtual function associated with $\bar{f}$ is among these 12.
+Upon a special request, our program printed these 12 virtual functions.
+It turns out that for all of them the intersection index $\langle\Delta_{7},\Delta_{8}\rangle$ of two cycles vanishing at non-real points is equal to 1.
+In particular, this is true for the real polynomial $\bar{f}$.
+Using the Lyashko–Looijenga map in the form of [12], (see the proof of Theorem 7 in [28]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig.
+9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point.
+Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$.
+With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta$.
+Denote this polynomial by $f!$, see Fig.
+9 (right).
+Its critical point corresponding to the critical value $\varepsilon-1$ and marked with 1 in this figure is a minimum point.
+The critical points with values marked with 2, 3, and 4 have Morse index 1, and the critical points marked with 7 and 8 have Morse index 2.
+According to Lemmas V.3.2 and V.3.3 of [22], the newborn critical point with critical value $1-\zeta$ (marked with 5 in the picture) has an odd Morse index, and the point with critical value $1+\zeta$ has an even Morse index.
+According to Proposition 40 of [28], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
+Proposition 40 of [28] implies that all real polynomials of class $\Xi_{2}$ with eight critical points, one of which is a minimum, have the Coxeter-Dynkin graph shown in Fig. 7, in such a way that
+  • the lowest vertex in this figure corresponds to the minimum point, the vertices in the middle row correspond to critical points with Morse index 1, and the upper row vertices correspond to points with Morse index 2;
+  • for each pair of vertices of the graph that are connected by an edge, the critical value of the critical point corresponding to the higher vertex in this figure is greater than the critical value of the critical point corresponding to the lower vertex.
+All vertices of this graph correspond to some critical points (and their critical values) of the polynomial $f!$.
+According to the previous conditions, the lowest vertex in Fig.
+7 corresponds to the value marked 1 in Fig.
+9 (right), the vertices in the middle row correspond to the values marked 2, 3, 4, and 5, and the vertices in the upper row correspond to the values marked 6, 7, and 8.
+The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5 of Fig.
+9 (right).
+Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [22].
+Therefore, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig.
+7 and one vertex of valence 3 in the middle row.
+The Coxeter-Dynkin graph of Fig. 7, together with
+  • (a)
+    the numbering of its vertices according to the order of the corresponding critical values,
+  • (b)
+    the number of negative critical values, and
+  • (c)
+    the Morse indices of the critical points corresponding to its vertices,
+defines uniquely a virtual function, see Remark 1.
+All virtual functions that can be defined by the graph of Fig.
+7 and information (a), (b), and (c) satisfying the above restrictions on numbering and Morse indices, belong to the same virtual component.
+Indeed, they can all be obtained from each other by transpositions of orders of vertices that are not connected by edges of the graph and corresponding to the critical values of the same sign.
+Such a transposition is a flip of virtual functions which corresponds to the elementary function surgery when one critical value overtakes another critical value at a distant critical point.
+Our program, when applied to one of the virtual functions from this set, says that this virtual component has 1216 vertices.
+$\Box$
+Let $L\sim S^{1}$ again be the set of all linear functions ${\mathbb{R}}^{3}\to{\mathbb{R}}$ of unit norm that vanish at the vector $(1,1,1)$.
+Lemma 48.
+For each $l\in L$, the polynomial
+(61) $F(l)\equiv x^{3}+y^{3}+z^{3}-6xyz+3l^{2}$
+has a critical point of class $D_{4}^{-}$ with zero critical value at the origin. All its other critical values at real critical points are positive.
+Figure 10. $D_{4}^{-}$ singularity and its perturbation
+Proof.
+The restriction of the polynomial (61) to the plane $\{l=0\}$ is a non-degenerate homogeneous cubic polynomial that vanishes on three real lines, see Fig.
+10 (left).
+This implies the first assertion of the lemma analogously to the proof of Lemma 39.
+Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$.
+The restriction of this polynomial to the line connecting 0 and $a$ is a polynomial of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0.
+$\Box$
+Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin.
+These numbers form a lower semi-continuous function on $L$.
+By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$.
+The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant.
+This union is connected and invariant under the action (44) of the group $S(3)$.
+By Lemma 47, its point corresponding to $l\equiv x-y$ belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$.
+Therefore, this entire component is $S(3)$-invariant.
+This concludes the proof of Theorem 42 for the virtual component with $\mbox{Card}=1216$.
+Components with $\mbox{Card}=1318$ and $844$.
+Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other critical points with critical value 0.
+By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$.
+All polynomials $F(l)$ of the form (61) belong to this stratum.
+Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum.
+Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin.
+Therefore, the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
+(62) $x^{2}y-y^{3}+\varkappa_{1}+\varkappa_{2}x+\varkappa_{3}y+\varkappa_{4}y^{2}$
+of the sample polynomial $x^{2}y-y^{3}$ of class $D_{4}^{-}$ in two variables with parameters $\varkappa_{i}$.
+Denote the fiber over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
+  • (a)
+    $f$ has four real critical points near the origin obtained by splitting the critical point of the polynomial $F(l)$: specifically, one minimum and three points with Morse index 1,
+  • (b)
+    $\mbox{Card}(f)=844$.
+Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) and condition (b) replaced by $\mbox{Card}(f)=1318$.
+Lemma 49.
+For each $l\in L$, the subset $V(l)$ is contractible, and the subset $W(l)$ consists of two contractible connected components.
+Preliminary remarks.
+By the versality property, the discriminant is equisingular along the smooth stratum $\{D_{4}^{-}\}$.
+Therefore, it is sufficient to prove Lemma 49 for only one particular $l\in L$, say for $l\equiv x-y$.
+The union of subsets $V(l)$ over all $l\in L$ is $S(3)$-invariant.
+Lemma 49 implies that it is also connected, and therefore the corresponding virtual component with $\mbox{Card}=844$ is associated with only one real connected component of the discriminant complement.
+Applying the involution (45) proves the assertion of Theorem 42 also for the other virtual component with $\mbox{Card}=844.$ The union of subsets $W(l)$ is also $S(3)$-invariant.
+It is fibered over the circle $L/{{\mathbb{Z}}_{2}}$ with the fiber consisting of two contractible components.
+In Lemma 50, we will see that the monodromy of this fibration permutes these components.
+Therefore, their union over all $l\in L$ is connected, and we again have only one real connected component.
+To prove Lemma 49, we will describe detail the geometry of the discriminant in the parameter space of the versal deformation (62) in detail.
+The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$.
+In the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig.
+11 (left), see [3], § 21.3.
+abcabcacbacbbacbaccabcabbcabcacbacba
+ccbbaa
+Figure 11. Caustic and Maxwell set for $D_{4}^{-}$ singularity
+This set divides the space ${\mathbb{R}}^{3}$ into three parts.
+The larger part consists of polynomials with only two real critical points.
+Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one point that is either a minimum or a maximum, depending on the pyramid.
+We are only interested in the polynomials (62) with a minimum point.
+For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values consists of five intervals.
+Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
+(63) $x^{2}y-y^{3}+\varkappa_{2}x+\varkappa_{4}y+\varkappa_{4}y^{2},$
+i.e., it can belong to one of the five intervals into which the critical values divide the line.
+For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal.
+The set of these non-generic polynomials (63) divides the pyramid into six parts.
+The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig.
+11 (center).
+The non-generic polynomials form three segments in it.
+In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints.
+The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig.
+10 (right).
+These polynomials, representing various sections of the pyramid, sweep out its central axis.
+An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
+Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$.
+This marking of critical points can be uniquely extended by continuity to all other polynomials within the pyramid.
+Each of the six cells into which the curvilinear triangle in Fig.
+11 (center) is split by the segments is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
+The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of nine connected domains.
+We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value.
+The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three shares of the pyramid: their intersections with a section of the pyramid are shown in Fig.
+11 (right).
+Each share, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the saddlepoint with that marking is the highest.
+The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding share shown in Fig.
+11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values.
+Clearly, all three of these domains are contractible.
+Proof of Lemma 49.
+For each $l\in L$, the orthogonal slice $U(l)$ of the discriminant stratum $\{D_{4}^{-}\}\subset\Xi_{2}$ at the point $\{F(l)\}$ is a versal deformation of the corresponding singularity of class $D_{4}^{-}$.
+The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant.
+Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph.
+These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
+Consider again the case $l\equiv x-y$.
+Let $f$ be a point in one of these three domains in the slice $U(l)$ at the point $F(l)$.
+We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\zeta$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points.
+Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$.
+For the same reasons as in the aforementioned proof, the corresponding critical points meet at a real critical point of class $A_{2}$.
+Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points.
+We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig.
+7.
+This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity.
+The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices.
+Only one of these three critical points has a positive critical value.
+Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex.
+This choice explicitly determines the virtual component associated with the obtained polynomial (and?
+consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
+Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant.
+It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig.
+7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases.
+$\Box$
+As $l$ moves along half of the circle $L$, the polynomial $F(l)$ and the corresponding slice $U(l)$ returns to themselves.
+The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this half-circle is a connected set that is $S(3)$-invariant.
+All of these domains consist of polynomials with $\mbox{Card}=844$.
+This proves the assertion of Theorem 42 for the virtual components with $\mbox{Card}=844$.
+Lemma 50.
+The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ of all polynomials $F(l)$ permutes the two components of the sets $W(l)$ mentioned in Lemma 49.
+$l=0$
+$l=0$
+Figure 12. The plane $\{l=0\}$ and the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
+Proof. The zero-level set of the polynomial $F(l)$ in the plane $\{l=0\}$ consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set
+(64) $x^{3}+y^{3}+z^{3}-6xyz=0,$
+which is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 (left).
+Consider the local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$ at the origin in ${\mathbb{R}}^{3}$ such that $l\approx\tilde{l}$ up to terms of second degree, and $F(l)$ has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class.
+The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin.
+The curves on this surface that are distinguished by the conditions $\tilde{p}=0$, $\tilde{p}=\tilde{q}$, and $\tilde{p}=-\tilde{q}$ are tangent to the three lines considered in the previous paragraph.
+Let $\hat{F}(l)\in U(l)$ be a perturbation of $F(l)$ that splits the $D_{4}^{-}$ singularity into a minimum point with critical value $\varepsilon$ and three critical points of Morse index 1 with critical value 0.
+In canonical coordinates, the zero-level set of the restriction of $\hat{F}(l)$ to the surface $\{\tilde{l}=0\}$ locally consists of three curves obtained by small deformations of the three curves for the function $F(l)$, and the three critical points occur at their intersections, see Fig.
+10 (right).
+Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the nearby domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at exactly this critical point is positive and the critical values at the other two such points are negative.
+Thus, there is a canonical one-to-one correspondence between these domains in $U(l)$ and the choices of some two out of three lines in the plane $\ker l$ considered in the previous paragraph.
+As $l$ moves along the half-circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute.
+Thus, the monodromy over this half-circle acts in the same way on the set of three related domains in $U(l)$.
+We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 49.
+Therefore, the permuted domains belong to the set $W(l)$.
+$\Box$
+Theorem 42 is thus proved for the virtual component with $\mbox{Card}=1318$.
+Component with $\mbox{Card}=156$.
+Lemma 51.
+Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig.
+12 ((right)).
+The polynomial
+(65) $x^{3}+y^{3}+z^{3}-6xyz+l^{2}$
+then has a singularity of class $D_{5}^{-}$ at the origin in ${\mathbb{R}}^{3}$. It has the form
+(66) $\zeta^{2}+\xi^{2}\eta-\eta^{4}$
+in some local coordinates at this point. The critical values at all other real critical points of ((65)) are positive.
+Proof.
+All singularities of corank 2 that can occur in the deformation (42) of $P_{8}^{2}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [19], [12].
+The restriction of the polynomial ((65)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$.
+According to Corollary 10.4.9 of [21], this singularity is not of class $E_{6}$.
+Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$.
+In the particular case when
+(67) $l\equiv 2x-5y+11z,$
+the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and Y=y3zY=y-3z reduce the polynomial (65) to a polynomial whose lower quasihomogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}Y^{2}z-\frac{9}{4}z^{4}$.
+A dilation of coordinates reduces it to the form (66).
+The set of all planes in ${\mathbb{R}}^{3}$ that are tangent to the same component of the zero-level set of (64) is connected.
+Therefore, the singularity will be of the same class for all other functions (65) corresponding to these planes .
+The positivity of other critical values follows from the same arguments as in the proof of Lemma 48.
+$\Box$
+Lemma 52.
+For any linear function $l$ as in Lemma 51, the set of polynomials with $\mbox{Ind}=-3$ consists of a single connected component in a small neighborhood of the corresponding polynomial ((65)).
+Proof.
+By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}$.
+Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (65).
+Any such slice realizes a versal deformation of the corresponding singularity.
+The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [25].
+For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$: it is the component containing the perturbations shown in Fig.
+13.
+According to Theorem 1.7 of [16], this component is contractible.
+$\Box$
+--$\Rightarrow$
+-$\Rightarrow$
+Figure 13. Perturbation of $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
+Lemma 53.
+All the polynomials in the set considered in Lemma 52 have $\mbox{Card}=156.$
+Proof. We can assume that $l$ is given by (67). Then the set in consideration contains a generic polynomial with the following properties:
+  • it has exactly six real critical points: three with Morse index 1 and three with Morse index 2,
+  • the critical values at all points with Morse index 1 are lower than all critical values at the points with Morse index 2,
+  • the intersection indices of vanishing cycles corresponding to the five lowest real critical values are defined by the matrix
+[…]
+Indeed, this polynomial can be obtained from the perturbation of (65) shown in the middle part of Fig.
+13 by a minor additional perturbation that properly orders the critical values.
+The intersection indices of its critical values are derived using the Gusein-Zade—A’Campo method [10].
+Our program shows that only six different virtual functions of $P_{8}^{2}$ class satisfy these conditions.
+It also shows that all six of these virtual functions appear in the virtual component with $\mbox{Card}=156.$ $\Box$
+The set of all polynomials (65) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant.
+The same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}\}$ stratum at the these points.
+This implies the statement of Theorem 42 for the virtual component with $\mbox{Card}=156$.
 $\Box$
@@
 Wavefronts are the discriminant sets of these families.
-Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, and the wave function is regular as well.
-However, over the discriminant sets of parameters, the level sets containing the integration cycles degenerate and the wave function may be singular.
-The singularity of a wave function in a neighborhood of a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point.
-The type of this asymptotic behavior depends primarily on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle.
-The homology classes depend on the wavefront complement components from which we approach the wavefront point.
+Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well.
+However, over the discriminant sets of parameters, the level sets degenerate and the wave function may become irregular.
+The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point.
+The type of this asymptotic behavior primarily depends on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle.
+These homology classes depend on the connected components of the wavefront complement from which we approach the wavefront point.
 Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
@@
 For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas.
-On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or the stably equivalent functions in n+2k1n+2k-1 variables generally has a different set.
+On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or a stably equivalent function in n+2k1n+2k-1 variables generally has a different set.
 Theorem 54.
@@
 $J_{10}^{1}$ 0 0 0 0
-In comparison with the results of [17], this table replaces the lower estimate $\geq 2$ for $P_{8}^{2}$ with the exact number 3, and replaces some five estimates of the form $\geq k$ with the same numbers $k$.
+Almost all of these local lacunas were described in [23], and it was conjectured that there are no additional local lacunas.
+The above explicit enumeration of connected components of the discriminant complements proves this conjecture for all singularity classes except for $P_{8}^{2}$.
+For $P_{8}^{2}$, we find one new local lacuna, thus replacing the number 2 conjectured in [23] by 3.
 In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig.
@@
 The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure.
+The lacunas for the $X_{9}^{-}$ singularities are obtained from the $X_{9}^{+}$ singularities by multiplying of all functions by 1-1.
 The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig.
@@
 5.
-The two local lacunas of the $P_{8}^{1}$ singularity at odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$.
+The two local lacunas of the $P_{8}^{1}$ singularity with odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$.
 Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$.
-The only one new local lacuna found now is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.
+The only new local lacuna found is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.

### v4

@@
 Abstract.
-We list all connected components of sets of non-discriminant functions within versal deformations of all parabolic function singularities.
-These singularities are the second most common family of singularity classes of smooth functions after simple singularities.
-Thus, we prove (and improve in one particular case) the corresponding conjectures from the previous work [24] with the same title.
-As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs.
-We also show that the complements of the discriminant varieties of $X_{9}^{+}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
+We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most common family of singularity classes of smooth functions after simple singularities).
+Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [25] with the same title.
+As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs (one of which is a new one).
+We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
 These results are applications of a general method for investigating non-singular perturbations of real function singularities.
@@
 [2], [3], [4], [9], [14], [21], [23], [24].
-Many important functions in these theories arise as integral transforms.
+Many functions studied in these theories are defined by integral transforms.
 Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals along the cycles in the zero-level sets of the corresponding functions.
 These integral functions are regular outside the discriminant varieties of the families.
-The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding discriminant function and on the connected component of the complement of the discriminant, from which we approach this point.
+The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant, from which we approach this point.
 Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
 Definition 1 (see [3]).
-A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},a)\to({\mathbb{R}},0)$ at its critical point.
+A point $a\in{\mathbb{R}}^{n}$ is a critical point of a smooth function ${\mathbb{R}}^{n}\to{\mathbb{R}}$ if all first partial derivatives of this function vanish at $a$.
+A function singularity is a germ of a regular analytic function $({\mathbb{R}}^{n},a)\to({\mathbb{R}},0)$ at its critical point with critical value 0.
 Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$ (i.e., they have the same expression in appropriate local coordinate systems centered at $a$ and $b$).
@@
 $f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
+The critical points of analytic functions ${\mathbb{R}}^{n}\to{\mathbb{R}}$ are equivalent if the function singularities obtained from their germs at these points by subtracting the critical values are equivalent.
 An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic.
@@
 For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation.
-The discriminant varieties of all versal deformations of stably equivalent singularities are topologically equivalent, and the connected components of their complements in the parameter spaces are in a one-to-one correspondence.
+The complements of discriminant varieties of all versal deformations of stably equivalent singularities are homotopically equivalent, in particular the connected components of these complements are in a one-to-one correspondence.
 Therefore, for any stable equivalence class of singularities it is sufficient to enumerate these components for only one versal deformation of one function of this class.
-The list of equivalence classes of function singularities begins with simple singularities, which were distinguished by V.
+The list of equivalence classes of function singularities begins with simple singularities, which were identified by V.
 Arnold, see [3].
-All of these singularities are stably equivalent to functions in two variables, and the entire geometric study of their deformations and discriminants reduces to the study of such functions.
+All of these singularities are stably equivalent to singularities of functions in two variables, and the entire geometric study of their deformations and discriminants reduces to studying such functions.
 The components of discriminant complements for simple singularities were listed earlier, see § 1.3.
@@
 The classification of their real forms includes more classes: $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$.
-Singularities of classes $X_{9}$ and $J_{10}$ are stably equivalent to the functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables.
+Singularities of classes $X_{9}$ and $J_{10}$ are stably equivalent to the singularities of functions in two variables, while the $P_{8}$ singularities are stably equivalent to the non-degenerate cubic forms in three variables.
 Any real function singularity in two variables of class $X_{9}^{*}$ or $J_{10}^{*}$, as well as any singularity in three variables of one of classes $P_{8}^{*}$, has the normal form indicated in Table 1 in appropriate local coordinates centered at the critical point.
-Additionally, this table provides the normal forms of simple singularities $A_{k},D_{k},$ and $E_{k}$, into which the parabolic singularities can split after small perturbations.
+Additionally, this table provides the normal forms of the simple singularities, $A_{k},D_{k},$ and $E_{k}$.
 Table 2. Numbers of rigid isotopy and topological classes of non-discriminant perturbations of parabolic singularities
@@
 Clearly, the topological type is an invariant of the connected components of the discriminant complements.
-Our “rigid isotopy” classification of non-discriminant perturbations separating all these components is only slightly stricter than the topological classification, see Table 2.
+Our “rigid isotopy” classification of non-discriminant perturbations, i.e., the classification whose classes are these components, is only slightly stricter than the topological classification, see Table 2.
 Counting the components of the discriminant complements of $P_{8}$ singularities is related to (though is different from) the isotopy classification of nonsingular cubic surfaces in ${\mathbb{R}}^{3}$, which is described in detail in [8].
@@
 1.2. On the method
-In [24], a system of invariants of connected components of the complements of the discriminant varieties of function singularities is described.
-Complete lists of possible values of these invariants for parabolic singularities are given in [24], and all of these values are realized by concrete polynomials.
-In the same paper, we also conjectured that the resulting lists of connected components are complete.
-These lists consist of 7, 15, 7, 14, 52, 11, and 33 components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively.
+In [25], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for parabolic singularities are found, and all of these values are realized by concrete polynomials.
+The remaining problem is to list all components with each value.
+In [25], a conjectural answer to this problem was proposed.
+It provides the lists consisting of 7, 15, 7, 14, 52, 11, and 33 connected components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively.
 Below, we prove that almost all of these lists are indeed complete.
 However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [25] consist of pairs of different connected components.
-Thus, the number 11 above is replaced in Table 2 by 13.
+Thus, the number 11 is replaced in Table 2 by 13.
 These results are summarized in Theorems 6, 14, 19, 29, 32, 38, and 42 below.
@@
 E.
-Looijenga [16] implicitly listed the local connected components of the complements of the discriminants of simple singularities.
+Looijenga [16] implicitly listed the local connected components of the complements of the discriminants of simple singularities, although the numbers of these components remained unknown.
 O.
@@
 For simple singularities with Milnor numbers at most six this fact also follows from the results of [19].
-Unfortunately, I only read the work [30] containing the enumeration of topological types of simple and parabolic function perturbations in two variables after the publication of [24] and [25].
+Unfortunately, I only read the work [31], which contains the enumeration of topological types of simple and parabolic singularity perturbations in two variables after the publication of [25] and [26].
 1.4. 1-cohomology of discriminant complements
@@
 In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [18], [23]).
-Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6].
+Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6], see also [2].
 Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [21].
@@
 The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$.
-The discriminant subvarieties of these spaces consist of functions that have at least one real critical point with zero critical value.
+The discriminant subvarieties of these spaces consist of polynomials that have at least one real critical point with zero critical value.
 These varieties divide the spaces $\Phi_{*}$ into connected components.
@@
 Thus, it also proves the above Proposition 3.
-These constructions provide upper bounds of the numbers of isotopy classes of generic Morse perturbations associated with virtual functions.
+These constructions and statements allow us to prove upper bounds of the numbers of isotopy classes of generic Morse perturbations associated with particular virtual functions.
 These bounds are given in Propositions 9, 16, and 21 for classes $X_{9}^{+}$, $X_{9}^{1}$ and $X_{9}^{2}$ separately.
@@
 Definition 3.
-For any generic polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and with an odd Morse index.
+For any generic polynomial $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$, the number $\mbox{Ind}(f)$ equals the number of its real critical points with a negative critical value and an even Morse index minus the number of critical points with a negative critical value and an odd Morse index.
 It is easy to see that this number is an invariant of the connected components of the discriminant complements, and also of associated virtual components.
@@
 Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions).
-Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$.
-(For this existence statement, the parabolicity of $X_{9}$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
+Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$.
+(For this existence statement, the parabolicity of $X_{9}$ singularities is essential.
+The proof of it is based on the results of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
 First, suppose that it is the same point $I(1)$.
-A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map: this is the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$.
-Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this automorphism.
+A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map.
+This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [28].
+Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or mapped to each other by this automorphism.
 In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$.
@@
 2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
-Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane.
+Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, whose points are arbitrary $k$-element subsets of the plane.
 For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$.
-The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
+The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\setminus\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
 Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals.
@@
 Theorem 13 (cf. [27], § 1.8.3).
-The class $W$ takes unit values on some 1-cycles in each of the connected components shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
+The class $W$ takes unit values on some 1-cycles in each of the connected components containing the polynomials shown in pictures ((d)), ((e)), and ((g)) of Fig. 1.
 Proof. These cycles are the families (5), (8), and (11). $\Box$
@@
 This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y).
-These involutions have no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
+This involution has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
 For any polynomial $f$ with a principal homogeneous part of the form (12), its zero-level set goes to the infinity along four branches: two vertical and two horizontal.
@@
 Each of these polynomials is invariant under the transformation $(x,y)\mapsto(y,x)$.
-The transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (16) to the polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
+The transformation $(x,y)\mapsto(-x,-y)$ maps each polynomial (16) to the analogous polynomial with the opposite value of the parameter $t$, thus it permutes two connected components represented by these polynomials.
 The same considerations hold for components obtained from these by the involution (13). This concludes the proof of Theorem 14. $\Box$
@@
 Remark 20.
-Picture ($c$) and its rotation by the angle $\frac{\pi}{2}$ are topologically equivalent, but the connected components of the discriminant complement realized by polynomials with these shapes are different.
+Picture ($c$) and the result of its rotation by the angle $\frac{\pi}{2}$ are topologically equivalent, but the connected components of the discriminant complement realized by polynomials with these shapes are different.
 Theorem 19 was conjectured in [25], and the difference of all the connected components mentioned there was proven.
@@
 (21) $\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.$
-Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [12].
+Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [12], p. 160.
 […]
@@
 Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement.
-These four components differ for the topological reasons.
+These four components differ for the topological reasons, see [25].
 For the remaining nine virtual components, the convenient realizations of the corresponding real connected components can be obtained from the polynomials (22)–(30) via the operation (17). $\Box$
-3. Components for $J_{10}$ singularities
+3. Components of discriminant complements for $J_{10}$ singularities
 3.1.
@@
 (32) $(x-Ay^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad A\in(-1,1),$
-with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, cf. [12].
+with ten parameters $\lambda_{1},\dots,\lambda_{9},$ and $A$, cf. [12], [30].
 Definition 7.
-A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (31) or (32) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal quasihomogeneous part of $f$ (of degree six with weights $\deg x=2,\deg y=1$) consists of three different complex parabolas or two parabolas and one line in ${\mathbb{C}}^{2}$.
+A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of the form (31) or (32) is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, all of its critical values are different and not equal to 0, and the complex zero-level set of the principal weighted homogeneous part of $f$ (of degree six with weights $\deg x=2,\deg y=1$) consists of three different complex parabolas or two parabolas and one line in ${\mathbb{C}}^{2}$.
 Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively.
@@
 (33) $f(x,y)\leftrightarrow f(x,-y)\ .$
-In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components of the discriminant complements in $\Omega_{*}$. $\Box$
+In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components of the discriminant complement in $\Omega_{*}$. $\Box$
 In the following pictures, it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
@@
 On the other hand, the involution (34) moves the component of picture $(f)$ to the same component as (33).
-An invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ realizing the picture $(e)$ from its mirror image f(x,y)f(x,-y) is as follows.
-Four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement.
-Let us take three arbitrary points inside the rightmost three ovals.
-There is a single parabola (or line) of the form $x=ay^{2}+by+c$ that passes through these three points.
-The restriction of the polynomial $f$ to this parabola is a polynomial of degree at most six in the coordinate $y$.
-This restriction has six roots, so it is of degree exactly six.
-This parabola cannot intersect the fourth oval, hence it passes either above or below it with respect to the $x$ coordinate.
-The position of this parabola relative to the fourth oval is independent of the choice of three points inside the ovals, and remains unchanged along the connected components of the discriminant complement.
-However, these relative positions are different for any polynomial $f(x,y)$ with the topological shape of the zero-level set shown in picture $(e)$ and for its mirror image f(x,y)f(x,-y).
+The following is an invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ from its mirror image f(x,y)f(x,-y).
+The four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement.
+Let us take an arbitrary point inside each of these four ovals.
+There is a single curve with the equation $x=ay^{3}+by^{2}+cy+d$ that passes through these four points.
+The coefficient $a$ of this equation cannot be zero.
+Indeed, otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial of degree at most six with eight roots.
+Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement.
+Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define the opposite signs.
 The same is true for polynomials that realize the picture (e~)($\tilde${e}).
@@
 6.
-Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal quasihomogeneous part will be of class $J_{10}^{3}$.
+Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal weighted homogeneous part will be of class $J_{10}^{3}$.
 The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values.
@@
 5.
-Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal quasihomogeneous part of the resulting polynomial will have the canonical form, see (32).
+Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32).
 Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$.
-The resulting polynomial realizes the picture $(e)$, and is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$.
+The resulting polynomial realizes the picture $(e)$.
+It is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$.
 $\Box$
@@
 4.1.
-Two classes of function singularities $P_{8}^{1}$ and $P_{8}^{2}$ in three variables have the canonical normal form
+Two classes $P_{8}^{1}$ and $P_{8}^{2}$ of function singularities in three variables have the canonical normal form
 (41) $x^{3}+y^{3}+z^{3}-3Axyz,\quad A\neq 1,$
 with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5].
-The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$.
+The zero-level sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$.
 Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin.
-The zero set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
+The zero-level set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
 The standard versal deformations of $P_{8}^{1}$ and $P_{8}^{2}$ singularities consist of the polynomials
@@
 Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$, respectively.
+It follows easily from Morse theory that the invariant Ind can take only the values 1,0,1,2,1,0,-1,-2, and 3-3 on polynomials of classes $\Xi_{1}$ and $\Xi_{2}$.
+The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
 4.2. The homological invariant of real and virtual components
@@
 4.3. The symmetry group and the relation between the real and virtual components
-The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values.
+The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data sets of virtual functions only include the parities of the Morse indices of critical points, and not their integer values.
 We also have the following analog of Proposition 3.
@@
 2.
-If two virtual functions of the class $\Xi_{1}$ ((respectively, $\Xi_{2})$ belong to the same virtual component, then they are associated with two generic polynomials of class $\Xi_{1}$ ((respectively, $\Xi_{2})$ that belong to the same connected component of the space of generic polynomials.
+If two virtual functions of the class $\Xi_{1}$ ((respectively, $\Xi_{2})$ belong to the same virtual component, then they are associated with two generic polynomials of type $\Xi_{1}$ ((respectively, $\Xi_{2})$ that belong to the same connected component of the space of generic polynomials.
 $\Box$
@@
 The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 35 (respectively, 36) repeats the deduction of Proposition 9 from Lemma 10.
+Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$.
 $\Box$
@@
 These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates.
-According to Proposition 34, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
+Therefore, according to Proposition 34, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
 The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2.
@@
 The polynomials
-(49) ${\bf T}\left(x^{3}+y^{3}+z^{3}+3(x+y+z)^{2}-(x^{2}+y^{2}+z^{2})/2+\frac{7}{36}\right)$
+(49) ${\bf T}\left(x^{3}+y^{3}+z^{3}+3(x+y+z)^{2}-\frac{1}{2}(x^{2}+y^{2}+z^{2})+\frac{7}{36}\right)$
 and
-(50) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(x+y+z)^{2}+(x^{2}+y^{2}+z^{2})/2-\frac{7}{36}\right)$
+(50) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(x+y+z)^{2}+\frac{1}{2}(x^{2}+y^{2}+z^{2})-\frac{7}{36}\right)$
 have $\mbox{Ind}=-3$.
@@
 where $\psi_{2}$ and $\psi_{1}$ are some homogeneous polynomials of degrees 2 and 1, respectively.
-The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [3], §12, this polynomial can be reduced to the normal form $\tilde{z}^{2}+\tilde{x}^{3}+\tilde{x}\tilde{y}^{2}$ of $D_{4}^{+}$ class in some local coordinates $\tilde{x},\tilde{y}$, and $\tilde{z}$.
+The lowest weighted homogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [3], §12, this polynomial can be reduced to the normal form $\tilde{z}^{2}+\tilde{x}^{3}+\tilde{x}\tilde{y}^{2}$ of $D_{4}^{+}$ class in some local coordinates $\tilde{x},\tilde{y}$, and $\tilde{z}$.
 2.
-Let $\alpha$ be the coefficient of the polynomial $\psi_{2}$ at the monomial $q_{l}^{2}$.
-Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
-Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest quasihomogeneous part
+Let $\alpha$ be the coefficient at the monomial $q_{l}^{2}$ of the polynomial $\psi_{2}$ in (54).
+Then the lowest weighted homogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
+Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest weighted homogeneous part
 $\tilde{l}^{2}-\varepsilon P_{l}^{2}+\left(\frac{\alpha^{2}}{4}+\frac{1}{4\varepsilon}\right)q_{l}^{4}\ .$
@@
 It is easy to calculate that for $l\equiv x-y$ it is equal to 1.
-The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points and the local maximum at the $A_{3}$ point.
-Therefore the critical value at its Morse critical point is negative.
+The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points; its critical point at the origin is a local maximum.
+Therefore, the critical value of $F(l)$ at the Morse critical point is negative.
 $\Box$
@@
 Definition 11.
-The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that component.
+The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that virtual component.
 The involution (45) of virtual components of classes $\Xi_{1}$ and $\Xi_{2}$ preserves this invariant.
@@
 Proof of this theorem takes the rest of this section.
-All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$ action.
+All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$-action.
 Thus, if $f$ is any of these polynomials and $C$ is a noncritical value of $f$, then Proposition 34 implies that the virtual component associated with the polynomial fCf-C is associated with only one real component of the discriminant complement.
@@
 According to Proposition 41 and Table 3, these are two components with $\mbox{Card}=94$.
-According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has $\mbox{HI}=(031)$.
+According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has the homology index $(031)$.
 Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants.
@@
 By Proposition 40 of [29], the intersection matrix of the canonical basis of vanishing cycles for any generic polynomial of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Coxeter-Dynkin graph shown in
-Figure 7.
+Figure 7. Coxeter–Dynkin graph for (58)
 Fig.
@@
 12536748
-Figure 8. Coxeter-Dynkin graph for (59
+Figure 8. Coxeter-Dynkin code for (59)
 3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is as follows:
@@
 Define the map $\Lambda:U\to{\mathbb{R}}^{8}$, mapping each point $\tilde{f}\in U$ to the sequence of its correspondingly ordered critical values.
-By the versality property, this map is a local diffeomorphism.
+By the versality property of the family $\Xi_{2}$, this map is a local diffeomorphism.
 Let $\varepsilon>0$ be a very small number.
@@
 $\Box$
-Due to the construction of the polynomial $I(1)$, the Card invariant of arbitrarily small generic perturbations of this polynomial equal to that of the initial polynomial $f_{0}$, that is, it is equal to 60.
+Due to the construction of the polynomial $I(1)$, the Card invariant of arbitrarily small generic perturbations of this polynomial is equal to that of the initial polynomial $f_{0}$, that is, it is equal to 60.
 This proves Proposition 45.
@@
 Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points.
-Of these 20 virtual functions, only 12 have these positive critical values at the critical points with Morse index 2.
+Of these 20 virtual functions, only 12 have all these positive critical values at the critical points with Morse index 2.
 Thus, the virtual function associated with $\bar{f}$ is among these 12.
@@
   • for each pair of vertices of the graph that are connected by an edge, the critical value of the critical point corresponding to the higher vertex in this figure is greater than the critical value of the critical point corresponding to the lower vertex.
-All vertices of this graph correspond to some critical points (and their critical values) of the polynomial $f!$.
+The virtual function associated with the polynomial $f!$ is determined by this graph with the above information on the Morse indices of vertices and the numbering of these vertices corresponding to the ascending order of critical points.
+The vertices with numbers 1–4 correspond to negative critical values, and vertices 5–8 correspond to positive critical values.
 According to the previous conditions, the lowest vertex in Fig.
-7 corresponds to the value marked 1 in Fig.
-9 (right), the vertices in the middle row correspond to the values marked 2, 3, 4, and 5, and the vertices in the upper row correspond to the values marked 6, 7, and 8.
+7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8.
 The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5 of Fig.
@@
 Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [23].
-Therefore, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig.
-7 and one vertex of valence 3 in the middle row.
-The Coxeter-Dynkin graph of Fig. 7, together with
-  • (a)
-    the numbering of its vertices according to the order of the corresponding critical values,
-  • (b)
-    the number of negative critical values, and
-  • (c)
-    the Morse indices of the critical points corresponding to its vertices,
-defines uniquely a virtual function, see Remark 1.
-All virtual functions that can be defined by the graph of Fig.
-7 and information (a), (b), and (c) satisfying the above restrictions on numbering and Morse indices, belong to the same virtual component.
-Indeed, they can all be obtained from each other by transpositions of orders of vertices that are not connected by edges of the graph and corresponding to the critical values of the same sign.
+In particular, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig.
+7 and one vertex of valence three in the middle row.
+All virtual functions with intersection matrix defined by the graph of Fig.
+7 and the Morse indices and numbering of vertices satisfying these restrictions belong to the same virtual component.
+Indeed, they can all be obtained from each other by transpositions of orders of vertices that are not connected by edges of the graph and correspond to the critical values of the same sign.
 Such a transposition is a flip of virtual functions which corresponds to the elementary function surgery when one critical value overtakes another critical value at a distant critical point.
@@
 $\Box$
+It remains to prove that the connected component of the discriminant complement containing the polynomial $f!$ is $S(3)$-invariant.
 Let $L\sim S^{1}$ again be the set of all linear functions ${\mathbb{R}}^{3}\to{\mathbb{R}}$ of unit norm that vanish at the vector $(1,1,1)$.
@@
 10 (left).
-This implies the first assertion of the lemma analogously to the proof of Lemma 39.
+The first assertion of the lemma follows from this analogously to the proof of Lemma 39.
 Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$.
-The restriction of this polynomial to the line connecting 0 and $a$ is a polynomial of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0.
+The restriction of the polynomial to the line through 0 and $a$ is a polynomial function of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0.
 $\Box$
 Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin.
-These numbers form a lower semi-continuous function on $L$.
+These numbers $c(l)$ form a lower semi-continuous function on $L$.
 By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$.
@@
 This union is connected and invariant under the action (44) of the group $S(3)$.
-By Lemma 47, its point corresponding to $l\equiv x-y$ belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$.
-Therefore, this entire component is $S(3)$-invariant.
+By Lemma 47, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$.
+The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant.
+Therefore, the entire component is $S(3)$-invariant.
 This concludes the proof of Theorem 42 for the virtual component with $\mbox{Card}=1216$.
@@
 Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin.
-Therefore, the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
+Therefore, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
 (62) $x^{2}y-y^{3}+\varkappa_{1}+\varkappa_{2}x+\varkappa_{3}y+\varkappa_{4}y^{2}$
 of the sample polynomial $x^{2}y-y^{3}$ of class $D_{4}^{-}$ in two variables with parameters $\varkappa_{i}$.
-Denote the fiber over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
+Denote the fiber of this tubular neighborhood over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
   • (a)
@@
     $\mbox{Card}(f)=844$.
-Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) and condition (b) replaced by $\mbox{Card}(f)=1318$.
+Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
 Lemma 49.
@@
 Therefore, their union over all $l\in L$ is connected, and we again have only one real connected component.
-To prove Lemma 49, we will describe detail the geometry of the discriminant in the parameter space of the versal deformation (62) in detail.
+To prove Lemma 49, we will describe the geometry of the discriminant in the parameter space of the versal deformation (62) in detail.
 The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$.
-In the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig.
+In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig.
 11 (left), see [3], § 21.3.
@@
 The larger part consists of polynomials with only two real critical points.
-Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one point that is either a minimum or a maximum, depending on the pyramid.
+Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one critical point that is either a minimum or a maximum, depending on the pyramid.
 We are only interested in the polynomials (62) with a minimum point.
@@
 (63) $x^{2}y-y^{3}+\varkappa_{2}x+\varkappa_{4}y+\varkappa_{4}y^{2},$
-i.e., it can belong to one of the five intervals into which the critical values divide the line.
+i.e., it can belong to one of the five intervals into which the critical values divide the line $\mathbb{R}$}.
 For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal.
@@
 10 (right).
-These polynomials, representing various sections of the pyramid, sweep out its central axis.
+These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis.
 An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
 Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$.
-This marking of critical points can be uniquely extended by continuity to all other polynomials within the pyramid.
-Each of the six cells into which the curvilinear triangle in Fig.
-11 (center) is split by the segments is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
-The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of nine connected domains.
+This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid.
+Each of the six cells, into which the curvilinear triangle in Fig.
+11 (center) is split by the segments, is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
+The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of eight connected domains.
 We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value.
@@
 Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph.
-These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
+These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials of class $\Xi_{2}$ having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
 Consider again the case $l\equiv x-y$.
@@
 Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$.
-For the same reasons as in the aforementioned proof, the corresponding critical points meet at a real critical point of class $A_{2}$.
+For the same reasons as in the proof of Lemma 47, the corresponding critical points meet at a real critical point of class $A_{2}$.
 Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points.
@@
 Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex.
-This choice explicitly determines the virtual component associated with the obtained polynomial (and?
-consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
+This choice explicitly determines the virtual component associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
 Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant.
@@
 $\Box$
-As $l$ moves along half of the circle $L$, the polynomial $F(l)$ and the corresponding slice $U(l)$ returns to themselves.
-The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this half-circle is a connected set that is $S(3)$-invariant.
+The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
+The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this circle is a connected set that is $S(3)$-invariant.
 All of these domains consist of polynomials with $\mbox{Card}=844$.
 This proves the assertion of Theorem 42 for the virtual components with $\mbox{Card}=844$.
+The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the same circle, with the fiber $W(l)$ consisting of two contractible domains.
 Lemma 50.
-The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ of all polynomials $F(l)$ permutes the two components of the sets $W(l)$ mentioned in Lemma 49.
+The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $D_{4}^{-}$ stratum at the corresponding polynomials $F(l)$.
 $l=0$
 $l=0$
-Figure 12. The plane $\{l=0\}$ and the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
+Figure 12. The plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
 Proof. The zero-level set of the polynomial $F(l)$ in the plane $\{l=0\}$ consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set
-(64) $x^{3}+y^{3}+z^{3}-6xyz=0,$
-which is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 (left).
-Consider the local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$ at the origin in ${\mathbb{R}}^{3}$ such that $l\approx\tilde{l}$ up to terms of second degree, and $F(l)$ has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class.
+(64) $x^{3}+y^{3}+z^{3}-6xyz=0$
+that is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$.
+The third line belongs to the component that is homeomorphic to the plane, see Fig.
+12 left (in which we schematically depicted the set (64) by the union of the quadratic cone and the horizontal plane).
+The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree.
 The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin.
-The curves on this surface that are distinguished by the conditions $\tilde{p}=0$, $\tilde{p}=\tilde{q}$, and $\tilde{p}=-\tilde{q}$ are tangent to the three lines considered in the previous paragraph.
-Let $\hat{F}(l)\in U(l)$ be a perturbation of $F(l)$ that splits the $D_{4}^{-}$ singularity into a minimum point with critical value $\varepsilon$ and three critical points of Morse index 1 with critical value 0.
-In canonical coordinates, the zero-level set of the restriction of $\hat{F}(l)$ to the surface $\{\tilde{l}=0\}$ locally consists of three curves obtained by small deformations of the three curves for the function $F(l)$, and the three critical points occur at their intersections, see Fig.
+The curves on this surface that are distinguished by the conditions
+(65) $\tilde{p}=0,\ \tilde{p}=\tilde{q},\ \mbox{and}\ \tilde{p}=-\tilde{q}$
+are tangent to the three lines considered in the previous paragraph.
+Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0.
+This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig.
+11 left.
+Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig.
+10 right.
+The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$.
+In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$.
+The three critical points with Morse index 1 occur at their intersections, see Fig.
 10 (right).
-Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the nearby domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at exactly this critical point is positive and the critical values at the other two such points are negative.
-Thus, there is a canonical one-to-one correspondence between these domains in $U(l)$ and the choices of some two out of three lines in the plane $\ker l$ considered in the previous paragraph.
-As $l$ moves along the half-circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute.
-Thus, the monodromy over this half-circle acts in the same way on the set of three related domains in $U(l)$.
+Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative.
+Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
+As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute.
+Thus, the monodromy over this circle acts in the same way on the set of three related domains in $U(l)$.
 We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 49.
@@
 Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig.
-12 ((right)).
+12 right.
 The polynomial
 (66) $x^{3}+y^{3}+z^{3}-6xyz+l^{2}$
-then has a singularity of class $D_{5}^{-}$ at the origin in ${\mathbb{R}}^{3}$. It has the form
+then has a singularity of class $D_{5}^{-}$ at the origin in ${\mathbb{R}}^{3}$. In particular, it has the form
 (67) $\zeta^{2}+\xi^{2}\eta-\eta^{4}$
@@
 Proof.
-All singularities of corank 2 that can occur in the deformation (42) of $P_{8}^{2}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [19], [12].
+All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12].
 The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$.
@@
 (68) $l\equiv 2x-5y+11z,$
-the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and Y=y3zY=y-3z reduce the polynomial (65) to a polynomial whose lower quasihomogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}Y^{2}z-\frac{9}{4}z^{4}$.
-A dilation of coordinates reduces it to the form (66).
+the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (66) to a polynomial whose lower weighted homogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$.
+A dilation of coordinates reduces it to the form (67), and technique of [3], §12, removes all higher weighted homogeneous terms of the Taylor expansion.
 The set of all planes in ${\mathbb{R}}^{3}$ that are tangent to the same component of the zero-level set of (64) is connected.
@@
 Lemma 52.
-For any linear function $l$ as in Lemma 51, the set of polynomials with $\mbox{Ind}=-3$ consists of a single connected component in a small neighborhood of the corresponding polynomial ((65)).
+For any linear function $l$ as in Lemma 51, the set of polynomials with $\mbox{Ind}=-3$ in a small neighborhood of the corresponding polynomial ((66)) consists of a single connected component.
 Proof.
@@
   • it has exactly six real critical points: three with Morse index 1 and three with Morse index 2,
-  • the critical values at all points with Morse index 1 are lower than all critical values at the points with Morse index 2,
+  • the critical values at all critical points with Morse index 1 are lower than all critical values at the points with Morse index 2,
   • the intersection indices of vanishing cycles corresponding to the five lowest real critical values are defined by the matrix
 […]
-Indeed, this polynomial can be obtained from the perturbation of (65) shown in the middle part of Fig.
+Indeed, a polynomial with these properties can be obtained from the perturbation of (66) shown in the middle part of Fig.
 13 by a minor additional perturbation that properly orders the critical values.
-The intersection indices of its critical values are derived using the Gusein-Zade—A’Campo method [10].
-Our program shows that only six different virtual functions of $P_{8}^{2}$ class satisfy these conditions.
-It also shows that all six of these virtual functions appear in the virtual component with $\mbox{Card}=156.$ $\Box$
+The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10].
+Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions.
+It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
 The set of all polynomials (66) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant.
-The same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}\}$ stratum at the these points.
+According to Lemma 52, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at these points.
+Indeed, this union is the space of a fiber bundle with contractible fibers over the space of such planes.
 This implies the statement of Theorem 42 for the virtual component with $\mbox{Card}=156$.
@@
 In this theory, generating families of functions play a crucial role.
-Wave functions are defined on the parameter spaces of these families via integrals along certain cycles in the zero-level sets of the corresponding functions.
+Wave functions are defined on the parameter spaces of these families via integrals of special differential forms along certain cycles in the zero-level sets of the corresponding functions.
 Wavefronts are the discriminant sets of these families.
@@
 Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
-If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [22]).
+If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [23], Chapter V).
 Versal deformations of stably equivalent function singularities can be naturally identified with each other.
-This identification establishes a one-to-one correspondence between the components of their discriminant complements.
-The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number of variables, and on the parities of the inertia indices of the quadratic part of the singularity.
+This identification establishes a one-to-one correspondence between the components of the discriminant complements of these deformations.
+The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number $N$ of variables, and on the parities of the positive inertia indices $i_{+}$ of the quadratic part of the singularity.
 For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas.
@@
 The above explicit enumeration of connected components of the discriminant complements proves this conjecture for all singularity classes except for $P_{8}^{2}$.
-For $P_{8}^{2}$, we find one new local lacuna, thus replacing the number 2 conjectured in [23] by 3.
+For $P_{8}^{2}$, it gives us one new local lacuna, thus replacing the number 2 conjectured in [24] by 3.
 In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig.

### v5

@@
 Abstract.
-We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most common family of singularity classes of smooth functions after simple singularities).
+We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities).
 Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [25] with the same title.
-As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs (one of which is a new one).
+As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs, in particular, find a new one.
 We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
-These results are applications of a general method for investigating non-singular perturbations of real function singularities.
+These results are applications of a general method for investigating and separating non-singular perturbations of real function singularities.
 An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
@@
 Primary: 14Q30. Secondary: 14B07, 14P25
+This work was supported by the Absorption Center in Science of the Ministry of Immigration and Absorption of the State of Israel
 1. Introduction
-$1.$
-Discriminant varieties of families of smooth functions depending on parameters are the sets of parameters for which the zero-level sets of the corresponding functions are singular.
+1.1. Motivation and terminology
+The discriminant variety of a family of smooth functions that depend on parameters is the set of parameters for which the zero-level sets of the corresponding functions are singular.
 These varieties appear in many problems of PDE theory and integral geometry, often under the name “wavefronts”, see e.g.
@@
 Many functions studied in these theories are defined by integral transforms.
-Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals along the cycles in the zero-level sets of the corresponding functions.
+Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals of certain differential forms along the cycles in the zero-level sets of the corresponding functions.
 These integral functions are regular outside the discriminant varieties of the families.
-The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant, from which we approach this point.
+The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant variety, from which we approach this point.
 Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
@@
 Two function singularities, $f(x_{1},\dots,x_{n})$ and $\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})$, are stably equivalent if they become equivalent after the summation with non-degenerate quadratic forms in additional variables, that is,
-$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-k})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m}).$
+$f(x_{1},\dots,x_{n})+Q_{2}(y_{1},\dots,y_{N-n})\sim\tilde{f}(\tilde{x}_{1},\dots,\tilde{x}_{m})+\tilde{Q}_{2}(\tilde{y}_{1},\dots,\tilde{y}_{N-m})$
+for some $N$ and quadratic polynomials $Q_{2},$ $\tilde{Q}_{2}$ of maximal ranks.
 The critical points of analytic functions ${\mathbb{R}}^{n}\to{\mathbb{R}}$ are equivalent if the function singularities obtained from their germs at these points by subtracting the critical values are equivalent.
-An $l$-parametric deformation of a function singularity $f$ is a family of analytic functions, $f_{\lambda}$, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables is regular analytic.
-The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a real critical point with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level set of the analytic continuation of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ is singular at this point).
+An $l$-parametric deformation of a function singularity $f$ is a family $\{f_{\lambda}\}$ of analytic functions, where $\lambda$ runs through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ is regular analytic in a neighborhood of the origin in ${\mathbb{R}}^{n+l}$.
+The discriminant variety of the deformation $F$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding functions $f_{\lambda}$ have real critical points with zero critical value near the origin in ${\mathbb{R}}^{n}$ (or, equivalently, the zero-level sets of the analytic continuations of $f_{\lambda}$ to ${\mathbb{C}}^{n}$ are singular at these points).
 For each isolated function singularity, studying all its deformations reduces to studying any its sufficiently large (so-called versal, see [3]) deformation.
-The complements of discriminant varieties of all versal deformations of stably equivalent singularities are homotopically equivalent, in particular the connected components of these complements are in a one-to-one correspondence.
-Therefore, for any stable equivalence class of singularities it is sufficient to enumerate these components for only one versal deformation of one function of this class.
+The complements of discriminant varieties of all versal deformations of stably equivalent singularities are homotopically equivalent, in particular there is a one-to-one correspondence between the connected components of these complements.
+Therefore, for any stable equivalence class of singularities it is sufficient to list these components for only one versal deformation of one function of this class.
 The list of equivalence classes of function singularities begins with simple singularities, which were identified by V.
@@
 By definition, two small non-discriminant perturbations $f_{\lambda}:{\mathbb{R}}^{2}\to{\mathbb{R}}$ of the same isolated function singularity have the same topological type if there is an isotopy between their zero-level sets in a neighborhood of the critical point that preserves the behavior of these sets at the boundary of the neighborhood.
-Clearly, the topological type is an invariant of the connected components of the discriminant complements.
+Clearly, the topological type is an invariant of the local connected components of the discriminant complements.
 Our “rigid isotopy” classification of non-discriminant perturbations, i.e., the classification whose classes are these components, is only slightly stricter than the topological classification, see Table 2.
@@
 This classification consists of five and ten classes for $P_{8}^{1}$ and $P_{8}^{2}$ singularities, respectively.
+1.2. Structure and interpretation of results
+Each class of parabolic singularities is unimodal, that is, it is swept out by a one-parametric family of orbits of the natural action of the group of local diffeomorphisms $({\mathbb{R}}^{n},0)\to({\mathbb{R}}^{n},0)$.
+In each case, we consider the canonical versal deformation of the corresponding class.
+This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, ($J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42).
+The polynomials with parabolic singularities of the corresponding class form a subset that is diffeomorphic to an open interval in this space.
+All orbits are represented by some points of these sets.
+Our main results are the lists of connected components, into which the discriminant varieties divide these spaces, see Theorems 6, 14, 19, 29, 32, 39, and 43.
+However, each of these results can be interpreted as a list of all local components of the discriminant complements in arbitrarily small neighborhoods of an arbitrary particular parabolic singularity of the corresponding class.
+Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of versal deformations.
+These groups are naturally isomorphic to the multiplicative group ${\mathbb{R}}^{1}_{+}$ of positive numbers.
+For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomial $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular preserves the principal homogeneous part of the polynomials.
+For $X_{9}$ singularities it acts as $\{t\}:f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $\{t\}:f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$.
+All of these actions preserve the discriminants and their complements.
+Parabolic singularities are the only invariant points of these actions.
+Each point of the discriminant complement is reduced by this action into an arbitrarily small neighborhood of some parabolic singularity.
+Additionally, according to [17], the discriminant variety is equisingular along the one-dimensional sets of parabolic singularities in parameter spaces of versal deformations.
+Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities in the one-parametric sets.
 1.3. On the method
@@
 Thus, the number 11 is replaced in Table 2 by 13.
-These results are summarized in Theorems 6, 14, 19, 29, 32, 38, and 42 below.
 1.4. On the discriminant complements for simple singularities
@@
 For simple singularities with Milnor numbers at most six this fact also follows from the results of [19].
-Unfortunately, I only read the work [31], which contains the enumeration of topological types of simple and parabolic singularity perturbations in two variables after the publication of [25] and [26].
+Unfortunately, I only read the work [31], which contains an enumeration of the topological types of perturbations of simple and parabolic singularities in two variables, after the publication of [25] and [26].
 1.5. 1-cohomology of discriminant complements
 E.
-Looijenga [16] proved that all local components of discriminant complements of real simple singularities are homotopically trivial.
-Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 40.
+Looijenga [16] proved that all local connected components of discriminant complements of real simple singularities are homotopically trivial.
+Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 41.
 1.6. An application to hyperbolic PDEs.
-In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [18], [23]).
-Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the lacunas were enumerated in [7] and [6], see also [2].
+In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [23]).
+Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the local lacunas were enumerated in [7] and [6], see also [2].
 Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [21].
-In [24], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that the list is complete.
+In [24], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete.
 We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
+A notation
+The symbol $\Box$ indicates either the conclusion of a proof or the absence of a proof, as in the case of immediate corollaries or statements supplied with references to the works where they are proven.
 2. Results for the $X_{9}^{*}$ classes
 2.1.
-Real parabolic function singularities of type $X_{9}$ form four stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$.
+Real parabolic function singularities of type $X_{9}$ form four one-parametric families of stable equivalence classes $X_{9}^{+}$, $X_{9}^{-}$, $X_{9}^{1}$, and $X_{9}^{2}$, see Table 1.
 The standard versal deformations of these singularities in all four cases have the same form: they are families of polynomials
@@
 The parameters $\{\lambda_{1},\dots,\lambda_{8},A\}$ identify the spaces $\Phi_{+}$ and $\Phi_{1}$ with ${\mathbb{R}}^{8}\times(-1,1)$, and the space $\Phi_{2}$ with ${\mathbb{R}}^{8}\times(-\infty,-1)$.
-The discriminant subvarieties of these spaces consist of polynomials that have at least one real critical point with zero critical value.
-These varieties divide the spaces $\Phi_{*}$ into connected components.
+The discriminant subvarieties divide the spaces $\Phi_{*}$ into connected components.
 Enumerating these components is our goal in this section.
@@
 Proposition 2.
-If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$, S(f~)S($\tilde${f}) are the same.
+If two generic polynomials, $f$ and $\tilde{f}$, belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$ and S(f~)S($\tilde${f}) are the same.
 $\Box$
@@
 However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$.
-Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves.
+Figure 1 essentially coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves.
 In our case, however, the isotopies are allowed only within the space $\Phi_{+}$ which is much smaller than the space of all degree four polynomials.
@@
 There are exactly seven virtual components of class $\Phi_{+}$.
-The shapes of the zero-level sets of some polynomials associated with virtual functions representing these virtual components are shown in Fig.
+The shapes of the zero-level sets of some polynomials of this class that are associated with virtual functions representing these virtual components are shown in Fig.
 1.
@@
 This group acts on the space $\Phi_{+}$ (each transformation $\Gamma:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ moves the polynomial $f$ to $f\circ\Gamma$).
-This action preserves the set of generic polynomials.
+This action preserves the set of generic polynomials and the set-valued invariant.
 Proposition 9.
@@
 In the next paragraph, we present the outline of this proof; we refer to [28] for all common details.
-Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function.
-Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions).
+Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the polynomials $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function.
+Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of polynomials with fewer than nine critical values).
 Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$.
 (For this existence statement, the parabolicity of $X_{9}$ singularities is essential.
-The proof of it is based on the results of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
+The proof of it is based on Proposition 2 of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2).
 First, suppose that it is the same point $I(1)$.
@@
 Let us prove that in fact each of these seven virtual components is associated with a single real connected component (that is, these components are $G_{0}$-invariant).
-Pictures $(a)$, $(c)$, $(b)$, $(f)$ of Fig. 1 are realized, respectively, by polynomials
+Pictures $(a)$, $(c)$, $(b)$, and $(f)$ of Fig. 1 are realized, respectively, by polynomials
 $x^{4}+y^{4}+t,$
@@
 $x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\qquad\alpha\in(0,1/4),$
+and
 $x^{4}+y^{4}-t(x^{2}+y^{2})+\alpha t^{2},\quad\alpha\in(1/4,1/2),$
@@
 The topological type of the curve shown in picture $(e)$ can be realized by any of the four polynomials
-(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa),$
+(6) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(x^{3}-3xy^{2})+\varkappa)$
+and
 (7) ${\bf T}(x^{4}+2Ax^{2}y^{2}+y^{4}\pm(y^{3}-3x^{2}y)+\varkappa)$
@@
 The transformation $(x,y)\mapsto(-x,-y)$ maps each such polynomial to the analogous polynomial with the opposite value of parameter $t$.
-In particular, it permutes the connected components described by pictures $(d)$ and $(e)$.
+In particular, it permutes the connected components of the discriminant complement that are described by pictures $(d)$ and $(e)$.
 Pictures $(f)$ and $(g)$ of Fig. 2 are realized by polynomials
@@
 2.6. Components for the $X_{9}^{2}$ singularities
-The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the complement of the discriminant in some way.
+The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement in some way.
 All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
 Lemma 23.
-Each polynomial of degree four in ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
+Each polynomial of degree four on ${\mathbb{R}}^{2}$ having a critical point of class $D_{6}$ with zero critical value and a critical point of class $A_{3}$ can be reduced by an affine transformation of ${\mathbb{R}}^{2}$ to the form
 (18) $\pm\left(t\left(\tilde{x}^{2}\tilde{y}+\frac{1}{3}\tilde{x}^{3}\right)+\frac{3}{4}\tilde{x}^{4}+3\tilde{x}^{3}\tilde{y}+\frac{3}{2}\tilde{x}^{2}\tilde{y}^{2}-\tilde{x}\tilde{y}^{3}\right)\ ,\quad t>0\ .$
 In particular, the j-invariant of its principal homogeneous part is equal to 5.
-Proof of Lemma 23. This lemma follows from Proposition 22 of [28] by the substitution $\tilde{x}=|a|^{\frac{-3}{4}}X$, $\tilde{y}=sign(a)|a|^{\frac{1}{4}}Y$, $t=|a|^{\frac{-5}{4}}$. $\Box$
+Proof of Lemma 23.
+This lemma follows from Proposition 22 of [28] by the substitution $\tilde{x}=|a|^{\frac{3}{4}}x$, $\tilde{y}=\mbox{sign}(a)|a|^{\frac{-1}{4}}y$, $t=|a|^{\frac{-5}{4}}$, and the sign $\pm$ in (18) equal to $\mbox{sign}(a)$.
+$\Box$
 Corollary 24.
@@
 If a polynomial $f(x,y)$ of degree four with principal homogeneous part
-(20) $x^{4}+2Ax^{2}y^{2}+y^{4}$
+(20) $x^{4}+2Ax^{2}y^{2}+y^{4},\quad A<-1,$
 has critical points of classes $D_{6}$ and $A_{3}$, then $A$ is a root of the equation
@@
 Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively.
-The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consist of ten elements.
+The virtual functions associated with generic polynomials of these types, the formal graphs of types $J_{10}^{1}$ and $J_{10}^{3}$, and their virtual components are defined in the same way as for the case $X_{9}$, except that the sets of vanishing cycles consist of ten elements.
 Proposition 26 (see [30]).
@@
 In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components of the discriminant complement in $\Omega_{*}$. $\Box$
-In the following pictures, it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
+In the following pictures (4)–(6), it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
 3.2. Components for $J_{10}^{1}$ singularities
@@
 $(x-2t)((x+t)^{2}+y^{4}-3ty^{2}+2t^{2}),$
+and
 $(x-y^{2}+2t)(x^{2}-2tx+y^{4}-5t^{2})-t^{3}/10$
@@
 The coefficient $a$ of this equation cannot be zero.
-Indeed, otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial of degree at most six with eight roots.
+Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with eight roots.
 Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement.
@@
 5.
-In all these cases except those shown in pictures $(h)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components.
+In all these cases except those shown in pictures $(g)$, $(l)$, and $(m)$, the polynomials $f(x,y)$ and f(x,y)-f(-x,y), which are related by the involution ((34)), represent different virtual components.
 $\Box$
@@
 The zero-level set of the resulting polynomial will have the desired shape shown in picture $(e)$ of Fig.
-5.
+5, see Fig.
+6 right.
 Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32).
@@
 The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
-4.2. The homological invariant of real and virtual components
+4.2. The homological invariant of real components
 Let $f$ be a polynomial in the class $\Xi_{1}$ or $\Xi_{2}$.
@@
 Consider three subsets in ${\mathbb{R}}^{3}$: $W_{0}=f^{-1}((-\infty,0])$, $W_{-}=f^{-1}((-\infty,T_{-}]),$ and $W_{+}=f^{-1}((-\infty,T_{+}])$.
-If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$) then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle).
+If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$), then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle).
 In particular, the reduced homology groups $\tilde{H}_{i}(W_{-})$ for $f\in\Xi_{2}$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
 Definition 9.
-The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f\in\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
+The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f$ of class $\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
 It is convenient to consider this homology group in the context of the exact homological sequence of the triple
@@
 We also have the following analog of Proposition 3.
-Proposition 33 (see [29], Corollary 8).
+Proposition 33 (see [29], Theorem 10).
+For any generic Morse polynomial $f$ of type $\Xi_{1}$ or $\Xi_{2}$, a virtual Morse function $\varphi$ associated with it, and a virtual Morse function $\tilde{\varphi}$ connected with $\varphi$ by an edge of the formal graph, there exists a generic Morse polynomial $\tilde{f}$ and a path in the space $\Xi_{1}$ or $\Xi_{2}$ connecting $f$ and $\tilde{f}$ and containing only one non-generic point at which it experiences a standard surgery of the same type as the edge $\varphi,\tilde{\varphi}$].
+Corollary 34.
 1. Any virtual function of the class $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
@@
 also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It defines an involution on both sets of real and virtual components. This involution preserves the invariant Ind.
-If $f$ and $\tilde{f}$ are two functions related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
+If $f$ and $\tilde{f}$ are two polynomials related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
 Proposition 35.
@@
 Proof.
-According to Lemma 35 of [29], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$.
-Three of these polynomials (mentioned in formula (26) of [29]) have signature (0,2) at the $A_{2}$ point.
-The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$.
+According to Lemma 38 of [29], there are only six polynomials of the form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$.
+Three of these polynomials (mentioned in formula (27) of [29] with a certain value of the parameter $\beta$) have signature (0,2) at the $A_{2}$ point.
+The remaining three polynomials have the form (46) up to the permutations of coordinates, see formula (28) in [29].
+In particular, they belong to one orbit of the group $S(3)$.
 $\Box$
 Lemma 37.
-The set of all polynomials of class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
+The set of all polynomials of of the form (42) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
 (47) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(1+\sqrt{3})xyz+(x+y+(1+\sqrt{3})z)^{2}\right),\$
@@
 Proof of Proposition 35.
-The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 35 (respectively, 36) repeats the deduction of Proposition 9 from Lemma 10.
+The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 36 (respectively, 37) repeats the deduction of Proposition 9 from Lemma 10.
 Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$.
@@
 Theorem 39.
-Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$.
+Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials in class $\Xi_{1}$.
 In particular, there are exactly seven such connected components.
@@
 For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$.
-According to the previous paragraph, the associated virtual components are also different.
+According to the last sentence of the previous paragraph, the associated virtual components are also different.
 The polynomials
@@
 Let $\alpha$ be the coefficient at the monomial $q_{l}^{2}$ of the polynomial $\psi_{2}$ in (54).
-Then the lowest weighted homogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
+Then the lowest weighted homogeneous part of the polynomial (52) with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$.
 Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest weighted homogeneous part
@@
 It is easy to calculate that for $l\equiv x-y$ it is equal to 1.
-The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points; its critical point at the origin is a local maximum.
-Therefore, the critical value of $F(l)$ at the Morse critical point is negative.
+The restriction of the polynomial (52) to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points; its critical point at the origin is a local maximum.
+Therefore, its critical value at the Morse critical point is negative.
 $\Box$
-Denote by $C(l)$ the critical value of the polynomial $F(l)$ at its Morse critical point. Then the polynomial
-$\tilde{F}(l)\equiv F(l)-\frac{C(l)}{2}$
-is non-discriminant and has $\mbox{Ind}=-1$.
-The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it.
-The union of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$.
+Denote by $C(l)$ the critical value of the polynomial (52) at its Morse critical point.
+Let $F(l)$ be the polynomial obtained from (52) by adding the constant $C(l)/2$.
+Then the polynomial $F(l)$ is non-discriminant and has $\mbox{Ind}=-1$.
+The latter polynomial is canonically determined by the linear form $l$ and depends continuously on it.
+The union of polynomials $F(l)$ over all $l\in L$ is connected and invariant under the action of the group $S(3)$.
 Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement.
@@
 Indeed, the topology of these sets remains constant along the component.
-For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
+For any polynomial $F(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
 Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$.
-These homology groups form a local system over the space of all such polynomials, i.e., over the connected component under consideration.
+These homology groups form a local system over the space of all such polynomials, i.e., over the connected component of the discriminant complement under consideration.
 In particular, the fundamental group of this component acts on these groups by monodromy.
-For each $l\in L$, one basic homology class of the zero-level set of the polynomial $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$.
-When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement.
+For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F_{l}$ is defined by the intersection of this set and the plane $\ker l$.
+When $l\in L$ runs a half of the circle $L$, the planes $\ker l$ run a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomials $F(l)$ run a closed cycle in the discriminant complement.
 The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation.
@@
 Definition 11.
-The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that virtual component.
+The invariant Card of a virtual component of the formal graph is the number of vertices in that virtual component.
 The involution (45) of virtual components of classes $\Xi_{1}$ and $\Xi_{2}$ preserves this invariant.
@@
 We will show that all virtual components with Card equal to 258, 336, 1648, 262, or 94 can be realized in this way.
-Hence, the assertion of Theorem 42 is true for these components.
+Hence, the assertion of Theorem 43 is true for these seven components.
 Components with $\mbox{\rm Card}=1648$.
@@
 The homology indices of these components are equal to $(000)$ and $(011)$.
-For example let $f$ be the $S(3)$-invariant polynomial
+For example, let $f$ be the $S(3)$-invariant polynomial
 (55) $x^{3}+y^{3}+z^{3}-4xyz.$
@@
 This polynomial is invariant under the involution (45).
-Therefore, according to the last statement of Proposition 41, the associated virtual component cannot have $\mbox{Card}=94$.
-Thus, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
+According to the last statement of Proposition 42, the associated virtual component cannot have $\mbox{Card}=94$.
+Therefore, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
 Components with $\mbox{\rm Card}=94$ and $336$. The polynomial
 (57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz+3(x^{2}+y^{2}+z^{2})$
-has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one point with Morse index $1$ and critical value $12$.
+has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one critical point with Morse index $1$ and critical value $12$.
 Therefore, the invariant $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{49}\right),$ to 2-2 for $C\in\left(\frac{88}{49},4\right),$ to 1 for $C\in\left(4,12\right),$ and to 0 for $C>12$.
@@
 Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$.
-The involution (45) moves this polynomial to that with $\mbox{HI}=(012)$.
+The involution (45) moves this polynomial to one with $\mbox{HI}=(012)$.
 In particular, these two polynomials represent two different components.
-They are $S(3)$-invariant, therefore by Proposition 34, the associated virtual components are also different.
+They are $S(3)$-invariant.
+Therefore, by Proposition 35, the associated virtual components are also different.
 According to Proposition 42 and Table 3, these are two components with $\mbox{Card}=94$.
@@
 Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{147}\right)$, to 2-2 for $C\in\left(\frac{88}{147},\frac{4}{3}\right)$, to 3-3 for $C\in\left(\frac{4}{3},4\right)$, and to 0 for $C>4$.
-Lemma 43.
+Proposition 44.
 […]
@@
 This graph does not contain edges of multiplicities other than 1 and 1-1.
-Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles is equal to 2-2.
+Upon a special request, our program verified that, for each virtual function of $P_{8}^{2}$ type from the virtual components with $\mbox{Card}=60$ or $156$, the intersection index of some two different vanishing cycles of the canonical basis is equal to 2-2.
 Therefore, the virtual component associated with our polynomial is the only remaining virtual component with $\mbox{Ind}=-3$.
@@
 Remark 45.
-It is easy to check that the polynomials fCf-C with $\mbox{Ind}=-2$ and $\mbox{Ind}=1$ give new realizations of some components with $\mbox{Card}=336$ and $94$, respectively.
+It is easy to check that the polynomials fCf-C with $f$ given by (58) and $C\in\left(\frac{88}{147},\frac{4}{3}\right)$ or $C\in\left(0,\frac{88}{147}\right)$ give new realizations of components with $\mbox{Card}=336$ and $94$, respectively.
 Components with $\mbox{Card}=60$.
@@
 12536748
-Figure 8. Coxeter-Dynkin code for (59)
+Figure 8. Coxeter-Dynkin graph for (59)
 3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is as follows:
@@
 4) the Morse indices of the points number 1, 2, 3, and 5 with respect to this order are odd, while the Morse indices of points 4, 6, 7, and 8 are even.
-According to Proposition 33, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data.
+According to Corollary 34, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data.
 According to Proposition 41 of [29], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2.
-Let $i:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $i(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $i(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
-Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $i(t)$ is the sequence of critical values of the polynomial $I(t)$.
+Therefore the polynomial $f_{0}$ has no local minima or maxima.
+Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
+Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$.
 As in the proof of Theorem 10 of [29], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined.
@@
 $\Box$
-The involution (45) moves the polynomial $I(1)$ to a polynomial in the other component of the discriminant complement with $\mbox{Card}=60$.
+By statement 3 of Proposition 42, the involution (45) moves the polynomial $I(1)$ to a polynomial in the different component of the discriminant complement with $\mbox{Card}=60$.
 Thus, this component is also represented by an $S(3)$-invariant polynomial.
@@
 Component with $\mbox{Card}=1216$. Denote by $\hat{f}$ the polynomial
-(60) $x^{3}+y^{3}+z^{3}-6xyz+3(x-y)^{2}$
+(60) $x^{3}+y^{3}+z^{3}-6xyz+3(x-y)^{2}.$
 This polynomial has a critical point of class $D_{4}^{-}$ with critical value 0 at the origin, two Morse critical points with Morse index 2 and a common critical value $\omega_{1}\approx 3.14$, and two non-real critical points with a common critical value $\omega_{2}\approx-11.63.$
@@
 Due to the versality of the space $\Xi_{2}$, for any sufficiently small $\varepsilon>0$ we can slightly perturb the polynomial $\hat{f}$ within the space $\Xi_{2}$ in such a way that the resulting polynomial $f_{\varepsilon}$ satisfies the following conditions:
-  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $2\varepsilon^{2}$,
+  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$,
   • the critical values at its two critical points with Morse index 2 are equal to $\omega_{1}$ and $\omega_{1}+\varepsilon^{2}$,
@@
 012341-1785116
-Figure 9. System of paths for $\bar{f}$ (left) and $f!$ (right)
+Figure 9. System of paths for polynomials $\bar{f}$ (left) and $f!$ (right)
 The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program.
@@
 In particular, this is true for the real polynomial $\bar{f}$.
-Using the Lyashko–Looijenga map in the form of [12], (see the proof of Theorem 7 in [29]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig.
+Using the Lyashko–Looijenga map in the form of [12] (see the proof of Theorem 7 in [29]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig.
 9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point.
 Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$.
-With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta$.
+With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta>0$.
 Denote this polynomial by $f!$, see Fig.
@@
   • for each pair of vertices of the graph that are connected by an edge, the critical value of the critical point corresponding to the higher vertex in this figure is greater than the critical value of the critical point corresponding to the lower vertex.
-The virtual function associated with the polynomial $f!$ is determined by this graph with the above information on the Morse indices of vertices and the numbering of these vertices corresponding to the ascending order of critical points.
-The vertices with numbers 1–4 correspond to negative critical values, and vertices 5–8 correspond to positive critical values.
-According to the previous conditions, the lowest vertex in Fig.
+The virtual function associated with the polynomial $f!$ is determined by this graph supplied with the information on the Morse indices of critical points corresponding to its vertices, and with the numbering of these vertices corresponding to the ascending order of critical values, i.e., the correspondence of these vertices with the black points in Fig.
+9 right.
+The graph and the Morse indices are known to us, while concerning the numbering we know the following.
+The lowest vertex in Fig.
 7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8.
-The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5 of Fig.
-9 (right).
+The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5.
 Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [23].
@@
 $\Box$
-It remains to prove that the connected component of the discriminant complement containing the polynomial $f!$ is $S(3)$-invariant.
+It remains to prove that the connected component of the discriminant complement that contains the polynomial $f!$ is $S(3)$-invariant.
 Let $L\sim S^{1}$ again be the set of all linear functions ${\mathbb{R}}^{3}\to{\mathbb{R}}$ of unit norm that vanish at the vector $(1,1,1)$.
@@
 The first assertion of the lemma follows from this analogously to the proof of Lemma 40.
-Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$.
+Let $a\in{\mathbb{R}}^{3}$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$.
 The restriction of the polynomial to the line through 0 and $a$ is a polynomial function of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0.
 $\Box$
+The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
 Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin.
-These numbers $c(l)$ form a lower semi-continuous function on $L$.
+These numbers $c(l)$ form a lower semi-continuous function on $L/{\mathbb{Z}}_{2}$.
 By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$.
@@
 Components with $\mbox{Card}=1318$ and $844$.
-Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other critical points with critical value 0.
+Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0.
 By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$.
@@
 of the sample polynomial $x^{2}y-y^{3}$ of class $D_{4}^{-}$ in two variables with parameters $\varkappa_{i}$.
-Denote the fiber of this tubular neighborhood over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
+Denote the fiber of this tubular neighborhood over the point $\{F(l)\}$, $l\in L$, by $U(l)$.
+For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
   • (a)
@@
 (63) $x^{2}y-y^{3}+\varkappa_{2}x+\varkappa_{4}y+\varkappa_{4}y^{2},$
-i.e., it can belong to one of the five intervals into which the critical values divide the line $\mathbb{R}$}.
+i.e., it can belong to one of the five intervals into which these critical values divide the line $\mathbb{R}$}.
 For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal.
@@
 Consider again the case $l\equiv x-y$.
-Let $f$ be a point in one of these three domains in the slice $U(l)$ at the point $F(l)$.
-We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\zeta$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points.
-Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$.
+Let $f$ be a point in one of these three domains in the corresponding slice $U(l)$.
+We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points.
+Then, acting as in the proof of Lemma 48, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\xi$ and $\tilde{\omega}_{1}$.
 For the same reasons as in the proof of Lemma 48, the corresponding critical points meet at a real critical point of class $A_{2}$.
@@
 Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex.
-This choice explicitly determines the virtual component associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
+This choice explicitly determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
 Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant.
@@
 $\Box$
-The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
-The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this circle is a connected set that is $S(3)$-invariant.
+The union of the contractible domains $V(l)\subset U(l)$ over all $l\in L$ is a connected set that is $S(3)$-invariant.
 All of these domains consist of polynomials with $\mbox{Card}=844$.
 This proves the assertion of Theorem 43 for the virtual components with $\mbox{Card}=844$.
-The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the same circle, with the fiber $W(l)$ consisting of two contractible domains.
+The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the circle $L/{\mathbb{Z}}_{2}$, with the fiber $W(l)$ consisting of two contractible domains.
 Lemma 51.
-The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $D_{4}^{-}$ stratum at the corresponding polynomials $F(l)$.
+The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
 $l=0$
@@
 The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin.
-The curves on this surface that are distinguished by the conditions
+The set $\{F(l)=0\}$ in this surface consists of three curves that are distinguished by the conditions
 (65) $\tilde{p}=0,\ \tilde{p}=\tilde{q},\ \mbox{and}\ \tilde{p}=-\tilde{q}$
-are tangent to the three lines considered in the previous paragraph.
+and are tangent to the three lines considered in the previous paragraph.
 Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0.
@@
 All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12].
-The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$.
+The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form.
+Therefore, its singularity at the origin is more complicated than $D_{4}$.
 According to Corollary 10.4.9 of [22], this singularity is not of class $E_{6}$.
@@
 Proof.
-By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}$.
+By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$.
 Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (66).
@@
 -$\Rightarrow$
-Figure 13. Perturbation of $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
+Figure 13. Perturbation of a $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
 Lemma 54.
 All the polynomials in the set considered in Lemma 53 have $\mbox{Card}=156.$
-Proof. We can assume that $l$ is given by (68). Then the set in consideration contains a generic polynomial with the following properties:
+Proposition 55.
+Let the linear function $l$ be given by ((68)). Then the set considered Lemmas 53, 54 contains a generic polynomial with the following properties:
   • it has exactly six real critical points: three with Morse index 1 and three with Morse index 2,
-  • the critical values at all critical points with Morse index 1 are lower than all critical values at the points with Morse index 2,
+  • the critical values at all critical points with Morse index 1 are lower than all critical values at the critical points with Morse index 2,
   • the intersection indices of vanishing cycles corresponding to the five lowest real critical values are defined by the matrix
 […]
-Indeed, a polynomial with these properties can be obtained from the perturbation of (66) shown in the middle part of Fig.
+Proof.
+A polynomial with these properties can be obtained from the perturbation of the polynomial (66) that is shown in the middle part of Fig.
 13 by a minor additional perturbation that properly orders the critical values.
 The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10].
+$\Box$
+Proof of Lemma 54.
+All these sets corresponding to different $l$ belong to the same connected component of the discriminant complement.
+Therefore, we can assume that $l$ is given by (68) and hence contains a polynomial satisfying the conditions listed in Proposition 55.
 Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions.
@@
 Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well.
-However, over the discriminant sets of parameters, the level sets degenerate and the wave function may become irregular.
+However, over the discriminant sets of parameters, the zero-level sets degenerate and the wave function may become irregular.
 The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point.
@@
 The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure.
-The lacunas for the $X_{9}^{-}$ singularities are obtained from the $X_{9}^{+}$ singularities by multiplying of all functions by 1-1.
+The lacunas for the $X_{9}^{-}$ singularities are obtained from these of $X_{9}^{+}$ singularities by multiplying of all functions by 1-1.
 The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig.

### v6

@@
 Abstract.
-We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities).
+We list all connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities).
 Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [26] with the same title.
-As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs, in particular, find a new one.
-We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
+As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs.
+We also show that the complements of the discriminant varieties of the versal deformations of $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
 These results are applications of a general method for investigating and separating non-singular perturbations of real function singularities.
-An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
+An important part of this method is a computer program that formalizes local Picard–Lefschetz theory and surgeries of Morse functions.
 Key words and phrases:
-discriminant, singularity, surgery, versal deformation, isotopy classification
+discriminant, singularity, surgery, versal deformation, isotopy classification, Lyashko–Looijenga covering, Picard–Lefschetz theory
 1991 Mathematics Subject Classification:
@@
 Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals of certain differential forms along the cycles in the zero-level sets of the corresponding functions.
-These integral functions are regular outside the discriminant varieties of the families.
+These integral functions are regular outside the discriminant varieties of the generating families.
 The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant variety, from which we approach this point.
-Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
+Thus, enumerating these components for different types of discriminant points is one of the first steps in studying the analytical properties of the integral transforms.
 Definition 1 (see [3]).
@@
 In each case, we consider the canonical versal deformation of the corresponding class.
-This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, ($J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42).
+This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, (respectively, $J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42).
 The polynomials with parabolic singularities of the corresponding class form a subset that is diffeomorphic to an open interval in this space.
-All orbits are represented by some points of these sets.
-Our main results are the lists of connected components, into which the discriminant varieties divide these spaces, see Theorems 6, 14, 19, 29, 32, 39, and 43.
+All orbits are represented by some points of these intervals.
+Our main results are the lists of connected components, into which the discriminant varieties divide the parameter spaces of these deformations, see Theorems 6, 14, 19, 30, 33, 40, and 44.
 However, each of these results can be interpreted as a list of all local components of the discriminant complements in arbitrarily small neighborhoods of an arbitrary particular parabolic singularity of the corresponding class.
-Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of versal deformations.
+Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of canonical versal deformations.
 These groups are naturally isomorphic to the multiplicative group ${\mathbb{R}}^{1}_{+}$ of positive numbers.
-For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomial $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular preserves the principal homogeneous part of the polynomials.
-For $X_{9}$ singularities it acts as $\{t\}:f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $\{t\}:f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$.
+For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomials $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular it preserves the principal homogeneous part of the polynomials.
+For $X_{9}$ singularities it acts as $f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$.
 All of these actions preserve the discriminants and their complements.
-Parabolic singularities are the only invariant points of these actions.
+Parabolic singularities are defined by the homogeneous or weighted homogeneous polynomials of top degree and are the only invariant points of these actions.
 Each point of the discriminant complement is reduced by this action into an arbitrarily small neighborhood of some parabolic singularity.
 Additionally, according to [17], the discriminant variety is equisingular along the one-dimensional sets of parabolic singularities in parameter spaces of versal deformations.
-Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities in the one-parametric sets.
+Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities of the same class.
 1.3. On the method
-In [25], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for parabolic singularities are found, and all of these values are realized by concrete polynomials.
-The remaining problem is to list all components with each value.
+In [26], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for all parabolic singularities are found, and all of these values for singularities of classes $X_{9}$ and $J_{10}$ are realized by concrete polynomials.
+The remaining problems are to list all components with each value of the invariant, and also to realize them for the $P_{8}$ singularities.
 In [26], a conjectural answer to this problem was proposed.
@@
 Looijenga [16] proved that all local connected components of discriminant complements of real simple singularities are homotopically trivial.
-Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 41.
+Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 42.
 1.6. An application to hyperbolic PDEs.
-In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [23]).
+In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [19], [24]).
 Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the local lacunas were enumerated in [7] and [6], see also [2].
@@
 A virtual function associated with a generic real polynomial $f$ is a collection of topological characteristics of the complex zero-level set of this polynomial.
-Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
+Specifically, these data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
   1. (1)
-    the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
+    the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold $V_{f}\subset{\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
   2. (2)
-    the string of intersection indices of these nine cycles with the set of real points,
+    the string of intersection indices of these nine cycles with the set of real points in $V_{f}$,
   3. (3)
@@
 In the statement of Proposition 2 of [29], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with non-degenerate principal homogeneous parts.
-However, the entire construction proving this proposition in [28] can be performed within the spaces of the versal deformations (1).
+However, the entire construction proving this proposition in [29] is performed within the spaces of the versal deformations (1).
 Thus, it also proves the above Proposition 3.
@@
 Theorem 6.
-There are exactly seven connected components in the space of non-discriminant functions of class $\Phi_{+}$.
+There are exactly seven connected components in the space of non-discriminant polynomials of class $\Phi_{+}$.
 The shapes of the zero-level sets of some polynomials representing these components are shown in Fig.
@@
 This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [29].
-Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or mapped to each other by this automorphism.
+Therefore, there are only two possibilities: the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or are mapped to each other by this automorphism.
 In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$.
@@
 For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$.
-The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\setminus\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
+The basic element of this group is induced from the generator of the group $H^{1}({\mathbb{C}}^{1}\setminus\{0\})\sim{\mathbb{Z}}$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\setminus\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$.
 Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals.
@@
 acts on the space $\Phi_{1}$.
-It preserves the discriminant variety and hence permutes the connected components of its complement.
+It preserves the discriminant variety and permutes the connected components of its complement.
 This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y).
@@
 In particular, it always alters the matching of the vertical and horizontal branches.
-++--++$(a)$
-++--++$(b)$
-++--++$(c)$
-++--++$(d)$
-++--++$(e)$
-++--++$(f)$
-++--++$(g)$
+++-++$(a)$
+++-++$(b)$
+++-++$(c)$
+++-++$(d)$
+++-++$(e)$
+++-++$(f)$
+++-++$(g)$
 Figure 2. Perturbations of $X_{9}^{1}$ singularity
@@
 Let $f$ be a polynomial that satisfies the conditions of this lemma.
-By Lemma 5 of [28], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12).
+By Lemma 5 of [29], the $j$-invariant of its principal homogeneous part is equal to 0.
+By the formula for $j$-invariants of quartic forms (see [18], §1.3), $A=\pm\frac{\sqrt{3}}{2}$ in (12).
 We can assume that the coordinates $(x,y)$ are centered at the $E_{7}$ point.
@@
 The deduction of this proof from Lemma 17 essentially repeats the proof of Proposition 9 with the following simplification.
-In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [28], § 4.5).
+In this case, the neighborhoods of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) have no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [29], § 4.5).
 Therefore, the statement of Proposition 16 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$.
@@
 and
-$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8)$
+$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8),$
 for any $t>0$.
@@
 2.6. Components for the $X_{9}^{2}$ singularities
-The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement in some way.
+The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement.
 All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
 All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$.
-So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
+The entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
 ----++$(a)$
@@
 (21) $\frac{(A^{2}+3)^{3}}{27(A^{2}-1)^{2}}=5.$
-Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [12], p. 160.
+Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [18]. $\Box$
 […]
-The rest of the proof of Proposition 21 repeats the proof of Proposition 3 in § 5.4 of [28]. $\Box$
+The rest of the proof of Proposition 21 repeats the proof of Proposition 9 with the simplification indicated in the proof of Proposition 16.
+Let $f_{0}$ be one of eight polynomials mentioned in in Lemma 22.
+The space $\Phi_{2}$ does not have automorphisms that commute with the Lyashko–Looijenga map and map this polynomial to itself, see the proof of Proposition 3 in § 5.4 of [29].
+Let $f$ and $\tilde{f}$ be two generic polynomials of class $\Phi_{2}$ with the same associated virtual function and sets of critical values and paths.
+Let $I$ be a generic path in $\Phi_{2}$ connecting $f$ with $f_{0}$, and $\tilde{I}$ be a path starting at $\tilde{f}$ and repeating all the critical values and surgeries of the path $I$.
+Then the endpoint $\tilde{f}_{0}$ of this path is well-defined and is also one of these eight points, and the entire path $\tilde{I}$ is mapped to $I$ by the element of the group $G_{0}$ that maps the point $\tilde{f}_{0}$ to $f_{0}$.
+$\Box$
 Proof of Theorem 19. Pictures ($a$) and ($b$) of Fig. 3 are realized, respectively, by the polynomials (22) and (23) with arbitrary $t>0$.
@@
 The virtual functions associated with generic polynomials of these types, the formal graphs of types $J_{10}^{1}$ and $J_{10}^{3}$, and their virtual components are defined in the same way as for the case $X_{9}$, except that the sets of vanishing cycles consist of ten elements.
-Proposition 26 (see [30]).
+The following analog of Propositions 2 and 3 follows from the same considerations as these statements.
+Proposition 26.
+For any class of polynomials $\Omega_{1}$ or $\Omega_{3}$, 1)) if two generic polynomials, $f$ and $\tilde{f}$, of this class belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$ and S(f~)S($\tilde${f}) are the same.
+2)) every virtual function of this class is associated with a generic real polynomial of the same class;
+3)) for any generic polynomial $f$ of this class, every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$.
+$\Box$
+Proposition 27 (see [31], Proposition 9).
 If two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are associated with the same virtual function, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
@@
 Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other.
-On the other hand, the involution (34) moves the component of picture $(f)$ to the same component as (33).
-The following is an invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ from its mirror image f(x,y)f(x,-y).
-The four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement.
+On the other hand, the involution (34) moves the component of picture $(f)$ to the same component as the involution (33).
+Every polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ belongs to a different connected component than its mirror image f(x,y)f(x,-y).
+Indeed, the four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement.
 Let us take an arbitrary point inside each of these four ovals.
@@
 The coefficient $a$ of this equation cannot be zero.
-Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with eight roots.
-Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement.
-Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define the opposite signs.
+Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with at least eight roots.
+Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and remains constant along the connected component of the discriminant complement.
+Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define opposite signs.
 The same is true for polynomials that realize the picture (e~)($\tilde${e}).
@@
 realize picture $(f)$.
-Figure 6.
+Figure 6. Realization of perturbation $(e)$
 To realize the picture $(e)$, we start with the polynomial
@@
 It will also have two critical points in the complex domain.
-Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain.
+Then we add a small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain.
 The zero-level set of the resulting polynomial will have the desired shape shown in picture $(e)$ of Fig.
@@
 4.1.
-Two classes $P_{8}^{1}$ and $P_{8}^{2}$ of function singularities in three variables have the canonical normal form
+The two classes $P_{8}^{1}$ and $P_{8}^{2}$ of function singularities in three variables have the canonical normal form
 (41) $x^{3}+y^{3}+z^{3}-3Axyz,\quad A\neq 1,$
 with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5].
-The zero-level sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$.
+The zero-level sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in the projective plane.
+In the case of two components, one of them is a convex oval.
+The point $(1:1:1)$ lies in the disc bounded by this oval, because both the curve and this point are invariant under the action of the group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$.
 Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin.
-The zero-level set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
+The zero-level set of a polynomial of class $P_{8}^{2}$ consists of a surface like this one and a surface that is homeomorphic to the standard quadratic cone.
 The standard versal deformations of $P_{8}^{1}$ and $P_{8}^{2}$ singularities consist of the polynomials
@@
 Definition 8.
-A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
+A polynomial $f_{\lambda,A}:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of its principal homogeneous part $f_{A}$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
 Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$, respectively.
-It follows easily from Morse theory that the invariant Ind can take only the values 1,0,1,2,1,0,-1,-2, and 3-3 on polynomials of classes $\Xi_{1}$ and $\Xi_{2}$.
-The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
 4.2. The homological invariant of real components
@@
 For $\Xi_{2}$, they are equal to $\mathbb{Z}$} for $i=1$ or $2$ and are trivial for all other $i$.
-Clearly, the invariant Ind is equal to the alternating sum of three elements of the homology index.
+Clearly, the homology index is an invariant of the connected components of discriminant complement, and the invariant Ind is equal to the alternating sum of three elements of the homology index.
+The values that the invariant Ind can take on polynomials of classes $\Xi_{1}$ and $\Xi_{2}$ are described in Propositions 39 and 43 below.
+These values are 1,0,1,2,1,0,-1,-2, and 3-3.
+The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f$ of a singularity (41) with $\mbox{\rm Ind}(f)$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, or three handles, or the disjoint union of the projective plane and the sphere)).
 4.3. The symmetry group and the relation between the real and virtual components
-The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data sets of virtual functions only include the parities of the Morse indices of critical points, and not their integer values.
+The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant polynomials $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data sets of virtual functions only include the parities of the Morse indices of critical points, and not their integer values.
 We also have the following analog of Proposition 3.
@@
 $\Box$
-The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ preserves the normal form (41) and acts on the spaces $\Xi_{1}$ and $\Xi_{2}$.
+The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ acts on the spaces $\Xi_{1}$ and $\Xi_{2}$ preserving the normal form (41) and the discriminant variety.
 This action is generated by the following transformations:
@@
 The first transformation is an involution, and the second is of order three.
-In particular, this group can act on the set of the connected components of the discriminant complements in the spaces $\Xi_{i}$.
+Consequently, this group acts on the set of the connected components of the discriminant complements in the spaces $\Xi_{*}$.
 The transformations (44) preserve all topological characteristics of the functions.
@@
 According to Lemma 38 of [30], there are only six polynomials of the form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$.
-Three of these polynomials (mentioned in formula (27) of [29] with a certain value of the parameter $\beta$) have signature (0,2) at the $A_{2}$ point.
+Three of these polynomials (mentioned in formula (27) of [30] with a certain value of the parameter $\beta$) have signature (0,2) of the second differential at the $A_{2}$ point.
 The remaining three polynomials have the form (46) up to the permutations of coordinates, see formula (28) in [30].
@@
 Lemma 38.
-The set of all polynomials of of the form (42) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
+The set of all polynomials of of the form ((42)) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
 (47) ${\bf T}\left(x^{3}+y^{3}+z^{3}-3(1+\sqrt{3})xyz+(x+y+(1+\sqrt{3})z)^{2}\right),\$
@@
 Proof of Proposition 36.
-The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 36 (respectively, 37) repeats the deduction of Proposition 9 from Lemma 10.
+The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 37 (respectively, 38) repeats the deduction of Proposition 9 from Lemma 10.
 Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$.
@@
 The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2.
-For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$.
+For the first component, this invariant is equal to $(1,0,0)$, and for the second it is equal to $(0,0,1)$.
 According to the last sentence of the previous paragraph, the associated virtual components are also different.
@@
 have $\mbox{Ind}=-3$.
-They are related to each other by the involution (45).
-Therefore, according to the final statement of Proposition 38, they realize two distinct virtual components.
-Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components.
+These polynomials are invariant under the entire group $S(3)$.
+Therefore, the virtual components associated with them are associated only with the real components of discriminant complements that contain these polynomials.
+These two polynomials are related to each other by the involution (45).
+Therefore, according to the final statement of Proposition 39, their associated virtual components are different.
 Thus, Theorem 40 is proved for all virtual components with Ind equal to 1,0,2,1,0,-2, and 3-3.
@@
 2)) For any sufficiently small $\varepsilon>0$, the polynomial
-(52) $F(l)\equiv x^{3}+y^{3}+z^{3}-l^{2}-\varepsilon p_{l}^{2}$
+(52) $x^{3}+y^{3}+z^{3}-l^{2}-\varepsilon p_{l}^{2}$
 has only two critical points in ${\mathbb{R}}^{3}$: a critical point of type $A_{3}$ at the origin, at which this polynomial is equal to
@@
 Denote by $C(l)$ the critical value of the polynomial (52) at its Morse critical point.
-Let $F(l)$ be the polynomial obtained from (52) by adding the constant $C(l)/2$.
+Let $F(l)$ be the polynomial obtained from (52) by subtracting the constant $C(l)/2$.
 Then the polynomial $F(l)$ is non-discriminant and has $\mbox{Ind}=-1$.
@@
 Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement.
+This completes the proof of Theorem 40.
 $\Box$
@@
 In particular, the fundamental group of this component acts on these groups by monodromy.
-For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F_{l}$ is defined by the intersection of this set and the plane $\ker l$.
+For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F(l)$ is defined by the intersection of this set and the plane $\ker l$.
 When $l\in L$ runs a half of the circle $L$, the planes $\ker l$ run a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomials $F(l)$ run a closed cycle in the discriminant complement.
@@
 For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$.
-The homology indices of these components are equal to $(000)$ and $(011)$.
+The homology indices of these components are equal to $(0,0,0)$ and $(0,1,1)$.
 For example, let $f$ be the $S(3)$-invariant polynomial
@@
 has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2.
-Thus, it has $\mbox{Ind}=1$ and homology index $(001)$.
+Thus, it has $\mbox{Ind}=1$ and homology index $(0,0,1)$.
 This polynomial is invariant under the involution (45).
@@
 The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions.
-Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$.
-The involution (45) moves this polynomial to one with $\mbox{HI}=(012)$.
+Therefore, the homology index $\mbox{HI}(f-1)$ is $(1,0,0)$.
+The involution (45) transforms this polynomial into one with $\mbox{HI}=(0,1,2)$.
 In particular, these two polynomials represent two different components.
 They are $S(3)$-invariant.
-Therefore, by Proposition 35, the associated virtual components are also different.
+The associated virtual components are therefore associated with single real connected components and are also different, though their Card invariants are the same.
 According to Proposition 43 and Table 3, these are two components with $\mbox{Card}=94$.
-According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has the homology index $(031)$.
+According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(0,2,0)$ or $(1,3,0)$, while its image under the involution (45) has the homology index $(0,3,1)$.
 Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants.
@@
 By Proposition 40 of [30], the intersection matrix of the canonical basis of vanishing cycles for any generic polynomial of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Coxeter-Dynkin graph shown in
-Figure 7. Coxeter–Dynkin graph for (58)
+Figure 7. Coxeter–Dynkin graph for the polynomial (58)
 Fig.
@@
 12536748
-Figure 8. Coxeter-Dynkin graph for (59)
+Figure 8. Coxeter-Dynkin graph of the matrix (59)
 3) the intersection matrix of the canonically oriented vanishing cycles, ordered by increasing critical values, is as follows:
@@
 According to Proposition 41 of [30], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2.
-Therefore the polynomial $f_{0}$ has no local minima or maxima.
-Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
-Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$.
+Therefore, the polynomial $f_{0}$ has no local minima or maxima.
+Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parameterized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$.
+Using the Lyashko–Looijenga covering, whose existence in the spaces $\Xi_{2}$ is shown in [12], one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$.
 As in the proof of Theorem 10 of [30], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined.
@@
 $\Box$
-By statement 3 of Proposition 42, the involution (45) moves the polynomial $I(1)$ to a polynomial in the different component of the discriminant complement with $\mbox{Card}=60$.
+By statement 3 of Proposition 43, the involution (45) moves the polynomial $I(1)$ to a polynomial in a different component of the discriminant complement with $\mbox{Card}=60$.
 Thus, this component is also represented by an $S(3)$-invariant polynomial.
@@
 Due to the versality of the space $\Xi_{2}$, for any sufficiently small $\varepsilon>0$ we can slightly perturb the polynomial $\hat{f}$ within the space $\Xi_{2}$ in such a way that the resulting polynomial $f_{\varepsilon}$ satisfies the following conditions:
-  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$,
+  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$;
   • the critical values at its two critical points with Morse index 2 are equal to $\omega_{1}$ and $\omega_{1}+\varepsilon^{2}$,
   • the critical values at the non-real critical points are equal to $\omega_{2}+i\varepsilon^{2}$ and $\omega_{2}-i\varepsilon^{2}$.
-Denote the polynomial $f_{\varepsilon}-1$ by $\bar{f}$.
-It belongs to the same component of the discriminant complement as the polynomial $\hat{f}-1$.
-The canonical system of paths in ${\mathbb{C}}^{1}$ defining the frame of vanishing cycles in $H_{2}(\bar{f}_{\varepsilon}^{-1}(0))$ then looks as shown in Fig.
-9 (left), where the numbers 1–8 denote the order of basic vanishing cycles corresponding to the critical values.
 […]
@@
 Figure 9. System of paths for polynomials $\bar{f}$ (left) and $f!$ (right)
-The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program.
+Denote the polynomial $f_{\varepsilon}-1$ by $\bar{f}$.
+It belongs to the same component of the discriminant complement as the polynomial $\hat{f}-1$.
+The canonical system of paths in ${\mathbb{C}}^{1}$ defining the frame of vanishing cycles in $H_{2}(\bar{f}^{-1}(0))$ then looks as shown in Fig.
+9 (left), where the numbers 1–8 denote the order of basic vanishing cycles corresponding to the critical values.
+The virtual function associated with the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program.
 Based on the information on the polynomial $\bar{f}$, let us try to find it in this list.
@@
 According to Proposition 43 of [30], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
-Proposition 43 of [29] implies that all real polynomials of class $\Xi_{2}$ with eight critical points, one of which is a minimum, have the Coxeter-Dynkin graph shown in Fig. 7, in such a way that
+Proposition 43 of [30] implies that the intersection form of any real polynomial of class $\Xi_{2}$ with eight critical points, one of which is a minimum, is defined by the Coxeter-Dynkin graph shown in Fig.
+7, in such a way that
   • the lowest vertex in this figure corresponds to the minimum point, the vertices in the middle row correspond to critical points with Morse index 1, and the upper row vertices correspond to points with Morse index 2;
@@
 The virtual function associated with the polynomial $f!$ is determined by this graph supplied with the information on the Morse indices of critical points corresponding to its vertices, and with the numbering of these vertices corresponding to the ascending order of critical values, i.e., the correspondence of these vertices with the black points in Fig.
-9 right.
+9 (right).
 The graph and the Morse indices are known to us, while concerning the numbering we know the following.
@@
 Such a transposition is a flip of virtual functions which corresponds to the elementary function surgery when one critical value overtakes another critical value at a distant critical point.
+This surgery is not related with crossing the discriminant.
 Our program, when applied to one of the virtual functions from this set, says that this virtual component has 1216 vertices.
@@
 For each $l\in L$, the polynomial
-(61) $F(l)\equiv x^{3}+y^{3}+z^{3}-6xyz+3l^{2}$
+(61) $F(l)\equiv x^{3}+y^{3}+z^{3}-6xyz+3\sqrt{2}l^{2}$
 has a critical point of class $D_{4}^{-}$ with zero critical value at the origin. All its other critical values at real critical points are positive.
@@
 $\Box$
-The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
+The set of all polynomials $F(l)$, $l\in L$, is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
 Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin.
 These numbers $c(l)$ form a lower semi-continuous function on $L/{\mathbb{Z}}_{2}$.
-By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$.
+By the previous lemma, this function is positive everywhere.
+Therefore, it is separated from zero by a constant $\delta>0$.
 The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant.
 This union is connected and invariant under the action (44) of the group $S(3)$.
-By Lemma 48, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$.
+By Lemma 49, for $l\equiv\frac{x-y}{\sqrt{2}}$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$.
 The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant.
@@
 Components with $\mbox{Card}=1318$ and $844$.
-Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0.
+Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant variety in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0.
 By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$.
@@
 Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum.
-Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin.
-Therefore, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
-(62) $x^{2}y-y^{3}+\varkappa_{1}+\varkappa_{2}x+\varkappa_{3}y+\varkappa_{4}y^{2}$
-of the sample polynomial $x^{2}y-y^{3}$ of class $D_{4}^{-}$ in two variables with parameters $\varkappa_{i}$.
-Denote the fiber of this tubular neighborhood over the point $\{F(l)\}$, $l\in L$, by $U(l)$.
-For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
+Suppose this tubular neighborhood is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is the parameter space of a versal deformation of the singularity of the function $F(l)$ at the origin.
+Then, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
+(62) $X^{3}-3XY^{2}+Z^{2}+\varkappa_{1}+\varkappa_{2}X+\varkappa_{3}Y+\varkappa_{4}(X^{2}+Y^{2})$
+of the sample polynomial $X^{3}-3XY^{2}+Z^{2}$ of class $D_{4}^{-}$ with parameters $\varkappa_{i}$.
+Definition 12.
+For each $l\in L$, denote the fiber of this tubular neighborhood over the point $\{F(l)\}$ by $U(l)$, and define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
   • (a)
-    $f$ has four real critical points near the origin obtained by splitting the critical point of the polynomial $F(l)$: specifically, one minimum and three points with Morse index 1,
+    $f$ has four real critical points near the origin that are obtained by splitting the critical point of the polynomial $F(l)$: specifically, one minimum and three critical points with Morse index 1,
   • (b)
     $\mbox{Card}(f)=844$.
-Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
+Similarly, define the subset $W(l)\subset U(l)$ as the union of polynomials $f$ satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
+Remark 51.
+According to Proposition 43 and Lemma 50, the set $V(l)\cup W(l)$ is the union of all polynomials in $U(l)$ with $\mbox{Ind}=-1$ that satisfy condition (a).
 Lemma 52.
@@
 By the versality property, the discriminant is equisingular along the smooth stratum $\{D_{4}^{-}\}$.
-Therefore, it is sufficient to prove Lemma 50 for only one particular $l\in L$, say for $l\equiv x-y$.
+Therefore, it is sufficient to prove Lemma 52 for only one particular $l\in L$, say, for $l\equiv\frac{x-y}{\sqrt{2}}$.
 The union of subsets $V(l)$ over all $l\in L$ is $S(3)$-invariant.
@@
 The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$.
-In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig.
+In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse polynomials looks as shown in Fig.
 11 (left), see [3], § 21.3.
@@
 The larger part consists of polynomials with only two real critical points.
-Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one critical point that is either a minimum or a maximum, depending on the pyramid.
+Two symmetric “pyramids” are filled with polynomials (62) that have three critical points with Morse index 1 and one critical point with Morse index 0 or 2, depending on the pyramid.
 We are only interested in the polynomials (62) with a minimum point.
-For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values consists of five intervals.
+For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values or parameters consists of five intervals.
 Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
-(63) $x^{2}y-y^{3}+\varkappa_{2}x+\varkappa_{4}y+\varkappa_{4}y^{2},$
+(63) $X^{3}-3XY^{2}+Z^{2}+\varkappa_{2}X+\varkappa_{3}Y+\varkappa_{4}(X^{2}+Y^{2}),$
 i.e., it can belong to one of the five intervals into which these critical values divide the line $\mathbb{R}$}.
@@
 The set of these non-generic polynomials (63) divides the pyramid into six parts.
-The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig.
+The intersection of the pyramid with a plane orthogonal to its axis looks as the curvilinear triangle shown in Fig.
 11 (center).
 The non-generic polynomials form three segments in it.
-In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints.
-The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig.
+In particular, the central point of this triangle is represented by the polynomial of the form (63) with $\varkappa_{2}=\varkappa_{3}=0$ and equal critical values at all three points with Morse index 1.
+These critical values are equal to $\frac{4}{27}\varkappa_{4}^{3}$.
+The orthogonal plane sections of the pyramid are defined by the conditions $\varkappa_{4}=C$ for different constants $C>0$.
+The set of all polynomials (62) having a minimum point and three critical points with critical value zero consists of one curve in ${\mathbb{R}}^{4}$ satisfying the conditions
+(64) $\varkappa_{2}=\varkappa_{3}=0,\quad\varkappa_{1}=-\frac{4}{27}\varkappa_{4}^{3},\quad\varkappa_{4}>0.$
+The zero-level sets of all such polynomials in the plane $\{Z=0\}$ look as in Fig. 10 (right).
+Choose an arbitrary polynomial inside the pyramid and label its three critical points of index 1 as $a,b,$ and $c$.
+This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid.
+Each of the six domains, into which the curvilinear triangle in Fig.
+11 (center) is split by the segments, is characterized by the order of the critical values of the corresponding polynomials at the critical points labeled in this way.
+The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of nine connected domains: the number of domains with 0, 1, 2, 3, or 4 negative critical values is equal to 1, 1, 3, 3, or 1, respectively.
+We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value.
+The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three sectors of the pyramid: their intersections with a section of the pyramid are shown in Fig.
 11 (right).
-These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis.
-An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
-Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$.
-This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid.
-Each of the six cells, into which the curvilinear triangle in Fig.
-11 (center) is split by the segments, is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
-The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of eight connected domains.
-We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value.
-The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three shares of the pyramid: their intersections with a section of the pyramid are shown in Fig.
-11 (right).
-Each share, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the saddlepoint with that marking is the highest.
-The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding share shown in Fig.
+Each sector, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the critical point with that marking is the highest.
+The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding sector shown in Fig.
 11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values.
 Clearly, all three of these domains are contractible.
+The set of polynomials (64) belongs to the boundaries of all these domains.
+Thus, we have the following statement.
+Proposition 53.
+There are natural one-to-one correspondences between the following three-element sets:
+  1. (1)
+    The set of connected components of the space of non-discriminant polynomials ((62)) having a minimum point and three critical points with Morse index 1, only one of whose critical values is positive;
+  2. (2)
+    The set of critical points with Morse index 1 of an arbitrary polynomial ((62)) having a minimum point;
+  3. (3)
+    The set of pairs of lines of the zero-level set of the polynomial $X^{3}-3XY^{2}+Z^{2}$ in the plane $\{Z=0\}.$
+Proof.
+The sets mentioned in item 2 for all polynomials (62) having a minimum point are in a standard one-to-one correspondence with each other.
+This correspondence is defined by the continuity within the contractible set of all such polynomials.
+Thus, it is sufficient to define the correspondence $1\leftrightarrow 2$ assuming that the polynomial mentioned in item 2 belongs to this component.
+In this case, we associate with the connected component the critical point with a positive critical value.
+Also, it is sufficient to define the correspondence $2\leftrightarrow 3$ assuming that the polynomial (62) satisfies the conditions (64).
+In this case, the zero-level set of such a polynomial in the plane $\{Z=0\}$ looks as in Fig.
+10 (right).
+The desired correspondence associates with each critical point the pair of lines of Fig.
+10 (left) that are parallel to the lines in Fig.
+10 (right) passing through this point.
+$\Box$
 Proof of Lemma 52.
@@
 The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant.
-Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph.
-These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials of class $\Xi_{2}$ having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
-Consider again the case $l\equiv x-y$.
-Let $f$ be a point in one of these three domains in the corresponding slice $U(l)$.
-We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points.
+This map defines a bijection between the union of three domains mentioned in item 1 of Proposition 53 and the union $V(l)\cup W(l)$, see Definition 12.
+Consider again the case $l\equiv\frac{x-y}{\sqrt{2}}$.
+Let $f$ be a point of the set $V(l)\cup W(l)$ it the corresponding slice $U(l)$.
+We assume that this point $f$ is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points.
 Then, acting as in the proof of Lemma 49, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\xi$ and $\tilde{\omega}_{1}$.
@@
 Only one of these three critical points has a positive critical value.
-Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex.
-This choice explicitly determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
-Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant.
-It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig.
-7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases.
+Depending on the component containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like rightmost vertex.
+This choice uniquely determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement).
+Indeed, the information on this polynomial known to us determines all the elements of the virtual function up to the surgeries not crossing the discriminant: all these surgeries consist in collisions of the critical values at distant critical points.
+Starting from an arbitrary element of this virtual component, our program listed all of its elements, in particular found its Card invariant.
+It turns out that $\mbox{Card}=844$ when the rightmost vertex in Fig.
+7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other two cases.
 $\Box$
@@
 Lemma 54.
-The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
+The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the orthogonal four-dimensional slices $U(l)$ of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
 $l=0$
 $l=0$
-Figure 12. The plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
-Proof. The zero-level set of the polynomial $F(l)$ in the plane $\{l=0\}$ consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set
+Figure 12. Plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
+Proof. For any $l\in L$, the subset in ${\mathbb{R}}^{3}$ defined by equations
+(65) $l=0,F(l)=0$
+consists of three lines, see Fig. 10. These lines have different nature. Two of them belong to the component of the set where
 (66) $x^{3}+y^{3}+z^{3}-6xyz=0$
@@
 The third line belongs to the component that is homeomorphic to the plane, see Fig.
-12 left (in which we schematically depicted the set (64) by the union of the quadratic cone and the horizontal plane).
-The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree.
-The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin.
-The set $\{F(l)=0\}$ in this surface consists of three curves that are distinguished by the conditions
-(65) $\tilde{p}=0,\ \tilde{p}=\tilde{q},\ \mbox{and}\ \tilde{p}=-\tilde{q}$
-and are tangent to the three lines considered in the previous paragraph.
-Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0.
-This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig.
-11 left.
-Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig.
-10 right.
-The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$.
-In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$.
-The three critical points with Morse index 1 occur at their intersections, see Fig.
-10 (right).
-Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative.
-Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
-As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute.
+12 (left) (in which we schematically depicted the set (66) by the union of the quadratic cone and the horizontal plane).
+Proposition 53 reduces the monodromy action on the set of components of the space $V(l)\cup W(l)\subset U(l)$ to the action on the set of pairs of these lines.
+For any $\varkappa=(\varkappa_{1},\dots,\varkappa_{4})$, denote by $g_{\varkappa}$ the corresponding polynomial (62).
+According to the definition of versal deformations, there is a neighborhood $\tilde{U}$ in the space ${\mathbb{R}}^{4}$ of parameters $\varkappa$, a diffeomorphism $\theta:\tilde{U}\to U(l)$, and a family of local diffeomorphisms $\Gamma_{\varkappa}$ from a neighborhood of the origin in the space of arguments $(X,Y,Z)$ of polynomials (62) to a neighborhood of the origin in the space of arguments $(x,y,z)$ of polynomials of class $\Xi_{2}$, such that the diffeomorphisms $\Gamma_{\varkappa}$ depend smoothly on the parameters $\varkappa\in\tilde{U}$, and for any $\varkappa\in\tilde{U}$ we have
+(67) $g_{\varkappa}(X,Y,Z)\equiv f_{\theta(\varkappa)}\circ\Gamma_{\varkappa}(X,Y,Z).$
+The diffeomorphism $\Gamma_{0}$ can be considered as a change of local coordinates that reduces the polynomial $F(l)$ to the normal form $X^{3}-3XY^{2}+Z^{2}$.
+The kernel of the second differential of a polynomial is independent of the coordinates, therefore, this diffeomorphism maps the plane $\{Z=0\}$ to a surface that is tangent to the plane $\{l=0\}$.
+It also maps the three lines distinguished by the equation $g_{0}=0$ in the plane $\{Z=0\}$ to three curves that are tangent to three lines forming the set (65).
+The map $\theta$ defines a bijection between the set of polynomials $g_{\varkappa}\subset\tilde{U}$ having a minimum point and $\mbox{Ind}=-1$, on the one hand, and the union of the subsets $V(l)$ and $W(l)$ of the slice $U(l)$, on the other hand.
+According to Proposition 53, we obtain a one-to-one correspondence between the pairs of lines forming the set (65) and the set of connected components of the space $V(l)\cup W(l)$.
+This correspondence does not depend on the choice of maps $\Gamma_{\varkappa}$ and $\theta$ reducing the versal deformations (62) and $U(l)$ to one another.
+Indeed, let $\bar{f}\in U(l)$ be an arbitrary polynomial very close to $F(l)$, which has a minimum point and three critical points $a,b,$ and $c$ with Morse index 1 and critical value 0.
+(Such points form a single curve in $U(l)$ which is the image of the curve (64) under the map $\theta$.) To each of the points $a,b,c$, a connected component of the set $V(l)\cup W(l)$ corresponds: it contains all polynomials neighboring to $\bar{f}$ and such that the critical values of their critical points neighboring to the selected critical point of $\bar{f}$ is positive, and the critical values at the other two critical points with Morse index 1 are negative.
+Also, a pair of lines forming the set (65) corresponds to each point $a,b,$ and $c$.
+These two lines are nearly parallel to the two edges of the triangle $\triangle_{abc}\subset{\mathbb{R}}^{3}$ that contain this point.
+The composition of these two correspondences defines an invariant description of the correspondence between the components of $V(l)\cup W(l)$ and the pairs of lines in set (65).
+Thus, the monodromy action of the circle $L/{{\mathbb{Z}}_{2}}$ on the set of pairs of lines in (65) coincides with its action on the set of corresponding connected components of the set $V(l)\cup W(l)$.
+As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these lines in $\ker l$ move in such a way that one pair of them returns to itself, while the other two pairs permute.
 Thus, the monodromy over this circle acts in the same way on the set of three related domains in $U(l)$.
@@
 Lemma 55.
-Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig.
-12 right.
+Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary non-zero linear function, whose kernel is tangent to the component of the surface ((66)) that is homeomorphic to the quadratic cone, see Fig.
+12 ((right)).
 The polynomial
@@
 (69) $\zeta^{2}+\xi^{2}\eta-\eta^{4}$
-in some local coordinates at this point. The critical values at all other real critical points of ((66)) are positive.
+in some local coordinates at this point. The critical values of the polynomial ((68)) at all its other real critical points are positive.
 Proof.
-All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12].
+All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ or $E_{6}$, see [21], [12].
 The restriction of the polynomial ((68)) to the plane $\ker l$ is a degenerate cubic form.
@@
 (70) $l\equiv 2x-5y+11z,$
-the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (66) to a polynomial whose lower weighted homogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$.
+the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (68) to a polynomial whose lower weighted homogeneous part is $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$.
 A dilation of coordinates reduces it to the form (69), and technique of [3], §12, removes all higher weighted homogeneous terms of the Taylor expansion.
-The set of all planes in ${\mathbb{R}}^{3}$ that are tangent to the same component of the zero-level set of (64) is connected.
-Therefore, the singularity will be of the same class for all other functions (66) corresponding to these planes.
+The set of all polynomials (68) such that the plane $\ker l$ is tangent to the same component of the zero-level set of (66) is connected.
+Therefore, the singularity will be of the same class for all other functions (68).
 The positivity of other critical values follows from the same arguments as in the proof of Lemma 50.
@@
 Lemma 56.
-For any linear function $l$ as in Lemma 52, the set of polynomials with $\mbox{Ind}=-3$ in a small neighborhood of the corresponding polynomial ((66)) consists of a single connected component.
+For any linear function $l$ as in Lemma 55, the set of polynomials with $\mbox{Ind}=-3$ has exactly one local connected component near the corresponding polynomial ((68)) in the space $\Xi_{2}$.
 Proof.
-By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$.
+By the versality of the deformation (42), the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$.
 Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (68).
@@
 The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [27].
-For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$: it is the component containing the perturbations shown in Fig.
-13.
-According to Theorem 1.7 of [16], this component is contractible.
-$\Box$
+For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$.
 --$\Rightarrow$
@@
 Figure 13. Perturbation of a $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
+Specifically, in the case of any singularity in two variables that is equivalent to $\xi^{2}\eta-\eta^{4}$ it is the component containing the perturbations whose zero-level sets are as shown in Fig.
+13 (right).
+$\Box$
 Lemma 57.
@@
 Proposition 58.
-Let the linear function $l$ be given by ((68)). Then the set considered Lemmas 53, 54 contains a generic polynomial with the following properties:
-  • it has exactly six real critical points: three with Morse index 1 and three with Morse index 2,
-  • the critical values at all critical points with Morse index 1 are lower than all critical values at the critical points with Morse index 2,
+Let the linear function $l$ be given by ((70)). Then the set considered in Lemmas 56 and 57 contains a generic polynomial with the following properties:
+  • it has exactly six real critical points: three with Morse index 1 and negative critical values, and three with Morse index 2 and positive critical values;
   • the intersection indices of vanishing cycles corresponding to the five lowest real critical values are defined by the matrix
@@
 Proof.
-A polynomial with these properties can be obtained from the perturbation of the polynomial (66) that is shown in the middle part of Fig.
-13 by a minor additional perturbation that properly orders the critical values.
-The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10].
+According to the Gusein-Zade—A’Campo method [10], the intersection matrix of five cycles vanishing in the splitting of $D_{5}^{-}$ singularity shown in Fig.
+13 is given by the Coxeter–Dynkin diagram , where white (respectively, black) circles correspond to the critical points with Morse index 1 and negative critical value (respectively, with Morse index 2 and positive critical value).
+Using the Lyashko–Looijenga map, we can achieve that their critical values are ordered as follows: 53421, which corresponds to the matrix (71).
+All these perturbations can be performed in an arbitrarily small neighborhood of the polynomial (68), so that they will not create real critical points in addition to six critical points of this polynomial.
 $\Box$
 Proof of Lemma 57.
-All these sets corresponding to different $l$ belong to the same connected component of the discriminant complement.
-Therefore, we can assume that $l$ is given by (68) and hence contains a polynomial satisfying the conditions listed in Proposition 55.
-Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions.
-It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
-The set of all polynomials (66) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant.
-According to Lemma 53, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at these points.
-Indeed, this union is the space of a fiber bundle with contractible fibers over the space of such planes.
+All these sets corresponding to different planes $\ker l$ belong to the same connected component of the discriminant complement.
+Therefore, we can assume that $l$ is given by (70), and any neighborhood of the polynomial (68) contains a polynomial satisfying the conditions listed in Proposition 58.
+Our program, upon a special request, listed all virtual functions of $P_{8}^{2}$ class satisfying these conditions.
+There are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
+The set of all polynomials (68) with $\ker l$ tangent to the cone is connected and $S(3)$-invariant.
+According to Lemma 56, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at the points corresponding to these polynomials.
+Indeed, this union is the space of a fiber bundle with connected base and fibers.
 This implies the statement of Theorem 44 for the virtual component with $\mbox{Card}=156$.
 $\Box$
+Conjecture 59.
+The connected components with Card equal to 1216, 1318, 844, or 156 have non-trivial first homology groups.
 5. Petrovskii lacunas
@@
 Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
-If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [23], Chapter V).
+If the wavefront point corresponds to a function with isolated singularity, then whether a neighboring component of the wavefront complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary non-discriminant function from that component (see, for example, [24], Chapter V).
 Versal deformations of stably equivalent function singularities can be naturally identified with each other.
 This identification establishes a one-to-one correspondence between the components of the discriminant complements of these deformations.
-The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number $N$ of variables, and on the parities of the positive inertia indices $i_{+}$ of the quadratic part of the singularity.
+The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number $N$ of variables, and on the parity of the positive inertia index $i_{+}$ of the quadratic part of the singularity.
 For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas.
@@
 Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$.
-The only new local lacuna found is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.
+The only new local lacuna is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.